### v1

# Geometry and cohomology of compactified Deligne–Lusztig varieties

Abstract.
For connected reductive groups, we prove that the cohomology of the structure sheaf and the canonical sheaf for compactified Deligne–Lusztig varieties associated to a tuple $(w_{1},\ldots,w_{r})$ of Weyl group elements depends only on the $F$-support of $(w_{1},\ldots,w_{r})$.
Key words and phrases:
Deligne–Lusztig variety; Finite groups of Lie type; Coherent cohomology
2020 Mathematics Subject Classification:
Primary: 20G40, Secondary: 14G17
1. Introduction
1.1. Background
Let $q$ be a power of a prime $p$ and let $k$ be an algebraic closure of the finite field $\mathbb{F}_{q}$. Let $G_{0}$ be a connected reductive $\mathbb{F}_{q}$-group and $G$ its base change to $k$. Denote by $F:G\to G$ the corresponding $q$-Frobenius. The group of fixed points $G^{F}:=G(\mathbb{F}_{q})$ of $F$ is also known as a finite group of Lie type in the literature.
Let $T_{0}\subset B_{0}$ be a maximal torus and a Borel subgroup of $G_{0}$ and $T\subset B$ their base change to $k$. Denote by $W:=N(T)/T$ the corresponding Weyl group and let $S\subseteq W$ be the set of simple reflections. For any $w\in W$, the associated Deligne–Lusztig variety is denoted by $X(w)$, and there is a smooth compactification $\overline{X}(w)$ associated to each reduced expression of $w$. In [DL], Deligne and Lusztig considered the $\ell$-adic cohomology of $X(w)$ and Galois coverings of $X(w)$ to study the irreducible $G^{F}$-representations over fields of characteristic zero.
The de Rham and mod $p$ cohomology of $X(w)$ and $\overline{X}(w)$ are less understood than their $\ell$-adic cohomology. However, when $G={\rm GL}_{2}$ and $w$ is the Coxeter element, the de Rham cohomology of $\overline{X}(w)$ is not mysterious because $\overline{X}(w)$ is isomorphic to the projective line.
The simplicity in the situation of $\overline{X}(w)$ for $G={\rm GL}_{2}$ does not carry over to higher dimensions. The smooth compactification $\overline{X}(w)$ is constructed in loc. cit. as the transversal intersection between the graph of the Frobenius and the desingularisation of the Schubert variety associated to $w$ constructed by Demazure and Hansen. Unlike the desingularisation of Schubert varieties, $\overline{X}(w)$ is not naturally an iterated $\mathbb{P}^{1}$-fibration. Thus the cohomology groups of $\overline{X}(w)$ cannot be computed in the same way.
In [Wan], we described the cohomology of the structure sheaf of $\overline{X}(w)$ for any $w\in W$ in the case of $G={\rm GL}_{n}$. Our main theorem showed that the cohomology of the structure sheaf of $\overline{X}(w)$ is always isomorphic to that of $\overline{X}(c)$, where $c$ is a Coxeter element in a parabolic subgroup of $W$ with ${\text{supp}}(c)={\text{supp}}(w):=\{s\in S|s\leq w\}$. As a consequence of our main theorem, we were able to describe the compactly supported integral $p$-adic cohomology of any Deligne–Lustig variety $X(w)$ for ${\rm GL}_{n}$.
1.2. Main results
The aim of this paper is to extend the main theorem of [Wan] to all connected reductive groups $G_{0}$.
In this general setting, we encounter a special class of elements in $W$ which enjoy nice properties [GKP], [HN]. Namely, they are the minimal length elements in the $F$-conjugacy classes of $W$ (cf. Section 4.2). In any given $F$-conjugacy class $C$, we denote by $C_{\rm min}$ the subset of minimal length elements. For example, if $G$ is ${\rm GL}_{n}$ or an inner form of ${\rm GL}_{n}$, such elements are precisely the Coxeter elements in parabolic subgroups of $W$.
We show that the study of the cohomology of the structure sheaf and canonical sheaf of $\overline{X}(w)$ for any $w\in W$ can always be reduced to that of some minimal length element in an $F$-conjugacy class.
Theorem 1.1.
Let $w\in W$. There exists an $F$-conjugacy class $C$ in $W$ such that for any $x\in C_{\rm min}$, there are $G^{F}$-equivariant isomorphisms
$H^{k}\Big{(}{\overline{X}(w)},\mathcal{O}_{{\overline{X}(w)}}\Big{)}\overset{\sim}{\longrightarrow}H^{k}\Big{(}{\overline{X}(x)},\mathcal{O}_{{\overline{X}(x)}}\Big{)}$
and
$H^{k}\Big{(}{\overline{X}(w)},\omega_{{\overline{X}(w)}}\Big{)}\overset{\sim}{\longrightarrow}H^{k}\Big{(}{\overline{X}(x)},\omega_{{\overline{X}(x)}}\Big{)}$
for all $k\geq 0$. Moreover, there are $G^{F}$-equivariant isomorphisms
$H^{k}\Big{(}{\overline{X}(x)},\mathcal{F}\Big{)}\overset{\sim}{\longrightarrow}{\rm Ind}_{P_{I}(\mathbb{F}_{q})}^{G(\mathbb{F}_{q})}H^{k}\Big{(}\overline{X}_{L_{I}}(x),\left.\mathcal{F}\right|_{\overline{X}_{L_{I}}(x)}\Big{)},$
for all $k\geq 0$, where $I$ is the $F$-support of $x$, $P_{I}$ and $L_{I}$ are the associated standard parabolic and Levi subgroup, and $\mathcal{F}$ is the structure sheaf (resp. canonical sheaf) of ${\overline{X}(x)}$.
One should note that the $F$-conjugacy class $C$ in the above theorem may not contain $w$. However, for any given $w$, one can determine explicitly the minimal length elements in this $F$-conjugacy class by following our recipe given in the proof of 5.1.
Smooth compactifications of $X(w)$ given by different reduced expressions of $w$ are not necessarily isomorphic. To avoid this ambiguity, we work with Deligne–Lusztig varieties $X(w_{1},\ldots,w_{r})$ and compactified Deligne–Lusztig varieties $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ associated to a tuple of Weyl group elements cf. 2.2. These notions also make the construction of partial compactifications of Deligne–Lusztig varieties more straightforward and uniform. One can always recover statements for classical Deligne–Lusztig varieties by setting $r=1$.
In this generality, we give a connectedness criterion for compactified Deligne–Lusztig varieties $X(\underline{w_{1}},\ldots,\underline{w_{r}})$, generalising a result of Bonnafé and Rouquier [BR]. We also show that they have $F$-rational singularities and thus are normal, Cohen–Macaulay, and pseudo-rational. Furthermore, we provide a list of partial compactifications $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ of Deligne–Lusztig varieties which are smooth.
The structure of the paper is as follows. We fix our notations for Deligne–Lusztig varieties and compactified Deligne–Lusztig varieties and describe their geometric properties in Section 2. In Section 3, we prove that $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is smooth when each $w_{1},\ldots,w_{r}$ comes from a braid relation in the Weyl group $W$, and we establish relations between cohomology groups for the structure sheaf and canonical sheaf on the associated compactified Deligne–Lusztig varieties.
Section 4 contains the constructions of morphisms which are $\mathbb{P}^{1}$-fibrations between certain compactified Deligne–Lusztig varieties. They are associated with $F$-cyclic conjugations in the Weyl group. In 4.6, we relate the cohomology of the structure sheaf (resp. canonical sheaf) between these compactified Deligne–Lusztig varieties. Furthermore, we deduce that the cohomology of the structure sheaf (resp. canonical sheaf) for $\overline{X}(w)$ and $\overline{X}(F_{1}^{m}(w))$ are isomorphic for all $m\geq 0$, where $F_{1}$ is a partial Frobenius morphism (2.5). The key observation in this section is that the partial Frobenius morphism $F_{1}$ induces an automorphism of the Weyl group, even though it has no well-defined inverse as a morphism of schemes.
In Section 5, we begin with proving the statement in 5.1 under the assumption that $F(S)=S$. The remaining cases come from Suzuki and Ree groups, which are treated in Section 5.2. Finally, the theorem is proved in full generality in Corollary 5.3.
Acknowledgements
The author thanks Jochen Heinloth and Georg Linden for their comments on an earlier draft of this paper. Thanks to Ulrich Görtz for helpful conversations.
Parts of the work are carried out when the author was a member of the trimester program "The Arithmetic of the Langlands Program" at the Hausdorff Research Institute for Mathematics in Bonn and when the author was a member of the Research Training Group 2553 in Essen, both funded by the DFG (Deutsche Forschungsgemeinschaft).
2. Compactified Deligne–Lusztig varieties
We begin with fixing notations related to Deligne–Lusztig varieties associated to a tuple of Weyl group elements. By adapting classical proofs, we give criterions for compactified Deligne–Lusztig varieties to be connected and smooth, respectively.
2.1. Definitions and basic properties
Let $G$, $B$, $T$, $F:G\to G$, $W$, $S$ be as in Section 1.1. Denote by $\ell:W\to\mathbb{N}\cup\{0\}$ the Bruhat length and $\leq$ the Bruhat order on $W$.
Recall that by the Bruhat decomposition, the diagonal $G$-action on the product $G/B\times G/B$ yields a bijection between the Weyl group $W$ and $G\backslash\left(G/B\times G/B\right)$. The orbit corresponding to $w\in W$ is denoted by ${O(w)}$. It can be regarded as a fibre bundle over the flag variety $G/B$ with fibres isomorphic to the Schubert cell $BwB/B$.
Similarly, for a given $r$-tuple $(w_{1},\ldots,w_{r})\in W^{r}$, one can construct an iterated fibre bundle over $G/B$ by taking fibre products of the orbits $O(w_{i})$, $i=1,\ldots,r$, which we denote by
$O(w_{1},\ldots,w_{r}):=O(w_{1})\times_{G/B}\cdots\times_{G/B}O(w_{r}).$
We use such tuple of elements $w_{1},\ldots,w_{r}\in W$ for the definition of partial compactifications. They will in particular be useful for avoiding ambiguity in having different reduced expressions of a Weyl group element.
Definition 2.1.
Let $w_{1},\ldots,w_{r}\in W$. We define the compactified orbit associated to $(w_{1},\ldots,w_{r})$ as the $k$-scheme
$O(\underline{w_{1}},\ldots,\underline{w_{r}}):=\overline{O(w_{1})}\times_{G/B}\cdots\times_{G/B}\overline{O(w_{r})},$
where for each $i=1,\ldots,r$, $\overline{O(w_{i})}$ is the Zariski closure of $O(w_{i})$ in $G/B\times G/B$.
The compactified orbit $O(\underline{w_{1}},\ldots,\underline{w_{r}})$ can be viewed as an iterated fibre bundle over $G/B$ with fibres isomorphic to Schubert varieties $\overline{Bw_{i}B}/B$. Thus $O(\underline{w_{1}},\ldots,\underline{w_{r}})$ is reduced, separated, normal, Cohen Macaulay $k$-schemes locally of finite type. The algebraic group $G$ acts diagonally on $O(w_{1},\ldots,w_{r})$ and $O(\underline{w_{1}},\ldots,\underline{w_{r}})$.
Whenever $(w_{1}^{\prime},\ldots,w_{r}^{\prime})$ is a tuple satisfying $w_{i}^{\prime}\leq w_{i}$ for all $i$, there is a $G$-equivariant embedding $O(w_{1}^{\prime},\ldots,w_{r}^{\prime})\hookrightarrow O(\underline{w_{1}},\ldots,\underline{w_{r}})$ cf.[DMR, Proposition 2.2.16]. Moreover, there is a stratification
$O(\underline{w_{1}},\ldots,\underline{w_{r}})=\bigcup_{w_{i}^{\prime}\leq w_{i},\text{ for all }i}O(w_{1}^{\prime},\ldots w_{r}^{\prime}).$
Denote by $F:G/B\to G/B$ the map induced by the Frobenius $F:G\to G$ and $\Gamma_{F}\subseteq G/B\times G/B$ its graph.
Definition 2.2.
Let $w_{1},\ldots,w_{r}\in W$. We define the Deligne–Lusztig variety associated to $(w_{1},\ldots,w_{r})$ as the fibre product
$X(w_{1},\ldots,w_{r}):=O(w_{1},\ldots,w_{r})\times_{G/B\times G/B}\Gamma_{F}$
with respect to the projection map $\text{pr}_{0,r}:O(w_{1},\ldots,w_{r})\to G/B\times G/B$ and the inclusion map $\Gamma_{F}\hookrightarrow G/B\times G/B$. Similarly, we define the compactified Deligne–Lusztig variety associated to $(w_{1},\ldots,w_{r})$ as
$X(\underline{w_{1}},\ldots,\underline{w_{r}}):=O(\underline{w_{1}},\ldots,\underline{w_{r}})\times_{G/B\times G/B}\Gamma_{F}.$
Deligne–Lusztig varieties $X(w_{1},\ldots,w_{r})$ and compactified Deligne–Lusztig varieties $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ are separated $k$-schemes locally of finite type. The diagonal $G^{F}$-action on $O(w_{1},\ldots,w_{r})$ (resp. $O(\underline{w_{1}},\ldots,\underline{w_{r}})$) induces a diagonal $G^{F}$-action on $X(w_{1},\ldots,w_{r})$ (resp. $X(\underline{w_{1}},\ldots,\underline{w_{r}})$).
Whenever $(w_{1}^{\prime},\ldots,w_{r}^{\prime})$ is a tuple satisfying $w_{i}^{\prime}\leq w_{i}$ for all $i$, there is a $G^{F}$-equivariant embedding $X(w_{1}^{\prime},\ldots,w_{r}^{\prime})\hookrightarrow X(\underline{w_{1}},\ldots,\underline{w_{r}})$. Moreover, we have a stratification
$X(\underline{w_{1}},\ldots,\underline{w_{r}})=\bigcup_{w_{i}^{\prime}\leq w_{i},\text{ for all }i}X(w_{1}^{\prime},\ldots w_{r}^{\prime}).$
When $s_{i}\in S$ for all $i=1,\ldots,r$, the tuple $(s_{1},\ldots,s_{r})$ can be regarded as an element in the free monoid generated by $S$, which we denote by $S^{*}$. By Lemma 2.3 below, $X(\underline{s_{1}},\ldots,\underline{s_{r}})$ is smooth. To simplify the notation, given any $w=s_{1}\cdots s_{m}\in S^{*}$ with $s_{i}\in S,i=1,\ldots m$, we denote $X(s_{1},\ldots,s_{m})$ by $X^{S^{*}}(w)$ and $X(\underline{s_{1}},\ldots,\underline{s_{m}})$ by $\overline{X}^{S^{*}}(w)$. The notations $O^{S^{*}}(w)$ and $\overline{O}^{S^{*}}(w)$ for the orbits are defined analogously.
Furthermore, denote by $\alpha:S^{*}\to W$ be the natural projection morphism of monoids. Let $\preceq$ be the Bruhat order in $S^{*}$ and let $\ell^{S^{*}}:S^{*}\to\mathbb{N}\cup\{0\}$ be the Bruhat length.
Lemma 2.3.
Let $w_{1},\ldots,w_{r}\in W$. The Deligne–Lusztig variety $X(w_{1},\ldots,w_{r})$ is smooth of dimension $\ell(w_{1})+\cdots+\ell(w_{r})$. If for all $i=1,\ldots,r$, the associated Schubert variety $\overline{{B}w_{i}{B}}/{B}$ is smooth, then $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is smooth of dimension $\ell(w_{1})+\cdots+\ell(w_{r})$.
Proof.
In the proof of [DL, Lemma 9.11], Deligne and Lusztig showed that the graph $\Gamma_{F}$ of the Frobenius is transversal to any smooth $G$-equivariant scheme $Y\to G/B\times G/B$. Here let $Y=O(w_{1},\ldots,w_{r})$ and $Y=O(\underline{w_{1}},\ldots,\underline{w_{r}})$, which are $G$-equivariant schemes over $G/B\times G/B$. By construction, $O(w_{1},\ldots,w_{r})$ is an iterated fibration over $G/B$ with fibres being Schubert cells, which are affine spaces. The compactified orbit $O(\underline{w_{1}},\ldots,\underline{w_{r}})$ is an iterated fibration over $G/B$ with fibres being Schubert varieties $\overline{{B}w_{i}{B}}/{B}$. Hence the smoothness of $X(w_{1},\ldots,w_{r})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ follows. ∎
Remark 2.4.
In general, compactified Deligne–Lusztig varieties are $F$-rational and thus pseudo-rational, normal, and Cohen–Macaulay. We will give a proof for this in 2.13.
2.2. The Lang map
There is an alternative description of Deligne–Lusztig varieties via the Lang map. Using this description, we show that $G^{F}$ acts transitively on the set of connected components of $X(w_{1},\ldots,w_{r})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}})$. Denote by $G^{r}$ the $r$-fold product $G\times\cdots\times G$ over ${\rm Spec}\ k$. This is again a $k$-group scheme.
Definition 2.5.
Let $F_{1}:G^{r}\to G^{r}$ be the partial Frobenius morphism given by
$(g_{0},\ldots,g_{r-1})\longmapsto(g_{1},\ldots,g_{r-1},F(g_{0})).$
Let $\mathcal{L}_{1}:G^{r}\to G^{r}$ be the Lang map associated to $F_{1}$ given by $x\mapsto x^{-1}F_{1}(x)$ for $x\in G^{r}$. In other words, we have
$\mathcal{L}_{1}:(g_{0},\ldots,g_{r-1})\longmapsto(g_{0}^{-1}g_{1},\ldots,g_{r-2}^{-1}g_{r-1},g_{r-1}^{-1}F(g_{0})).$
Similarly, we define the partial Frobenius morphism on $\left(G/B\right)^{r}$. Since $B$ is the base change of $B_{0}$ to ${\rm Spec}\ k$, we know that $G/B$ is the base change of $G_{0}/B_{0}$ to ${\rm Spec}\ k$ and has a corresponding $q$-Frobenius map $F:G/B\to G/B$. Denote by $f_{2}:\left(G/B\right)^{r}\to\left(G/B\right)^{r}$ the map given by
$(x_{0},\ldots,x_{r-2},x_{r-1})\longmapsto(x_{0},\ldots,x_{r-2},F(x_{r-1})),$
which is the identity on the first $(r-1)$-entries and Frobenius on the last entry. Denote by $f_{1}:\left(G/B\right)^{r}\to\left(G/B\right)^{r}$ the map
$(x_{0},\ldots,x_{r-1})\longmapsto(x_{1},\ldots,x_{r-1},x_{0}),$
which shifts the coordinates by $1$. The composition $f_{1}^{r}$ is the identity.
Definition 2.6.
We define the partial Frobenius morphism $F_{1}:\left(G/B\right)^{r}\to\left(G/B\right)^{r}$ to be the composition $f_{2}\circ f_{1}$.
The partial Frobenius morphism $F_{1}$ on $\left(G/B\right)^{r}$ is an étale morphism and the composition $F_{1}^{r}$ is the $q$-Frobenius morphism on $\left(G/B\right)^{r}$.
Lemma 2.7.
The Lang map $\mathcal{L}_{1}:G^{r}\to G^{r}$ is surjective and is a Galois cover with Galois group $G^{F}$.
Proof.
Note that $G$ is geometrically connected and smooth, so the product $G^{r}$ is connected and smooth. Observe that the composition of the partial Frobenius $F_{1}^{r}$ yields the Frobenius on $G^{r}$, so [CE, Theorem 7.1] applied to $G^{r}$ yields that $\mathcal{L}_{1}$ is surjective and étale. The kernel of $\mathcal{L}_{1}$ is $\Delta(G^{F})$, where $\Delta:G\to G^{r}$ is the diagonal embedding. Thus $\mathcal{L}_{1}$ is a finite morphism.
Since the diagonal (left) action of $G^{F}$ is a free action and induces automorphisms of the cover, and the number of automorphisms is the same as the degree of $\mathcal{L}_{1}$, we know that $\mathcal{L}_{1}$ is a Galois covering with Galois group $G^{F}$ [SP, 03SF]. ∎
Lemma 2.8.
For any $w_{1},\ldots,w_{r}\in W$, there are $G^{F}$-equivariant isomorphisms
$X(w_{1},\ldots,w_{r})\overset{\sim}{\longrightarrow}\mathcal{L}_{1}^{-1}\left({B}w_{1}{B}\times_{{\rm Spec}\ k}\cdots\times_{{\rm Spec}\ k}{B}w_{r}{B}\right)/({B})^{r}$
and
$X(\underline{w_{1}},\ldots,\underline{w_{r}})\overset{\sim}{\longrightarrow}\mathcal{L}_{1}^{-1}\left(\overline{{B}w_{1}{B}}\times_{{\rm Spec}\ k}\cdots\times_{{\rm Spec}\ k}\overline{{B}w_{r}{B}}\right)/({B})^{r}.$
Furthermore, $G^{F}$ acts transitively on the connected components of $X(w_{1},\ldots,w_{r})$ (resp. $X(\underline{w_{1}},\ldots,\underline{w_{r}})$). In particular, $X(w_{1},\ldots,w_{r})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ are equi-dimensional.
Proof.
We record the proof for $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ as the proof for $X(w_{1},\ldots,w_{r})$ is the same. In this proof, we denote $\mathcal{B}:=\overline{{B}w_{1}{B}}\times_{{\rm Spec}\ k}\cdots\times_{{\rm Spec}\ k}\overline{{B}w_{r}{B}}$. Observe that $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is isomorphic to a closed subscheme of $\left(G/B\right)^{r}$ via the projection ${\rm pr}_{0,\ldots,r-1}:\left(G/B\right)^{r+1}\to\left(G/B\right)^{r}$ to the first $r$-entries. The diagonal right $({B})^{r}$-action on $G^{r}$ is free. The subscheme $\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)$ of $G^{r}$ is stable under the right $({B})^{r}$-action, so we may take the quotient of $\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)$ with respect to the induced free right $({B})^{r}$-action. It follows that $\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$ is a closed subscheme of $\left(G/B\right)^{r}$. The isomorphism above then follows from verifying the definitions cf. [DL]. With the diagonal $G^{F}$-action on $\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$, the isomorphism is $G^{F}$-equivariant.
We show the second assertion for $\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$. Since the quotient map $\pi:\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)\to\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$ is equivariant under the left $G^{F}$-action and surjective, it induces a surjective morphism $G^{F}\backslash(\mathcal{L}_{1}^{-1}(\mathcal{B}))\to G^{F}\backslash\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$. Via Lemma 2.7 we know that
$G^{F}\backslash(\mathcal{L}_{1}^{-1}(\mathcal{B}))\overset{\sim}{\longrightarrow}\mathcal{B}$
is irreducible, so $G^{F}\backslash\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$ is irreducible and $G^{F}$ acts transitively on the set of connected components of $\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$. Therefore $\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$ is equidimensional. ∎
Remark 2.9.
The $G^{F}$-action on $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is not free. In $G/B$, the point ${B}$ already has nontrivial stabiliser $B^{F}=B(\mathbb{F}_{q})$. The point $(B,\ldots,B)$ in $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is also stabilised by $B^{F}=B(\mathbb{F}_{q})$.
2.3. Criterion for connectedness
There is a simple criterion for classical Deligne–Lusztig varieties to be connected, proved by Lusztig, Bonnafé–Rouquier [BR], Görtz [Gör], and many others. In this section we prove that $X(w_{1},\ldots,w_{r})$ (resp. $X(\underline{w_{1}},\ldots,\underline{w_{r}})$) is connected when the tuple $(w_{1},\ldots,w_{r})$ has full $F$-support by modifying the proof of [BR].
To begin, we fix the notations for the set of $F$-orbits in $S$ and the $F$-support for elements in $W$ and $S^{*}$.
Definition 2.10.
The $F$-orbit of $w\in W$ is the set ${{\rm Orb}_{F}(w)}:=\{F^{k}(w)|k\in\mathbb{Z}_{\geq 0}\}$. Under the assumption that $F(S)=S$, we define the set
$S_{F}:=\{{{\rm Orb}_{F}(s)}|s\in S\}$
of $F$-orbits in $S$. For $w\in W$, the $F$-support of $w$ is the set
${{\text{supp}}}_{F}(w):=\{\tilde{s}\in S_{F}|\text{there exists }t\in\tilde{s}\text{ such that }t\leq w\}.$
An element $w\in W$ is called of full $F$-support if ${{\text{supp}}}_{F}(w)=S_{F}$. For a tuple $(w_{1},\ldots w_{r})\in W^{r}$, we denote
${{\text{supp}}}_{F}(w_{1},\ldots,w_{r}):={{\text{supp}}}_{F}(w_{1})\cup\cdots\cup{{\text{supp}}}_{F}(w_{r}).$
Similarly, the $F$-support for any $v\in S^{*}$ with $v=s_{1}\cdots s_{m}$ is defined as
${{\text{supp}}}_{F}(v):={{\text{supp}}}_{F}(s_{1},\ldots,s_{m}).$
Remark 2.11.
We say that $w\in W$ is a Coxeter element when $w=s_{1}\cdots s_{r}$ is of full $F$-support and each $s_{i}\in S$ is from a distinct $F$-orbit.
Proposition 2.12.
[…]
Proof.
Take the product $w_{1}\cdots w_{r}$ in $S^{*}$ and denote it by $v$. This product depends on the choices of the reduced expressions of $w_{1},\ldots,w_{r}$, but for each such $v$, there is an isomorphism
$X^{S^{*}}(v)\overset{\sim}{\longrightarrow}X(w_{1},\ldots,w_{r})$
via projection. It also follows that $X(w_{1},\ldots,w_{r})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ are dense open subschemes of ${\overline{X}^{S^{*}}(v)}$. Thus it suffices to prove connectedness for ${\overline{X}^{S^{*}}(v)}$.
The proof showing that ${\overline{X}^{S^{*}}(v)}$ is connected is analogous to the proof of [BR] for classical Deligne–Lusztig varieties. Suppose ${{\text{supp}}}_{F}(w_{1},\ldots,w_{r})=S_{F}$ and thus ${{\text{supp}}}_{F}(v)=S_{F}$. Let $Z$ be the connected component of ${\overline{X}^{S^{*}}(v)}$ which contains the point $({B},\ldots,{B})$. Denote the stabiliser of $Z$ in $G^{F}$ by ${\rm Stab}(Z)$. Fix $v^{\prime}$ to be of minimal length such that $v^{\prime}\preceq v$ and ${{\text{supp}}}_{F}(v^{\prime})=S_{F}$. In particular, $v^{\prime}$ has the same expression as $\alpha(v^{\prime})\in W$, which is a Coxeter element. This implies that ${\overline{X}^{S^{*}}(v^{\prime})}$ is irreducible [Lus], [BR].
On the other hand, there is a $G^{F}$-equivariant embedding ${\overline{X}^{S^{*}}(v^{\prime})}\to{\overline{X}^{S^{*}}(v)}$. Let $Z^{\prime}$ be the connected component of ${\overline{X}^{S^{*}}(v^{\prime})}$ containing $({B},\ldots,{B})$. Then ${\rm Stab}(Z^{\prime})\subseteq{\rm Stab}(Z)$. However, ${\overline{X}^{S^{*}}(v^{\prime})}$ is irreducible, so ${\rm Stab}(Z^{\prime})=G^{F}$. Thus ${\overline{X}^{S^{*}}(v)}$ is irreducible.
Suppose ${{\text{supp}}}_{F}(w_{1},\ldots,w_{r})=I$ and so ${{\text{supp}}}_{F}(v)=I$. By [DMR, Proposition 2.3.8], we know that there is a $G^{F}$-equivariant isomorphism
${\overline{X}^{S^{*}}(v)}\overset{\sim}{\longrightarrow}G^{F}/U_{I}^{F}\times^{L_{I}^{F}}\overline{X}^{S^{*}}_{L_{I}}(v),$
where $U_{I}$ is the unipotent radical of $P_{I}$ and $L_{I}$ is the Levi subgroup such that $P_{I}=U_{I}\rtimes L_{I}$. Recall that $L_{I}$ is reductive, and $\overline{X}^{S^{*}}_{L_{I}}(v)$ is the compactified Deligne–Lusztig variety for $L_{I}$ corresponding to $v$. Consider the $G^{F}$-equivariant quotient morphism
$G^{F}/U_{I}^{F}\times^{L_{I}^{F}}\overline{X}^{S^{*}}_{L_{I}}(v)\longrightarrow G^{F}/P_{I}^{F},$
whose fibres are isomorphic to $\overline{X}^{S^{*}}_{L_{I}}(v)$. In this case, we know that $v$ has full $F$-support with respect to the Weyl group of $L_{I}$, so ${\overline{X}^{S^{*}}(v)}$ has $|G^{F}/P_{I}^{F}|$ connected components. ∎
2.4. Criterion for smoothness
In [Wan, Proposition A.6], we showed that some compactified Deligne–Lusztig varieties are strongly $F$-regular. This property generalises to arbitrary compactified Deligne–Lusztig varieties.
Proposition 2.13.
For $w_{1},\ldots,w_{r}\in W$, the compactified Deligne–Lusztig variety $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is strongly $F$-regular. In particular, $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is $F$-rational and thus normal and Cohen-Macaulay.
Proof.
The proof is similar to the proof of [Wan, Proposition A.6]. For each $i=1,\ldots,r$, we know that $\overline{{B}w_{i}{B}}$ is strongly $F$-regular [BT, Cor. 4.1, p. 958]. Thus the product
$\overline{{B}w_{1}{B}}\times_{{\rm Spec}\ k}\cdots\times_{{\rm Spec}\ k}\overline{{B}w_{r}{B}}$
is strongly $F$-regular by faithfully flat descent [HH, Theorem 5.5]. Recall that the Lang map $\mathcal{L}_{1}$ is étale, so its closed fibres are smooth and thus strongly $F$-regular cf. [Smi, Ch 3, Remark 1.6, Theorem 4.4(a)]. We may now apply [Abe, Theorem 3.6] and conclude that
$\mathcal{L}_{1}^{-1}\left(\overline{{B}w_{1}{B}}\times_{{\rm Spec}\ k}\cdots\times_{{\rm Spec}\ k}\overline{{B}w_{r}{B}}\right)$
is strongly $F$-regular. Finally, by Lemma 2.8 and the faithfully flat descent of strong $F$-regularity [HH, Theorem 5.5], $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is strongly $F$-regular. Strongly $F$-regular rings are $F$-rational, so $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is normal and Cohen-Macaulay by [HH, Theorem 4.2, 6.27]. ∎
Remark 2.14.
For $w\in W$, the Zariski closure $\overline{X(w)}$ in $G/B$ is not always smooth, but for any $v\in S^{*}$ that gives a reduced expression of $w$, the projection map ${\overline{X}^{S^{*}}(v)}\to\overline{X(w)}$ is a proper birational morphism [DL, Lemma 9.11].
Since $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is normal by 2.13, the connectedness criterion for $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ in 2.12 applies readily for irreducibility.
Since smoothness can be verified étale locally and satisfies faithfully flat descent, the same proof as in 2.13 gives an alternative proof for the smoothness result in Lemma 2.3.
Corollary 2.15.
Let $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ be as in 2.13. If for all $i=1,\ldots,r$, the Schubert variety $\overline{{B}w_{i}{B}}/{B}$ is smooth, then $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is smooth. In particular, ${\overline{X}^{S^{*}}(w)}$ is smooth for all $w\in S^{*}$.
Example 2.16.
If for all $i=1,\ldots,r$, $w_{i}$ is the longest element in $W_{{\text{supp}}_{F}(w_{i})}$, then $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is smooth.
3. Braid relations in the Weyl group and birational morphisms
3.1. Smoothness via braid relations
We consider some specific elements in the Weyl group for which the compactified Deligne–Lusztig variety is smooth. To begin, we recall the relations in the Weyl group of finite reductive groups cf. [GP].
All the relations in the group presentation of the Weyl group (as a finite Coxeter group $(W,S)$) can be written as
$(st)^{m_{st}}=1,\ s,t\in S,$ (3.1)
where $m_{st}\in\mathbb{Z}$ such that $m_{ss}=1$ and $m_{st}\geq 2$ if $s\neq t$. In the case of $s\neq t$, we call (3.1) a braid relation in the Weyl group $W$. The classification of finite groups of Lie type ([DM, Classification 4.3.6]) yields that for $s\neq t$, we have $m_{st}\in\{2,3,4,6\}$.
Remark 3.1.
Whenever $(st)^{m_{st}}=1,\ s,t\in S$, is a braid relation in $W$,
$(ts)^{m_{st}}=1$ (3.2)
also holds. This is a consequence of the relation $ss=1$ for all $s\in S$. We also have relations
$(F(s)F(t))^{m_{st}}=1\hskip 14.22636pt\text{and}\hskip 14.22636pt(F(t)F(s))^{m_{st}}=1$ (3.3)
corresponding to the braid relation (3.1). When there is no ambiguity, we refer to any equation among (3.1), (3.2) and (3.3) as a braid relation in $W$.
Definition 3.2.
Let $w\in S^{*}$. If $w=w_{1}w_{2}$, for some $w_{1},w_{2}\in S^{*}$, we say that $w_{i}$ divides $w$, $i=1,2$. Moreover, we call $w_{1}$ a prefix of $w$ and $w_{2}$ a suffix of $w$.
Definition 3.3.
Let $(st)^{m}=1$, with $\ s,t\in S$ and $s\neq t$, be a braid relation in $W$. We define
$P_{m}(s,t):=st\cdots\in S^{*}$ (3.4)
such that ${{\ell^{S^{*}}}(P_{m}(s,t))}=m$ and $P_{m}(s,t)$ is a prefix of $(st)^{m}$ in $S^{*}$. When there is no ambiguity, we also use $P_{m}(s,t)$ to denote its canonical projection $\alpha(P_{m}(s,t))\in W$.
Lemma 3.4.
Let $w_{1},\ldots,w_{r}\in W$. Suppose that for all $i\in\{1,\ldots,r\}$, we have $w_{i}=P_{m}(s,t)$ for some braid relation in $W$. Then the compactified Deligne–Lusztig variety $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is smooth.
Proof.
By Lemma 2.3 (resp. Corollary 2.15), it suffices to show that for all $i=1,\ldots,r$, the associated Schubert variety $\overline{{B}w_{i}{B}}/{B}$ is smooth.
We show this by using the classification of finite groups of Lie type and the classification of finite Coxeter groups. Denote $I=\{s,t\}\subseteq S$, $W_{I}\subseteq W$ the corresponding subgroup generated by $I$, and $P_{I}\subseteq G$, $L_{I}\subseteq P_{I}$ the parabolic and Levi subgroups respectively. Note that $L_{I}$ is reductive and $P_{I}/{B}$ is isomorphic to $L_{I}/({B}\cap L_{I})$, which is a flag variety and thus is smooth. We proceed by showing that $S(P_{m}(s,t))=P_{I}/{B}$.
Since $W_{I}$ is generated by two elements, the longest element is determined by the braid relation. When $L_{I}$ is of type $A_{1}\times A_{1}$ (resp. ${{\ell^{S^{*}}}(P_{2}(s,t))}=2$), $st\in W_{I}$ is the longest element. When $L_{I}$ is of type $A_{2}$ or ${}^{2}A_{2}$ (resp. ${{\ell^{S^{*}}}(P_{3}(s,t))}=3$), $sts\in W_{I}$ is the longest element. When $L_{I}$ is of type $B_{2}$ or ${}^{2}B_{2}$ (resp. ${{\ell^{S^{*}}}(P_{4}(s,t))}=4$), $stst\in W_{I}$ is the longest element. When $L_{I}$ is of type $G_{2}$ or ${}^{2}G_{2}$ (resp. ${{\ell^{S^{*}}}(P_{6}(s,t))}=6$), $ststst\in W_{I}$ is the longest element. Therefore $S(P_{m}(s,t))=P_{I}/{B}$ for $m=2,3,4,6$. ∎
Definition 3.5.
Denote by $B^{+}$ the monoid generated by $S$ and all the relations of the form
$P_{m}(s,t)=P_{m}(t,s)$
with $s,t\in S$, $t\neq s$, such that $(st)^{m}=1$ is a braid relation in $W$.
More details about this monoid can be found in [GP, §4].
3.2. Higher direct images of the structure sheaf for birational morphisms
Definition 3.6.
For any $v\in S^{*}$ such that $v=P_{m}(s,t)$ for a braid relation (3.1, 3.2, 3.3) in $W$, we set $R_{st}(v):=P_{m}(t,s)$.
Proposition 3.7.
Let $w_{1},\ldots,w_{r}\in W$ such that their corresponding Schubert varieties $\overline{{B}w_{i}{B}}/{B}$ are smooth for all $i=1,\ldots,r$. Suppose for $j\in\{1,\ldots,r\}$, the reduced expression of $w_{j}$ is of the form $w_{j}=P_{m}(s,t)$ for a braid relation in $W$. Then for all $k\geq 0$, there are $G^{F}$-equivariant isomorphisms
$H^{k}\left(X_{1},\mathcal{O}_{X_{1}}\right)\overset{\sim}{\longrightarrow}H^{k}\left(X_{2},\mathcal{O}_{X_{2}}\right)$
and
$H^{k}\left(X_{1},\omega_{X_{1}}\right)\overset{\sim}{\longrightarrow}H^{k}\left(X_{2},\omega_{X_{2}}\right),$
for $X_{i}={\overline{X}^{S^{*}}(w_{1}\cdots w_{r})}$, ${\overline{X}^{S^{*}}(w_{1}\cdots R_{st}(w_{j})\cdots w_{r})}$, $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ and $i=1,2$.
Proof.
By assumption, we have $w_{j}=R_{st}(w_{j})$ in $W$, so
$X(\underline{w_{1}},\ldots,\underline{w_{r}})=X(\underline{w_{1}},\ldots,\underline{R_{st}(w_{j})},\ldots,\underline{w_{r}}).$
For $X_{i}={\overline{X}^{S^{*}}(w_{1}\cdots w_{r})}$, ${\overline{X}^{S^{*}}(w_{1}\cdots R_{st}(w_{j})\cdots w_{r})}$, $X(\underline{w_{1}},\ldots,\underline{w_{r}})$, it follows from the assumptions and Lemma 2.3 that $X_{i}$ is smooth and projective. Note that we have $I:={\text{supp}}_{F}(w_{1}\cdots w_{r})={\text{supp}}_{F}(w_{1}\cdots R_{st}(w_{j})\cdots w_{r})={\text{supp}}_{F}(w_{1},\ldots,w_{r})$, and by 2.12 $X_{i}$ has $|G^{F}/P_{I}^{F}|$ irreducible components.
There are natural projection maps ${\rm pr}_{0,\ell(w_{i})}:\overline{O}^{S^{*}}(w_{i})\to O(\underline{w_{i}})$ for all $i=1,\ldots r$ and ${\rm pr}_{0,\ell(w_{j})}:\overline{O}^{S^{*}}(R_{st}(w_{j}))\to O(\underline{w_{j}})$. Similarly, the projection to the 00-th, $\ell(w_{1})$-th, $\ell(w_{1})+\ell(w_{2})$-th, . . ., and $\ell(w_{1})+\cdots+\ell(w_{r})$-th coordinate yields the vertical maps in the cartesian diagrams
[…]
and
[…]
The maps $f$ and $f^{\prime}$ are $G^{F}$-equivariant proper birational morphisms. It follows from the proof of Zariski’s main theorem ([Har, Corollary 11.4]) that $\phi_{*}\mathcal{O}_{X_{1}}\cong\mathcal{O}_{X_{2}}$, for $\phi=f$ (resp. $\phi=f^{\prime}$), $X_{1}={\overline{X}^{S^{*}}(w_{1}\cdots w_{r})}$ (resp. $X_{1}={\overline{X}^{S^{*}}(w_{1}\cdots R_{st}(w_{j})\cdots w_{r})}$) and $X_{2}=X(\underline{w_{1}},\ldots,\underline{w_{r}})$. Moreover, applying [CR, Corollary 3.2.10] yields $R\phi_{*}\mathcal{O}_{X_{1}}\cong\mathcal{O}_{X_{2}}$. Hence the Leray spectral sequence associated to $\phi$ degenerates. Since $\phi$ is $G^{F}$-equivariant, the isomorphism $\phi_{*}\mathcal{O}_{X_{1}}\cong\mathcal{O}_{X_{2}}$ is $G^{F}$-equivariant, thus functoriality of the associated Leray spectral sequence yields $G^{F}$-equivariant isomorphisms on cohomology groups
$H^{k}\left(X_{1},\mathcal{O}_{X_{1}}\right)\overset{\sim}{\longrightarrow}H^{k}\left(X_{2},\mathcal{O}_{X_{2}}\right),$
for all $k\geq 0$. The $G^{F}$-equivariant isomorphisms for the cohomology of the canonical sheaf follow from an equivariant version of Serre duality [Nir, Corollary 2.10]. ∎
4. $\mathbb{P}^{1}$-fibrations and Cyclic shifting
We construct $\mathbb{P}^{1}$-fibrations between compactified Deligne–Lusztig varieties and relate their cohomology for the structure sheaf and canonical sheaf. Moreover, we use these $\mathbb{P}^{1}$-fibrations to indirectly show that the partial Frobenius on compactified Deligne–Lusztig varieties does not affect the cohomology for the structure sheaf and canonical sheaf.
4.1. $\mathbb{P}^{1}$-fibrations over Deligne–Lusztig varieties
Lemma 4.1.
Let $w_{1},\ldots,w_{r}\in W$ and $s\in S$ such that $F(s)\in S$. Then there is a $G^{F}$-equivariant morphism of $k$-schemes
$\pi_{1}:X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})\longrightarrow X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)}),$
which is Zariski-locally a $\mathbb{P}^{1}$-bundle.
Proof.
For any $t\in S$, the projection map $O(\underline{t},\underline{t})\to O(\underline{t})$ given by $(B_{0},B_{1},B_{2})\mapsto(B_{0},B_{2})$ is a $\mathbb{P}^{1}$-bundle. Recall that it is a property of the Bruhat decomposition that if $(B_{0},B_{1})\in O(t)$ and $(B_{1},B_{2})\in O(t)$ then $(B_{0},B_{2})\in O(\underline{t})$, cf. [DL, §1.2]. By setting $t=F(s)$ and base change to $O(\underline{w_{1}},\ldots,\underline{w_{r}})$, we get again a $\mathbb{P}^{1}$-bundle
${\rm pr}_{F(s)}:O(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)},\underline{F(s)})\longrightarrow O(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)}).$
The partial Frobenius map induces a morphism of $k$-schemes
$F_{1}:X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})\longrightarrow X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)},\underline{F(s)}).$
As $X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)},\underline{F(s)})$ is a closed subscheme of $O(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)},\underline{F(s)})$, we may compose ${\rm pr}$ and $F_{1}$. It follows from the construction that
${\rm pr}_{F(s)}(X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)},\underline{F(s)}))\subseteq{\rm pr}_{0,r+1}^{-1}(\Gamma_{F}),$
where ${\rm pr}_{0,r+1}$ is the projection map
${\rm pr}_{0,r+1}:O(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})\longrightarrow G/B\times G/B.$
Thus the image of the composition ${\rm pr}_{F(s)}\circ F_{1}$ is contained in $X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$.
We define
$\pi_{1}:={\rm pr}_{F(s)}\circ F_{1}:X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})\longrightarrow X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)}),$
which is $G^{F}$-equivariant by construction.
For any $(B_{0},B_{1})\in O(\underline{s})$, we have $(FB_{0},FB_{1})\in O(\underline{F(s)})$. Recall that via the projection to the first $(r+2)$-entries $X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$ identifies with a closed subscheme of $(G/B)^{r+1}$. Under this identification, we have an embedding
$\iota:X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})\longrightarrow O(\underline{s})\times_{G/B}X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)}),$
given by $(B_{0},\ldots,B_{r+1})\mapsto((B_{0},B_{1}),(B_{1},\ldots,B_{r+1}))$.
Since $F(s)\in S$, for any $(B_{0},B_{1})\in O(\underline{s})$, if there exists a $B^{\prime}\in G/B$ with $(B^{\prime},FB_{1})\in O(\underline{F(s)})$, then we automatically have $(B^{\prime},FB_{0})\in O(\underline{F(s)})$. Thus $\iota$ is surjective with inverse given by $((B_{0}^{\prime},B_{1}^{\prime}),(B_{1}^{\prime},\ldots,B_{r+1}^{\prime}))\mapsto(B_{0}^{\prime},\ldots,B_{r+1}^{\prime})$. Note that here we have $(B_{r+1}^{\prime},FB_{0}^{\prime})\in O(\underline{F(s)})$. This yields that $X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$ is isomorphic to $O(\underline{s})\times_{G/B}X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$ and that $\pi_{1}$ is the composition of the isomorphism $\iota$ with the projection
$O(\underline{s})\times_{G/B}X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})\longrightarrow X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)}),$
which is a Zariski locally trivial $\mathbb{P}^{1}$-bundle with a trivialising section. ∎
Remark 4.2.
The reason for the fibre product of $X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$ with $O(\underline{s})$ yielding a Deligne–Lusztig variety is that the orbit of pairs $(B_{0},B_{1})$ determines the orbit of $(FB_{0},FB_{1})$. This allows us to verify that the tuples of the Borel subgroups live in the fibre product of the graph of Frobenius and a product of orbits. In general, taking a fibre product of a Deligne-Lusztig variety with $O(\underline{s})$ does not necessarily yield another Deligne–Lusztig variety.
Lemma 4.3.
Let $w_{1},\ldots,w_{r}\in W$ and $s\in S$ such that $F(s)\in S$. There are $G^{F}$-equivariant étale $\mathbb{P}^{1}$-bundles
$\pi_{2}:X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})},\underline{s})\longrightarrow X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}}),$
and
$F\circ\pi_{1}:X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})},\underline{s})\longrightarrow X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)}).$
Proof.
Define $\pi_{2}$ by
$(B_{0},B_{1},\ldots,B_{r+1},FB_{0})\longmapsto(B_{r+1},FB_{1},\ldots,FB_{r+1}).$
This is equivariant under the diagonal $G^{F}$-action. Since $(B_{0},B_{1})\in O(\underline{s})$, we know that $(FB_{0},FB_{1})\in O(F(\underline{s}))$. By construction we have $(B_{r+1},FB_{0})\in O(F(\underline{s}))$, so $(B_{r+1},FB_{1})\in O(F(\underline{s}))$ [DL, §1.2].
Let $Z$ be given by the fibre product
[…]
where $F=F_{1}^{r+1}$ is the Frobenius map. Points in $Z$ are denoted as tuples
$((B_{0},B_{1},\ldots,B_{r+1},FB_{0}),(B_{0}^{\prime},B_{1}^{\prime},\ldots,B_{r}^{\prime},FB_{0}^{\prime})),$
where $(B_{0},B_{1},\ldots,B_{r+1},FB_{0})\in X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})},\underline{s})$ and
$(B_{0}^{\prime},B_{1}^{\prime},\ldots,B_{r}^{\prime},FB_{0}^{\prime})\in X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})})$ such that $FB_{0}^{\prime}=B_{r+1}$, $FB_{1}^{\prime}=FB_{1},\ldots,FB_{r}^{\prime}=FB_{r},F^{2}B_{0}^{\prime}=FB_{r+1}$. Here the only condition for $B_{0}$ is that $(B_{0},B_{0}^{\prime})\in O(\underline{F^{-1}(s)})$.
The morphism $d_{1}$ also fits into another commutative diagram
[…]
where ${\rm pr}_{1}:O(\underline{s})\to G/B$ is given by $(B_{0},B_{1})\mapsto B_{1}$, and $r_{2}$ is given by
$((B_{0},B_{1},\ldots,B_{r+1},FB_{0}),(B_{0}^{\prime},B_{1}^{\prime},\ldots,B_{r}^{\prime},FB_{0}^{\prime}))\longmapsto(B_{0},B_{0}^{\prime}).$
The diagram (4.2) yields an injective morphism
$\iota:Z\longrightarrow O(\underline{s})\times_{G/B}X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})}).$
On the other hand, given $(B_{0}^{\prime},B_{1}^{\prime},\ldots,B_{r}^{\prime},FB_{0}^{\prime})\in X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})})$ and arbitrary $B_{0}\in G/B$ with $(B_{0},B_{0}^{\prime})\in O(\underline{F^{-1}(s)})$, we have $(B_{0},B_{1}^{\prime})\in O(\underline{F^{-1}(s)})$ and $(FB_{0},FB_{0}^{\prime})\in O(\underline{s})$. This gives
$((B_{0},B_{1}^{\prime},\ldots,B_{r}^{\prime},FB_{0}^{\prime},FB_{0}),(B_{0}^{\prime},B_{1}^{\prime},\ldots,B_{r}^{\prime},FB_{0}^{\prime}))\in Z,$
and that $\iota$ is surjective. Hence $\iota$ is an isomorphism and $d_{1}$ is a Zariski $\mathbb{P}^{1}$-bundle. Let $\{U_{\alpha}\}_{\alpha}$ be a Zariksi covering of $X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})})$ over which $d_{1}$ trivialises. Since $F:X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})})\to X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}})$ is surjective and étale, $\{F(U_{\alpha})\}_{\alpha}$ is an étale covering of $X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}})$ over which $\pi_{2}$ trivialises as a $\mathbb{P}^{1}$-bundle.
Let $\{V_{\alpha}\}_{\alpha}$ be a Zariski covering of $X(\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})},\underline{s})$ over which $\pi_{1}$ trivialises. Since $F:X(\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})},\underline{s})\to X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$ is surjective and étale, $\{F(V_{\alpha})\}_{\alpha}$ is an étale covering of $X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$ over which $F\circ\pi_{1}$ trivialises as a $\mathbb{P}^{1}$-bundle.
Remark 4.4.
As morphisms for $k$-schemes, the absolute Frobenius may not have a well-defined inverse, but the induced morphism $F:W\to W$ is an automorphism of the Weyl group. This is one of the key facts used in Lemma 4.3. Our constructions are analogous to the geometric constructions for classical Deligne–Lusztig varieties in the proof of [DL, Theorem 1.6].
Remark 4.5.
The maps constructed in Lemma 4.1 and Lemma 4.3 fit into a commutative diagram
[…]
Although there is a morphism $F_{1}:X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}})\to X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$, the morphism $F_{1}$ does not directly yield a clear relation between the cohomology of the structure sheaf for $X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$. The reason is that $F_{1}$ has degree divisible by $p$, which implies that $F_{1,*}\mathcal{O}_{X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}})}$ is not isomorphic to $\mathcal{O}_{X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})}$.
However, the étale $\mathbb{P}^{1}$-bundles $\pi_{1}$ and $\pi_{2}$ are projective and flat with geometrically connected and geometrically reduced fibres, so we could establish a clear relation between the cohomology of the structure sheaves of the source and the target under pushforward. This leads to the result of the following proposition that relates the cohomology of the structure sheaf (resp. canonical sheaf) between $X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$.
Proposition 4.6.
Let $w_{1},\ldots,w_{r}\in W$ and $s\in S$ such that $F(s)\in S$. There are $G^{F}$-equivariant isomorphisms for all $k\geq 0$,
$H^{k}\big{(}X_{1},\mathcal{O}_{X_{1}}\big{)}\overset{\sim}{\longrightarrow}H^{k}\big{(}X_{2},\mathcal{O}_{X_{2}}\big{)}$
and
$H^{k}\big{(}X_{1},\omega_{X_{1}}\big{)}\overset{\sim}{\longrightarrow}H^{k}\big{(}X_{2},\omega_{X_{2}}\big{)},$
where for $i=1,2$, $X_{i}=X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}})$, $X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$, $X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$, and $X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})},\underline{s})$.
Proof.
We continue to use the notations of Lemma 4.3. By Lemma 2.8, 2.13, and 2.12, the schemes $X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$ are normal, reduced, and equidimensional with $|G^{F}/P_{I}^{F}|$ irreducible components, where $I={\text{supp}}_{F}(s,w_{1},\ldots,w_{r})$.
From the proof of Lemma 4.3 we know that
$d_{1}:Z\to X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})})$
is a Zariski $\mathbb{P}^{1}$-bundle. It follows that all fibres of $d_{1}$ are isomorphic to $\mathbb{P}^{1}$ and that $d_{1}$ preserves connected components. Thus we may assume without loss of generality that $Z$ and $X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})})$ are irreducible. Since $X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})})$ is normal and all fibres of $d_{1}$ are geometrically reduced, a criterion for flatness [GW, Theorem 14.131] implies that $d_{1}$ is flat and thus $\pi_{2}$ is flat by base change.
Similarly, we have a Zariski $\mathbb{P}^{1}$-bundle
$\pi_{1}:X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})},\underline{s})\longrightarrow X(\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})},\underline{s}),$
with $X(\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})},\underline{s})$ normal, so $\pi_{1}$ is flat by loc. cit. Hence the composition
$F\circ\pi_{1}:X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})},\underline{s})\longrightarrow X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{s})$
is flat.
Denote $f:=\pi_{2}$ (resp. $F\circ\pi_{1}$), $X_{2}:=X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}})$ (resp. $X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{s})$) and $X_{1}:=X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})},\underline{s})$. The rest of the proof is similar to [Wan, Proposition 4.4]. Consider the Leray spectral sequence for $\mathcal{O}_{X_{2}}$:
$E_{2}^{i,j}=H^{i}\left(X_{2},R^{j}f_{*}\mathcal{O}_{X_{1}}\right)\implies H^{i+j}\left(X_{1},\mathcal{O}_{X_{1}}\right).$
Since $f$ is proper, and all fibres of $f$ are isomorphic to $\mathbb{P}^{1}_{k}$, we have $f_{*}\mathcal{O}_{X_{1}}\cong\mathcal{O}_{X_{2}}$ [SP, Lemma 0E0L]. This isomorphism is $G^{F}$-equivariant since $f$ is $G^{F}$-equivariant.
For all $x\in X_{2}$, we have $X_{1}\times_{X_{2}}x\cong\mathbb{P}^{1}_{k}$, so $H^{j}(X_{1}\times_{X_{2}}x,\mathcal{O}_{X_{1}\times_{X_{2}}x})=0$ for all $j>0$. Hence the stalk of $R^{j}f_{*}(\mathcal{O}_{X_{1}})$ at $x$ is 00 for all $x\in X_{2}$ and $j>0$ by Grauert’s theorem cf. [Vak, Theorem 25.1.6]. We may conclude that $R^{j}f_{*}\mathcal{O}_{X_{1}}=0$ for all $j>0$. Therefore the Leray spectral sequence degenerates and we have isomorphisms
$H^{i}\left(X_{2},\mathcal{O}_{X_{2}}\right)\overset{\sim}{\longrightarrow}H^{i}\left(X_{1},\mathcal{O}_{X_{1}}\right),$
for all $i\geq 0$, which are $G^{F}$-equivariant by functoriality of the Leray spectral sequence.
The same argument applies for the Zariski locally trivial $\mathbb{P}^{1}$-bundle
$\pi_{1}^{\prime}:X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})\longrightarrow X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)}).$
The $G^{F}$-equivariant isomorphisms for the cohomology of $\omega_{X_{1}}$ and $\omega_{X_{2}}$ follow from Serre duality for Deligne–Mumford stacks [Nir, Corollary 2.10].
4.2. $F$-conjugacy classes and $F$-cyclic shifts
Two elements $w,w^{\prime}\in W$ are $F$-conjugate if $w^{\prime}=v^{-1}wF(v)$ for some $v\in W$. The $F$-conjugacy classes in $W$ give a partition of $W$. An element $w\in W$ is of minimal length if it has the smallest Bruhat length in its $F$-conjugacy class.
Minimal length elements were studied by Geck, Kim, and Pfeiffer [GKP], as well as He and Nie [HN]. Under the assumption $F(S)=S$, these two papers showed that for any $w\in W$, there exists a minimal length element $w^{\prime}$ in the $F$-conjugacy class of $w$ such that $w\to w^{\prime}$ (cf. 4.7). In combination with 4.6, this result gives an inductive approach for studying cohomology groups of Deligne–Lusztig varieties. Let us recall the definition of $w\to w^{\prime}$.
Definition 4.7.
We write $w\to w^{\prime}$ if there exist $w_{1},\ldots,w_{r}\in W$ and $s_{1},\ldots,s_{r-1}\in S$ such that $w_{1}=w,w_{r}=w^{\prime}$, $w_{i}=s_{i-1}w_{i-1}F(s_{i-1})$, and $\ell(w_{i})\leq\ell(w_{i-1})$ for all $i=2,\ldots,r$. We say that $w$ and $w^{\prime}$ are conjugate by $F$-cyclic shifts and write $w\approx w^{\prime}$ if there exist $w_{1},\ldots,w_{r}\in W$ and $s_{1},\ldots,s_{r-1}\in S$ such that $w_{1}=w,w_{r}=w^{\prime}$, $w_{i}=s_{i-1}w_{i-1}F(s_{i-1})$, and $\ell(w_{i})=\ell(w_{i-1})$ for all $i=2,\ldots,r$.
Remark 4.8.
Let $w,w^{\prime}\in W$ such that $w\approx w^{\prime}$. If $v\in S^{*}$ gives a fixed expression for $w$, then $w\approx w^{\prime}$ guarantees that there exists $w_{1},w_{2}\in S^{*}$ such that $v=w_{1}w_{2}$ and $w_{2}F(w_{1})\in S^{*}$ gives a reduced expression for $w^{\prime}$.
The assumption $F(S)=S$ only excludes the cases when $G$ is of type ${}^{2}B_{2}$, ${}^{2}F_{4}$, or ${}^{2}G_{2}$, which will be treated separately in Section 5.2.
Proposition 4.9.
Assume $F(S)=S$. Let $w=w_{1}w_{2}$ and $w^{\prime}=w_{2}F(w_{1})$ for some $w_{1},w_{2}\in S^{*}$. For all $k\geq 0$, we have $G^{F}$-equivariant isomorphisms
$H^{k}\Big{(}{\overline{X}^{S^{*}}(w)},\mathcal{O}_{{\overline{X}^{S^{*}}(w)}}\Big{)}\overset{\sim}{\longrightarrow}H^{k}\Big{(}{\overline{X}^{S^{*}}(w^{\prime})},\mathcal{O}_{{\overline{X}^{S^{*}}(w^{\prime})}}\Big{)}$
and
$H^{k}\Big{(}{\overline{X}^{S^{*}}(w)},\omega_{{\overline{X}^{S^{*}}(w)}}\Big{)}\overset{\sim}{\longrightarrow}H^{k}\Big{(}{\overline{X}^{S^{*}}(w^{\prime})},\omega_{{\overline{X}^{S^{*}}(w^{\prime})}}\Big{)}.$
Proof.
This follows from 4.6. We have $w^{\prime}=F_{1}^{{{\ell^{S^{*}}}(w_{1})}}(w)$ and the result follows from composing the corresponding isomorphisms on cohomology ${{\ell^{S^{*}}}(w_{1})}$-times. ∎
Corollary 4.10.
Assume $F(S)=S$. Let $w\in S^{*}$ such that $w=w_{1}ssw_{2}$ for some $s\in S,w_{1},w_{2}\in S^{*}$. There are $G^{F}$-equivariant isomorphisms for all $k\geq 0$,
$H^{k}\Big{(}{\overline{X}^{S^{*}}(w)},\mathcal{O}_{{\overline{X}^{S^{*}}(w)}}\Big{)}\overset{\sim}{\longrightarrow}H^{k}\Big{(}{\overline{X}^{S^{*}}(sw_{2}F(w_{1})F(s))},\mathcal{O}_{{\overline{X}^{S^{*}}(sw_{2}F(w_{1})F(s))}}\Big{)},$ (4.3)
and
$H^{k}\Big{(}{\overline{X}^{S^{*}}(w)},\omega_{{\overline{X}^{S^{*}}(w)}}\Big{)}\overset{\sim}{\longrightarrow}H^{k}\Big{(}{\overline{X}^{S^{*}}(sw_{2}F(w_{1})F(s))},\omega_{{\overline{X}^{S^{*}}(sw_{2}F(w_{1})F(s))}}\Big{)}.$ (4.4)
Moreover,
$H^{k}\big{(}{\overline{X}^{S^{*}}(w)},\mathcal{F}\big{)}\overset{\sim}{\longrightarrow}H^{k}\big{(}{\overline{X}^{S^{*}}(w_{2}F(w_{1})F(s))},\mathcal{F}\big{)}\overset{\sim}{\longrightarrow}H^{k}\big{(}{\overline{X}^{S^{*}}(sw_{2}F(w_{1}))},\mathcal{F}\big{)},$ (4.5)
for all $k\geq 0$ and $\mathcal{F}$ being the corresponding structure sheaf and canonical sheaf.
Proof.
As $sw_{2}F(w_{1})F(s)=F_{1}^{{{\ell^{S^{*}}}(w_{1}s)}}(w)$, the isomorphisms (4.3) and (4.4) follow from 4.9, and the isomorphisms (4.5) follow from composing (4.3) (resp. (4.4)) with the isomorphisms in 4.6. ∎
Example 4.11.
In the case when $w=ss$, we have the following isomorphisms on cohomology for for all $k\geq 0$ and for $\mathcal{F}$ being the structure sheaf (resp. canonical sheaf)
$H^{k}\big{(}{\overline{X}^{S^{*}}(ss)},\mathcal{F}\big{)}\overset{\sim}{\longrightarrow}H^{k}\big{(}{\overline{X}^{S^{*}}(sF(s))},\mathcal{F}\big{)}\overset{\sim}{\longrightarrow}H^{k}\big{(}{\overline{X}^{S^{*}}(F(s))},\mathcal{F}\big{)}\overset{\sim}{\longrightarrow}H^{k}\big{(}{\overline{X}^{S^{*}}(s)},\mathcal{F}\big{)}.$
5. Cohomology of compactified Deligne–Lustig varieties
5.1. Main theorem and applications
Theorem 5.1.
Assume $F(S)=S$. For each $w\in S^{*}$, there exists an $F$-conjugacy class $C$ in $W$ such that for any $x\in C_{\rm min}$ and any reduced expression of $x$, there are $G^{F}$-equivariant isomorphisms
$H^{k}\Big{(}{\overline{X}^{S^{*}}(w)},\mathcal{O}_{{\overline{X}^{S^{*}}(w)}}\Big{)}\overset{\sim}{\longrightarrow}H^{k}\Big{(}\overline{X}(x),\mathcal{O}_{\overline{X}(x)}\Big{)}$
and
$H^{k}\Big{(}{\overline{X}^{S^{*}}(w)},\omega_{{\overline{X}^{S^{*}}(w)}}\Big{)}\overset{\sim}{\longrightarrow}H^{k}\Big{(}\overline{X}(x),\omega_{\overline{X}(x)}\Big{)},$
[…]
$H^{k}\Big{(}\overline{X}(x),\mathcal{F}\Big{)}\overset{\sim}{\longrightarrow}{\rm Ind}_{P_{I}(\mathbb{F}_{q})}^{G(\mathbb{F}_{q})}H^{k}\Big{(}\overline{X}_{L_{I}}(x),\left.\mathcal{F}\right|_{\overline{X}_{L_{I}}(x)}\Big{)},$
where $\mathcal{F}$ is the structure sheaf (resp. canonical sheaf) of ${\overline{X}(x)}$.
Proof.
We give the proof for the structure sheaf. The case for the canonical sheaf follows from Serre duality for Deligne–Mumford stacks [Nir, Corollary 2.10].
By 3.7, we may assume ${\overline{X}^{S^{*}}(w)}$ has the same cohomology as an element in $B^{+}$ (3.5). In other words, let $u^{\prime}\in B^{+}$ be the image of $w$ under $S^{*}\twoheadrightarrow B^{+}$ and let $u\in S^{*}$ be a lift of $u^{\prime}$ with the same expression as $u^{\prime}$, then we have
$H^{k}\Big{(}{\overline{X}^{S^{*}}(u)},\mathcal{O}_{{\overline{X}^{S^{*}}(u)}}\Big{)}\overset{\sim}{\longrightarrow}H^{k}\Big{(}{\overline{X}^{S^{*}}(w)},\mathcal{O}_{{\overline{X}^{S^{*}}(w)}}\Big{)}$
for all $k\geq 0$.
Using the description of the submonoid of reduced elements in $B^{+}$ [GP, Remark 4.1.1, Exercise 4.1], we know that the criterion for $u$ to come from a reduced expression in the Weyl group is that $u$ has no subexpression of the form $ss$ for all $s\in S$. It follows from Corollary 4.10 that there exists $v\in S^{*}$ such that $v$ has no subexpression of the form $ss$ for all $s\in S$ and for all $k\geq 0$,
$H^{k}\Big{(}{\overline{X}^{S^{*}}(v)},\mathcal{O}_{{\overline{X}^{S^{*}}(v)}}\Big{)}\overset{\sim}{\longrightarrow}H^{k}\Big{(}{\overline{X}^{S^{*}}(w)},\mathcal{O}_{{\overline{X}^{S^{*}}(w)}}\Big{)}.$
In other words, we may assume $v\in S^{*}$ gives a reduced expression of $\alpha(v)$ in $W$. By [GKP, Theorem 2.6], there exists $v^{\prime}\in W$ with $\alpha(v)\to v^{\prime}$ and $\ell(v^{\prime})\leq\ell(\alpha(v))$. If $\alpha(v)$ is already the $F$-minimal length element in its conjugacy class, then there is nothing to prove.
In general, $\alpha(v)\to v^{\prime}$ gives $v_{1}^{\prime},\ldots,v_{m}^{\prime}\in W,s_{1},\ldots s_{m-1}\in S$ such that $v_{1}^{\prime}=\alpha(v),v_{m}^{\prime}=v^{\prime}$ and $v_{i}^{\prime}=s_{i-1}v_{i-1}^{\prime}F(s_{i-1})$ for all $i=2,\ldots,m$. Let $j\in\{1,\ldots,m\}$ be the smallest integer such that $\ell(v_{j})<\ell(v_{j-1})$. For all $i=1,\ldots j$, let $v_{i}\in S^{*}$ be the canonical lift of the corresponding reduced expression of $v_{i}^{\prime}$. It follows from 4.9 that ${\overline{X}^{S^{*}}(v_{i})}$ all have the same cohomology for the structure sheaf as ${\overline{X}^{S^{*}}(v)}$ for $i=2,\ldots,j-1$. On the other hand, we have $v_{j-1}=s_{j-1}v_{j}F(s_{j-1})$ in $S^{*}$. Via 4.6 we know that ${\overline{X}^{S^{*}}(v)}$ has the same cohomology as ${\overline{X}^{S^{*}}(s_{j-1}v_{j})}$ and ${\overline{X}^{S^{*}}(v_{j}F(s_{j-1}))}$. In this case, we reset $v$ to be $s_{j-1}v_{j}$. This process will only terminate and yields a minimal length element $x$ cf. loc. cit..
Any two minimal length elements $x$, $x^{\prime}$ in an $F$-conjugacy class $C$ are conjugate by $F$-cyclic shifts by [GKP, Theorem 2.6] and the corresponding cohomology of ${\overline{X}^{S^{*}}(x)}$ and ${\overline{X}^{S^{*}}(x^{\prime})}$ for the structure sheaf (resp. canonical sheaf) are isomorphic by 4.9.
For the last assertion, there are $G^{F}$-equivariant morphisms (cf. [DMR, Proposition 2.3.8])
${\overline{X}^{S^{*}}(x)}\overset{\sim}{\longrightarrow}G^{F}/U_{I}^{F}\times^{L_{I}^{F}}{\overline{X}^{S^{*}}_{L_{I}}(x)}\overset{\pi_{I}}{\longrightarrow}G^{F}/P_{I}^{F},$
where $P_{I}=\bigcup_{v\in W_{I}}{B}v{B}$, $U_{I}$ is the unipotent radical of $P_{I}$, and $L_{I}$ is the Levi subgroup of $G$ satisfying $P_{I}=L_{I}\ltimes U_{I}$. Since $G^{F}$ acts on the fibres of $\pi_{I}$ transitively, we have an $G^{F}$-equivariant isomorphisms
$H^{k}\Big{(}{\overline{X}^{S^{*}}(x)},\mathcal{O}_{{\overline{X}^{S^{*}}(x)}}\Big{)}\overset{\sim}{\longrightarrow}{\rm Ind}_{P_{I}(\mathbb{F}_{q})}^{G(\mathbb{F}_{q})}H^{k}\Big{(}\overline{X}^{S^{*}}_{L_{I}}(x),\mathcal{O}_{\overline{X}^{S^{*}}_{L_{I}}(x)}\Big{)}.$
for all $k\geq 0$. ∎
Remark 5.2.
With the same notations as above, it follows from the proof of the theorem that ${\rm supp}_{F}(x)={\rm supp}_{F}(w)$ for all $x\in C_{\rm min}$. Note that $x$ is of full $F$-support in $L_{I}$, so ${\overline{X}(x)}$ is irreducible by 2.12.
Corollary 5.3.
Assume $F(S)=S$. Let $v_{1},\ldots,v_{r}\in W$. There exists an $F$-conjugacy class $C$ in $W$ such that for any $x\in C_{\rm min}$ and any reduced expression of $x$, there are $G^{F}$-equivariant isomorphisms
$H^{k}\Big{(}X(\underline{v_{1}},\ldots,\underline{v_{r}}),\mathcal{O}_{X(\underline{v_{1}},\ldots,\underline{v_{r}})}\Big{)}\overset{\sim}{\longrightarrow}H^{k}\Big{(}\overline{X}(x),\mathcal{O}_{\overline{X}(x)}\Big{)}$
and
$H^{k}\Big{(}X(\underline{v_{1}},\ldots,\underline{v_{r}}),\omega_{X(\underline{v_{1}},\ldots,\underline{v_{r}})}\Big{)}\overset{\sim}{\longrightarrow}H^{k}\Big{(}\overline{X}(x),\omega_{\overline{X}(x)}\Big{)},$
[…]
$H^{k}\Big{(}\overline{X}(x),\mathcal{F}\Big{)}\overset{\sim}{\longrightarrow}{\rm Ind}_{P_{I}(\mathbb{F}_{q})}^{G(\mathbb{F}_{q})}H^{k}\Big{(}\overline{X}_{L_{I}}(x),\left.\mathcal{F}\right|_{\overline{X}_{L_{I}}(x)}\Big{)},$
where $\mathcal{F}$ is the structure sheaf (resp. canonical sheaf) of $\overline{X}(x)$.
Proof.
Fix reduced expressions for $v_{1},\ldots,v_{r}\in W$, and take lifts $v_{1}^{\prime},\ldots,v_{r}^{\prime}\in S^{*}$ corresponding to these reduced expressions. By 5.1, it suffices to show that the cohomology of $X(\underline{v_{1}},\ldots,\underline{v_{r}})$ for the structure sheaf is given by the cohomology of ${\overline{X}^{S^{*}}(v_{1}^{\prime}\cdots v_{r}^{\prime})}$.
Via projection, there is a proper birational morphism $f:{\overline{X}^{S^{*}}(v_{1}^{\prime}\cdots v_{r}^{\prime})}\to X(\underline{v_{1}},\ldots,\underline{v_{r}})$ of projective and strongly $F$-regular schemes (2.13). Thus there is a $G^{F}$-equivariant isomorphism
$\phi_{*}\mathcal{O}_{{\overline{X}^{S^{*}}(v_{1}^{\prime}\cdots v_{r}^{\prime})}}\overset{\sim}{\longrightarrow}\mathcal{O}_{X(\underline{v_{1}},\ldots,\underline{v_{r}})},$
by the proof of Zariski’s main theorem in [Har, Corollary 11.4].
The definition of local rings being strongly $F$-regular implies that they are $F$-rational. Smith has shown that $F$-rational implies pseudo-rational [Smi, Theorem 3.1]. Moreover, recall that compactified Deligne–Lusztig varieties are normal and Cohen–Macaulay (2.13). By [Kov, Theorem 8.13], the higher direct image $R^{i}\phi_{*}\mathcal{O}_{{\overline{X}^{S^{*}}(v_{1}^{\prime}\cdots v_{r}^{\prime})}}$ for $i>0$ vanish.
Functoriality of the associated Leray spectral sequence yields $G^{F}$-equivariant isomorphisms on cohomology groups
$H^{k}\Big{(}X(\underline{v_{1}},\ldots,\underline{v_{r}}),\mathcal{O}_{X(\underline{v_{1}},\ldots,\underline{v_{r}})}\Big{)}\overset{\sim}{\longrightarrow}H^{k}\Big{(}{\overline{X}^{S^{*}}(v_{1}^{\prime}\cdots v_{r}^{\prime})},\mathcal{O}_{{\overline{X}^{S^{*}}(v_{1}^{\prime}\cdots v_{r}^{\prime})}}\Big{)},$
for all $k\geq 0$. The $G^{F}$-equivariant isomorphisms for the cohomology of the canonical sheaf follow from Serre duality for Deligne–Mumford stacks [Nir, Corollary 2.10].
The following proposition shows that taking the image of a compactified Deligne–Lusztig variety in a partial flag variety does not affect its cohomology for the structure sheaf and canonical sheaf. In particular, Corollary 5.3 applies.
Proposition 5.4.
Assume $F(S)=S$. Let $P$ be a parabolic subgroup of $G$ containing ${B}$ and denote $\pi_{P}:G/B\to G/P$ the projection map. Then there are $G^{F}$-equivariant isomorphisms
$H^{k}\Big{(}X(\underline{v}),\mathcal{F}_{X(\underline{v})}\Big{)}\overset{\sim}{\longrightarrow}H^{k}\Big{(}\pi_{P}(X(\underline{v})),\mathcal{F}_{\pi_{P}(X(\underline{v}))}\Big{)},$
for all $k\geq 0$, where $\mathcal{F}$ is the structure sheaf (resp. canonical sheaf).
Proof.
The map $\pi_{P}:G/B\to G/P$ is Zariski locally trivial with fibres isomorphic to $P/{B}\cong L/(L\cap{B})$, where $L$ is the Levi subgroup of $P$. The induced map $\left.\pi_{P}\right|_{X(\underline{v})}:X(\underline{v})\to\pi_{P}(X(\underline{v}))$ is proper and Zariski locally trivial with fibres isomorphic to $L/(L\cap{B})$. The Levi subgroup $L$ is a reductive group scheme over $k$ with finitely many connected components. Without loss of generality, we assume that $L$ is connected. Moreover, $L/(L\cap{B})$ is a flag variety over $k$, so its cohomology for the structure sheaf is given by Kempf’s vanishing theorem: $H^{j}(L/(L\cap{B}),\mathcal{O}_{L/(L\cap{B})})=0$ for $j>0$ and $H^{0}(L/(L\cap{B}),\mathcal{O}_{L/(L\cap{B})})=k$. Hence the proof of 4.6 applies. ∎
5.2. Suzuki and Ree groups
In the classification of finite reductive groups, the Frobenius endomorphism does not necessarily fix the length of elements of $S$ for groups of type ${}^{2}B_{2}$, ${}^{2}F_{4}$, and ${}^{2}G_{2}$.
Analogous to [DL, §11], we work with groups of type ${}^{2}B_{2}$, ${}^{2}F_{4}$ (resp. ${}^{2}G_{2}$) assuming $k$ is of characteristic $2$ (resp. characteristic $3$). In this situation, let $q$ be an odd power of $\sqrt{2}$ (resp. $\sqrt{3}$) and let $G$ be defined over $\mathbb{F}_{q^{2}}$. Let $F^{\prime}:G\to G$ be a Frobenius root (cf. [DM, §4.2]) such that $F^{\prime 2}$ is the the $q^{2}$-Frobenius on $G$. In order to make $F^{\prime}$ fix the length of the simple roots, fix a $F^{\prime}$-stable maximal torus $T^{*}\subseteq G$ and a Borel ${B}$ containing $T^{*}$.
A Frobenius root shares all the properties of the Frobenius morphism we used in our proofs. In particular, it is étale and its corresponding Lang map is surjective. A systematic referece for Frobenius roots can be found in loc. cit. Therefore, once we fix the datum above, 5.1 and Corollary 5.3 readily yield $G^{F^{\prime}}$-equivariant isomorphism of cohomology groups of compactified Deligne–Lustig varieties. Note that the group $G^{F^{\prime}}$ for type ${}^{2}B_{2}$ is known as the Suzuki group and $G^{F^{\prime}}$ is known as the Ree groups for type ${}^{2}F_{4}$ and ${}^{2}G_{2}$.

### v2

@@
 Abstract.
-For connected reductive groups, we prove that the cohomology of the structure sheaf and the canonical sheaf for compactified Deligne–Lusztig varieties associated to a tuple $(w_{1},\ldots,w_{r})$ of Weyl group elements depends only on the $F$-support of $(w_{1},\ldots,w_{r})$.
+For connected reductive groups together with a Frobenius root $F$ that preserves the set of simple reflections, we show that for every tuple $(w_{1},\ldots,w_{r})$ of Weyl group elements there exists a minimal length element in an $F$-conjugacy class in $W$ such that the cohomology of the structure sheaf and the canonical sheaf for the associated compactified Deligne–Lusztig varieties are isomorphic respectively.
 Key words and phrases:
@@
 1.1. Background
-Let $q$ be a power of a prime $p$ and let $k$ be an algebraic closure of the finite field $\mathbb{F}_{q}$.
-Let $G_{0}$ be a connected reductive $\mathbb{F}_{q}$-group and $G$ its base change to $k$.
-Denote by $F:G\to G$ the corresponding $q$-Frobenius.
-The group of fixed points $G^{F}:=G(\mathbb{F}_{q})$ of $F$ is also known as a finite group of Lie type in the literature.
-Let $T_{0}\subset B_{0}$ be a maximal torus and a Borel subgroup of $G_{0}$ and $T\subset B$ their base change to $k$.
+Let $q$ be a power of a prime $p$ and let $k_{0}=\mathbb{F}_{q^{d}}$ for a positive integer $d$.
+Let $k$ be an algebraic closure of $\mathbb{F}_{p}$ containing $k_{0}$.
+Let $G_{0}$ be a connected reductive $k_{0}$-group with $q^{d}$-Frobenius $F_{0}:G_{0}\to G_{0}$.
+Let $G$ be the base change of $G_{0}$ to $k$ and let $F:G\to G$ be an endomorphism such that $F^{d}=F_{0}\otimes_{k_{0}}{\rm Id}$.
+The group of fixed points $G^{F}$ is known as a finite group of Lie type.
+Let $T\subset B$ be an $F$-stable maximal torus and an $F$-stable Borel subgroup of $G$.
 Denote by $W:=N(T)/T$ the corresponding Weyl group and let $S\subseteq W$ be the set of simple reflections.
-For any $w\in W$, the associated Deligne–Lusztig variety is denoted by $X(w)$, and there is a smooth compactification $\overline{X}(w)$ associated to each reduced expression of $w$.
-In [DL], Deligne and Lusztig considered the $\ell$-adic cohomology of $X(w)$ and Galois coverings of $X(w)$ to study the irreducible $G^{F}$-representations over fields of characteristic zero.
-The de Rham and mod $p$ cohomology of $X(w)$ and $\overline{X}(w)$ are less understood than their $\ell$-adic cohomology.
-However, when $G={\rm GL}_{2}$ and $w$ is the Coxeter element, the de Rham cohomology of $\overline{X}(w)$ is not mysterious because $\overline{X}(w)$ is isomorphic to the projective line.
-The simplicity in the situation of $\overline{X}(w)$ for $G={\rm GL}_{2}$ does not carry over to higher dimensions.
+For any $w\in W$, the associated Deligne–Lusztig variety is denoted by $X(w)$.
+Throughout the introduction, we fix a reduced expression of $w$ and denote by $\overline{X}(w)$ the corresponding smooth compactification of $X(w)$.
+In [DL], Deligne and Lusztig considered the $\ell$-adic cohomology of $X(w)$ and certain Galois coverings of $X(w)$ to study the irreducible $G^{F}$-representations over fields of characteristic zero.
+The de Rham and mod $p$ cohomology of $X(w)$ and $\overline{X}(w)$ are less understood than their $\ell$-adic counterpart.
+However, for $G^{F}={\rm GL}_{2}(\mathbb{F}_{q})$ and $w$ the Coxeter element, the de Rham cohomology of $\overline{X}(w)$ is not mysterious because $\overline{X}(w)$ is isomorphic to the projective line.
+The simplicity in the situation of $\overline{X}(w)$ for ${\rm GL}_{2}(\mathbb{F}_{q})$ does not carry over to higher dimensions.
 The smooth compactification $\overline{X}(w)$ is constructed in loc.
 cit.
-as the transversal intersection between the graph of the Frobenius and the desingularisation of the Schubert variety associated to $w$ constructed by Demazure and Hansen.
-Unlike the desingularisation of Schubert varieties, $\overline{X}(w)$ is not naturally an iterated $\mathbb{P}^{1}$-fibration.
+as the transversal intersection between the graph of $F$ and the desingularisation of the Schubert variety associated to $w$ constructed by Demazure and Hansen.
+Unlike these desingularisations of Schubert varieties, $\overline{X}(w)$ is not naturally an iterated $\mathbb{P}^{1}$-fibration.
 Thus the cohomology groups of $\overline{X}(w)$ cannot be computed in the same way.
-In [Wan], we described the cohomology of the structure sheaf of $\overline{X}(w)$ for any $w\in W$ in the case of $G={\rm GL}_{n}$.
+In [Wang22], we described the cohomology of the structure sheaf of $\overline{X}(w)$ for any $w\in W$ in the case of ${\rm GL}_{n}(\mathbb{F}_{q})$.
 Our main theorem showed that the cohomology of the structure sheaf of $\overline{X}(w)$ is always isomorphic to that of $\overline{X}(c)$, where $c$ is a Coxeter element in a parabolic subgroup of $W$ with ${\text{supp}}(c)={\text{supp}}(w):=\{s\in S\ |\ s\leq w\}$.
-As a consequence of our main theorem, we were able to describe the compactly supported integral $p$-adic cohomology of any Deligne–Lustig variety $X(w)$ for ${\rm GL}_{n}$.
+As a consequence of this theorem, we were able to describe the compactly supported integral $p$-adic cohomology of any Deligne–Lusztig variety $X(w)$ for ${\rm GL}_{n}(\mathbb{F}_{q})$.
 1.2. Main results
-The aim of this paper is to extend the main theorem of [Wan] to all connected reductive groups $G_{0}$.
-In this general setting, we encounter a special class of elements in $W$ which enjoy nice properties [GKP], [HN].
-Namely, they are the minimal length elements in the $F$-conjugacy classes of $W$ (cf.
-Section 4.2).
-In any given $F$-conjugacy class $C$, we denote by $C_{\rm min}$ the subset of minimal length elements.
-For example, if $G$ is ${\rm GL}_{n}$ or an inner form of ${\rm GL}_{n}$, such elements are precisely the Coxeter elements in parabolic subgroups of $W$.
-We show that the study of the cohomology of the structure sheaf and canonical sheaf of $\overline{X}(w)$ for any $w\in W$ can always be reduced to that of some minimal length element in an $F$-conjugacy class.
-Theorem 1.1.
-Let $w\in W$. There exists an $F$-conjugacy class $C$ in $W$ such that for any $x\in C_{\rm min}$, there are $G^{F}$-equivariant isomorphisms
-$H^{k}\Big{(}{\overline{X}(w)},\mathcal{O}_{{\overline{X}(w)}}\Big{)}\overset{\sim}{\longrightarrow}H^{k}\Big{(}{\overline{X}(x)},\mathcal{O}_{{\overline{X}(x)}}\Big{)}$
-and
-$H^{k}\Big{(}{\overline{X}(w)},\omega_{{\overline{X}(w)}}\Big{)}\overset{\sim}{\longrightarrow}H^{k}\Big{(}{\overline{X}(x)},\omega_{{\overline{X}(x)}}\Big{)}$
-for all $k\geq 0$. Moreover, there are $G^{F}$-equivariant isomorphisms
-$H^{k}\Big{(}{\overline{X}(x)},\mathcal{F}\Big{)}\overset{\sim}{\longrightarrow}{\rm Ind}_{P_{I}(\mathbb{F}_{q})}^{G(\mathbb{F}_{q})}H^{k}\Big{(}\overline{X}_{L_{I}}(x),\left.\mathcal{F}\right|_{\overline{X}_{L_{I}}(x)}\Big{)},$
-for all $k\geq 0$, where $I$ is the $F$-support of $x$, $P_{I}$ and $L_{I}$ are the associated standard parabolic and Levi subgroup, and $\mathcal{F}$ is the structure sheaf (resp.
-canonical sheaf) of ${\overline{X}(x)}$.
-One should note that the $F$-conjugacy class $C$ in the above theorem may not contain $w$.
-However, for any given $w$, one can determine explicitly the minimal length elements in this $F$-conjugacy class by following our recipe given in the proof of 5.1.
-Smooth compactifications of $X(w)$ given by different reduced expressions of $w$ are not necessarily isomorphic.
-To avoid this ambiguity, we work with Deligne–Lusztig varieties $X(w_{1},\ldots,w_{r})$ and compactified Deligne–Lusztig varieties $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ associated to a tuple of Weyl group elements cf.
-2.2.
-These notions also make the construction of partial compactifications of Deligne–Lusztig varieties more straightforward and uniform.
-One can always recover statements for classical Deligne–Lusztig varieties by setting $r=1$.
-In this generality, we give a connectedness criterion for compactified Deligne–Lusztig varieties $X(\underline{w_{1}},\ldots,\underline{w_{r}})$, generalising a result of Bonnafé and Rouquier [BR].
-We also show that they have $F$-rational singularities and thus are normal, Cohen–Macaulay, and pseudo-rational.
-Furthermore, we provide a list of partial compactifications $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ of Deligne–Lusztig varieties which are smooth.
+The aim of this paper is to extend the main theorem of [Wang22] to all pairs $(G,F)$ for which $F(S)=S$.
+The assumption on $F$, for example, is satisfied when $G$ is split over $k_{0}$, or when $G^{F}$ is the group of unitary matrices in ${\rm GL}_{n}(\mathbb{F}_{q^{2}})$.
+If $G^{F}$ is a finite simple group of Lie type arising from a pair $(G,F)$ as in in Section 1.1, then the classification of such groups implies that $F(S)=S$ is satisfied.
+In fact, $F$ would be the Frobenius endomorphism (i.e.
+$d=1$) except for the cases for Suzuki and Ree groups, for which $d=2$ and additional assumptions on $q$ are required (Section 5.2).
+In our setting, we encounter the minimal length elements in the $F$-conjugacy classes of $W$ (cf.
+Section 4.2), which is a special class of elements in $W$ with nice properties [GeckKimPfeifferMinlength, Theorem 2.6], [HeNieminlength, Theorem 3.1].
+For example, if $G^{F}$ is ${\rm GL}_{n}(\mathbb{F}_{q})$ or the group of unitary matrices in ${\rm GL}_{n}(\mathbb{F}_{q^{2}})$, then such elements are precisely the Coxeter elements in parabolic subgroups of $W$.
+We show that the study of the cohomology of the structure sheaf and the canonical sheaf of $\overline{X}(w)$ for any $w\in W$ can always be reduced to that of a minimal length element in an $F$-conjugacy class in $W$.
+Theorem 5.1.1.
+Assume $F(S)=S$ and let $w\in W$. There exists an $F$-conjugacy class $C$ in $W$ and a minimal length element $x_{0}\in C$ such that for all $i\geq 0$, there are $G^{F}$-equivariant isomorphisms
+$H^{i}\left({\overline{X}(w)},\mathcal{F}_{{\overline{X}(w)}}\right)\overset{\sim}{\longrightarrow}H^{i}\left({\overline{X}(x_{0})},\mathcal{F}_{{\overline{X}(x_{0})}}\right),$
+where $\mathcal{F}\in\{\mathcal{O},\omega\}$ is either the structure sheaf or the canonical sheaf.
+Note that the $F$-conjugacy class $C$ in the theorem may not contain $w$.
+However, for any given $w$, one can determine $x_{0}$ and $C$ explicitly by following our recipe given in the proof of Theorem 5.1.1.
+We will also show that the cohomology groups in the theorem do not depend on the choice of reduced expressions of $w$ and $x_{0}$ and their corresponding smooth compactifications.
+On the other hand, smooth compactifications of $X(w)$ given by different reduced expressions of $w$ are not necessarily isomorphic.
+To avoid this ambiguity, we work with compactified Deligne–Lusztig varieties $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ associated to a tuple of Weyl group elements, cf.
+Definition 2.1.1, [DMR, Definition 2.3.2].
+This notion makes the construction of compactifications of Deligne–Lusztig varieties more straightforward and uniform.
+For example, setting $r=1$ yields the Zariski closure $X(\underline{w_{1}})=\overline{X(w_{1})}$ of $X(w_{1})$ in $G/B$.
+Setting $w_{1},\ldots,w_{r}\in S$ such that $w_{1}\cdots w_{r}\in W$ is a reduced expression yields the smooth compactification $X(\underline{w_{1}},\ldots,\underline{w_{r}})=\overline{X}(w_{1}\cdots w_{r})$ defined by Deligne and Lusztig.
+In this generality, we give a connectedness criterion for compactified Deligne–Lusztig varieties $X(\underline{w_{1}},\ldots,\underline{w_{r}})$, generalising a result of Bonnafé and Rouquier [BonnafeRouquierOntheirred].
+We also show that they have $F_{p}$-rational singularities (Section 2.4) and thus are normal, Cohen–Macaulay, and pseudo-rational.
+Furthermore, we provide a list of compactifications $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ of Deligne–Lusztig varieties which are smooth.
 The structure of the paper is as follows.
 We fix our notations for Deligne–Lusztig varieties and compactified Deligne–Lusztig varieties and describe their geometric properties in Section 2.
-In Section 3, we prove that $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is smooth when each $w_{1},\ldots,w_{r}$ comes from a braid relation in the Weyl group $W$, and we establish relations between cohomology groups for the structure sheaf and canonical sheaf on the associated compactified Deligne–Lusztig varieties.
+In Section 3, we show that $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is smooth when each $w_{i},i=1,\ldots,r,$ comes from a braid relation in the Weyl group $W$, and we establish relations between cohomology groups for the structure sheaf and the canonical sheaf on the associated compactified Deligne–Lusztig varieties.
 Section 4 contains the constructions of morphisms which are $\mathbb{P}^{1}$-fibrations between certain compactified Deligne–Lusztig varieties.
 They are associated with $F$-cyclic conjugations in the Weyl group.
-In 4.6, we relate the cohomology of the structure sheaf (resp.
+In Proposition 4.1.4, we relate the cohomology of the structure sheaf (resp.
 canonical sheaf) between these compactified Deligne–Lusztig varieties.
-Furthermore, we deduce that the cohomology of the structure sheaf (resp.
-canonical sheaf) for $\overline{X}(w)$ and $\overline{X}(F_{1}^{m}(w))$ are isomorphic for all $m\geq 0$, where $F_{1}$ is a partial Frobenius morphism (2.5).
-The key observation in this section is that the partial Frobenius morphism $F_{1}$ induces an automorphism of the Weyl group, even though it has no well-defined inverse as a morphism of schemes.
-In Section 5, we begin with proving the statement in 5.1 under the assumption that $F(S)=S$.
-The remaining cases come from Suzuki and Ree groups, which are treated in Section 5.2.
-Finally, the theorem is proved in full generality in Corollary 5.3.
-Acknowledgements
-The author thanks Jochen Heinloth and Georg Linden for their comments on an earlier draft of this paper. Thanks to Ulrich Görtz for helpful conversations.
-Parts of the work are carried out when the author was a member of the trimester program "The Arithmetic of the Langlands Program" at the Hausdorff Research Institute for Mathematics in Bonn and when the author was a member of the Research Training Group 2553 in Essen, both funded by the DFG (Deutsche Forschungsgemeinschaft).
+Finally, we prove Theorem 5.1.1 and extend the result to arbitrary compactified Deligne–Lusztig varieties of the form $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ in Section 5.
 2. Compactified Deligne–Lusztig varieties
 We begin with fixing notations related to Deligne–Lusztig varieties associated to a tuple of Weyl group elements.
-By adapting classical proofs, we give criterions for compactified Deligne–Lusztig varieties to be connected and smooth, respectively.
+This construction has been treated in [DL, §9] and [DMR, §2].
+We review here the definitions and properties that are relevant for our context.
+Moreover, by adapting classical proofs, we give criterions for compactified Deligne–Lusztig varieties to be connected and show that they are $F_{p}$-rational.
 2.1. Definitions and basic properties
-Let $G$, $B$, $T$, $F:G\to G$, $W$, $S$ be as in Section 1.1. Denote by $\ell:W\to\mathbb{N}\cup\{0\}$ the Bruhat length and $\leq$ the Bruhat order on $W$.
+Let $G$, $F:G\to G$, $T\subset B$, $S\subset W$ be as in Section 1.1.
+In the literature, the endomorphism $F$ is referred to as a Frobenius root or a Steinberg endomorphism.
+Denote by $\ell:W\to\mathbb{N}\cup\{0\}$ the Bruhat length and $\leq$ the Bruhat order on $W$.
 Recall that by the Bruhat decomposition, the diagonal $G$-action on the product $G/B\times G/B$ yields a bijection between the Weyl group $W$ and $G\backslash\left(G/B\times G/B\right)$.
@@
 $O(w_{1},\ldots,w_{r}):=O(w_{1})\times_{G/B}\cdots\times_{G/B}O(w_{r}).$
-We use such tuple of elements $w_{1},\ldots,w_{r}\in W$ for the definition of partial compactifications.
-They will in particular be useful for avoiding ambiguity in having different reduced expressions of a Weyl group element.
-Definition 2.1.
-Let $w_{1},\ldots,w_{r}\in W$. We define the compactified orbit associated to $(w_{1},\ldots,w_{r})$ as the $k$-scheme
-$O(\underline{w_{1}},\ldots,\underline{w_{r}}):=\overline{O(w_{1})}\times_{G/B}\cdots\times_{G/B}\overline{O(w_{r})},$
-where for each $i=1,\ldots,r$, $\overline{O(w_{i})}$ is the Zariski closure of $O(w_{i})$ in $G/B\times G/B$.
-The compactified orbit $O(\underline{w_{1}},\ldots,\underline{w_{r}})$ can be viewed as an iterated fibre bundle over $G/B$ with fibres isomorphic to Schubert varieties $\overline{Bw_{i}B}/B$.
-Thus $O(\underline{w_{1}},\ldots,\underline{w_{r}})$ is reduced, separated, normal, Cohen Macaulay $k$-schemes locally of finite type.
+Moreover, we denote the fibre product of the Zariski closures $\overline{O(w_{i})}\subseteq G/B\times G/B$ of the orbits by
+$O(\underline{w_{1}},\ldots,\underline{w_{r}}):=\overline{O(w_{1})}\times_{G/B}\cdots\times_{G/B}\overline{O(w_{r})}.$
+We refer to them as the compactified orbit associated to $(w_{1},\ldots,w_{r})$ in the rest of the article.
 The algebraic group $G$ acts diagonally on $O(w_{1},\ldots,w_{r})$ and $O(\underline{w_{1}},\ldots,\underline{w_{r}})$.
-Whenever $(w_{1}^{\prime},\ldots,w_{r}^{\prime})$ is a tuple satisfying $w_{i}^{\prime}\leq w_{i}$ for all $i$, there is a $G$-equivariant embedding $O(w_{1}^{\prime},\ldots,w_{r}^{\prime})\hookrightarrow O(\underline{w_{1}},\ldots,\underline{w_{r}})$ cf.[DMR, Proposition 2.2.16].
-Moreover, there is a stratification
-$O(\underline{w_{1}},\ldots,\underline{w_{r}})=\bigcup_{w_{i}^{\prime}\leq w_{i},\text{ for all }i}O(w_{1}^{\prime},\ldots w_{r}^{\prime}).$
-Denote by $F:G/B\to G/B$ the map induced by the Frobenius $F:G\to G$ and $\Gamma_{F}\subseteq G/B\times G/B$ its graph.
-Definition 2.2.
+Since $O(\underline{w_{1}},\ldots,\underline{w_{r}})$ is an iterated fibre bundle over $G/B$ with fibres isomorphic to Schubert varieties $\overline{Bw_{i}B}/B$, cf.
+[Jantzen, F.23], it is an integral, separated, normal, Cohen–Macaulay $k$-scheme locally of finite type.
+A proof of normality can be found in [DMR, Lemma 2.2.13].
+The Cohen–Macaulay property behaves well under base change [stacks-project, Lemma 045T], so $O(\underline{w_{1}},\ldots,\underline{w_{r}})$ is Cohen–Macaulay follows from the classical result that Schubert varieties are Cohen–Macaulay (see the introduction of [Hashimoto]).
+Denote by $F:G/B\to G/B$ the map induced by the endomorphism $F:G\to G$ and $\Gamma_{F}\subseteq G/B\times G/B$ its graph.
+Since $F^{d}$ is a Frobenius endomorphism, we know that $F$ is surjective, finite ([DigneMichelBook, Proposition 8.1.2]), separated ([GoerWedhAlgGeom1, Proposition 9.13]), and purely inseparable ([EGA1, (3.7.6)]).
+Moreover, the induced map $F:W\to W$ is a group automorphism [DigneMichelBook, Lemma 4.2.21].
+Definition 2.1.1.
 Let $w_{1},\ldots,w_{r}\in W$. We define the Deligne–Lusztig variety associated to $(w_{1},\ldots,w_{r})$ as the fibre product
@@
 $X(\underline{w_{1}},\ldots,\underline{w_{r}})$).
+For any $w_{1},\ldots,w_{r}\in W$, the $k$-scheme $X(w_{1},\ldots,w_{r})$ is smooth of dimension $\ell(w_{1})+\cdots+\ell(w_{r})$.
+If for all $i=1,\ldots,r$, the associated Schubert variety $\overline{{B}w_{i}{B}}/{B}$ is smooth, then $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is smooth of dimension $\ell(w_{1})+\cdots+\ell(w_{r})$, cf.
+[DMR, Proposition 2.3.5, 2.3.6 (i)].
+In general, compactified Deligne–Lusztig varieties have $F_{p}$-rational singularities, as we will see in Proposition 2.4.1.
 Whenever $(w_{1}^{\prime},\ldots,w_{r}^{\prime})$ is a tuple satisfying $w_{i}^{\prime}\leq w_{i}$ for all $i$, there is a $G^{F}$-equivariant embedding $X(w_{1}^{\prime},\ldots,w_{r}^{\prime})\hookrightarrow X(\underline{w_{1}},\ldots,\underline{w_{r}})$.
@@
 When $s_{i}\in S$ for all $i=1,\ldots,r$, the tuple $(s_{1},\ldots,s_{r})$ can be regarded as an element in the free monoid generated by $S$, which we denote by $S^{*}$.
-By Lemma 2.3 below, $X(\underline{s_{1}},\ldots,\underline{s_{r}})$ is smooth.
-To simplify the notation, given any $w=s_{1}\cdots s_{m}\in S^{*}$ with $s_{i}\in S,i=1,\ldots m$, we denote $X(s_{1},\ldots,s_{m})$ by $X^{S^{*}}(w)$ and $X(\underline{s_{1}},\ldots,\underline{s_{m}})$ by $\overline{X}^{S^{*}}(w)$.
+For any $s\in S$, the Zariski closure $\overline{BsB}=B\langle s\rangle B$ is the parabolic subgroup corresponding to the subgroup $\langle s\rangle\subset W$, cf.
+[DMR, Corollary 2.2.10], so the compactified Deligne–Lusztig variety $X(\underline{s_{1}},\ldots,\underline{s_{r}})$ is always smooth.
+To simplify the notations, given any $w=s_{1}\cdots s_{m}\in S^{*}$ with $s_{i}\in S,i=1,\ldots m$, we denote $X(s_{1},\ldots,s_{m})$ by $X^{S^{*}}(w)$ and $X(\underline{s_{1}},\ldots,\underline{s_{m}})$ by $\overline{X}^{S^{*}}(w)$.
 The notations $O^{S^{*}}(w)$ and $\overline{O}^{S^{*}}(w)$ for the orbits are defined analogously.
@@
 Let $\preceq$ be the Bruhat order in $S^{*}$ and let $\ell^{S^{*}}:S^{*}\to\mathbb{N}\cup\{0\}$ be the Bruhat length.
-Lemma 2.3.
-Let $w_{1},\ldots,w_{r}\in W$.
-The Deligne–Lusztig variety $X(w_{1},\ldots,w_{r})$ is smooth of dimension $\ell(w_{1})+\cdots+\ell(w_{r})$.
-If for all $i=1,\ldots,r$, the associated Schubert variety $\overline{{B}w_{i}{B}}/{B}$ is smooth, then $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is smooth of dimension $\ell(w_{1})+\cdots+\ell(w_{r})$.
-Proof.
-In the proof of [DL, Lemma 9.11], Deligne and Lusztig showed that the graph $\Gamma_{F}$ of the Frobenius is transversal to any smooth $G$-equivariant scheme $Y\to G/B\times G/B$.
-Here let $Y=O(w_{1},\ldots,w_{r})$ and $Y=O(\underline{w_{1}},\ldots,\underline{w_{r}})$, which are $G$-equivariant schemes over $G/B\times G/B$.
-By construction, $O(w_{1},\ldots,w_{r})$ is an iterated fibration over $G/B$ with fibres being Schubert cells, which are affine spaces.
-The compactified orbit $O(\underline{w_{1}},\ldots,\underline{w_{r}})$ is an iterated fibration over $G/B$ with fibres being Schubert varieties $\overline{{B}w_{i}{B}}/{B}$.
-Hence the smoothness of $X(w_{1},\ldots,w_{r})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ follows.
-∎
-Remark 2.4.
-In general, compactified Deligne–Lusztig varieties are $F$-rational and thus pseudo-rational, normal, and Cohen–Macaulay. We will give a proof for this in 2.13.
 2.2. The Lang map
-There is an alternative description of Deligne–Lusztig varieties via the Lang map.
+We consider an alternative description of Deligne–Lusztig varieties via the Lang map.
 Using this description, we show that $G^{F}$ acts transitively on the set of connected components of $X(w_{1},\ldots,w_{r})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}})$.
@@
 This is again a $k$-group scheme.
-Definition 2.5.
-Let $F_{1}:G^{r}\to G^{r}$ be the partial Frobenius morphism given by
+We begin with fixing notations. Let $F_{1}:G^{r}\to G^{r}$ be the endomorphism given by
 $(g_{0},\ldots,g_{r-1})\longmapsto(g_{1},\ldots,g_{r-1},F(g_{0})).$
+Note that $F_{1}$ can be viewed as a Frobenius root for the algebraic group $G^{r}$.
 Let $\mathcal{L}_{1}:G^{r}\to G^{r}$ be the Lang map associated to $F_{1}$ given by $x\mapsto x^{-1}F_{1}(x)$ for $x\in G^{r}$. In other words, we have
 $\mathcal{L}_{1}:(g_{0},\ldots,g_{r-1})\longmapsto(g_{0}^{-1}g_{1},\ldots,g_{r-2}^{-1}g_{r-1},g_{r-1}^{-1}F(g_{0})).$
-Similarly, we define the partial Frobenius morphism on $\left(G/B\right)^{r}$.
-Since $B$ is the base change of $B_{0}$ to ${\rm Spec}\ k$, we know that $G/B$ is the base change of $G_{0}/B_{0}$ to ${\rm Spec}\ k$ and has a corresponding $q$-Frobenius map $F:G/B\to G/B$.
-Denote by $f_{2}:\left(G/B\right)^{r}\to\left(G/B\right)^{r}$ the map given by
-$(x_{0},\ldots,x_{r-2},x_{r-1})\longmapsto(x_{0},\ldots,x_{r-2},F(x_{r-1})),$
-which is the identity on the first $(r-1)$-entries and Frobenius on the last entry. Denote by $f_{1}:\left(G/B\right)^{r}\to\left(G/B\right)^{r}$ the map
-$(x_{0},\ldots,x_{r-1})\longmapsto(x_{1},\ldots,x_{r-1},x_{0}),$
-which shifts the coordinates by $1$. The composition $f_{1}^{r}$ is the identity.
-Definition 2.6.
-We define the partial Frobenius morphism $F_{1}:\left(G/B\right)^{r}\to\left(G/B\right)^{r}$ to be the composition $f_{2}\circ f_{1}$.
-The partial Frobenius morphism $F_{1}$ on $\left(G/B\right)^{r}$ is an étale morphism and the composition $F_{1}^{r}$ is the $q$-Frobenius morphism on $\left(G/B\right)^{r}$.
-Lemma 2.7.
-The Lang map $\mathcal{L}_{1}:G^{r}\to G^{r}$ is surjective and is a Galois cover with Galois group $G^{F}$.
-Proof.
-Note that $G$ is geometrically connected and smooth, so the product $G^{r}$ is connected and smooth.
-Observe that the composition of the partial Frobenius $F_{1}^{r}$ yields the Frobenius on $G^{r}$, so [CE, Theorem 7.1] applied to $G^{r}$ yields that $\mathcal{L}_{1}$ is surjective and étale.
-The kernel of $\mathcal{L}_{1}$ is $\Delta(G^{F})$, where $\Delta:G\to G^{r}$ is the diagonal embedding.
-Thus $\mathcal{L}_{1}$ is a finite morphism.
-Since the diagonal (left) action of $G^{F}$ is a free action and induces automorphisms of the cover, and the number of automorphisms is the same as the degree of $\mathcal{L}_{1}$, we know that $\mathcal{L}_{1}$ is a Galois covering with Galois group $G^{F}$ [SP, 03SF].
-∎
-Lemma 2.8.
+Lemma 2.2.1.
 For any $w_{1},\ldots,w_{r}\in W$, there are $G^{F}$-equivariant isomorphisms
@@
 $X(\underline{w_{1}},\ldots,\underline{w_{r}})\overset{\sim}{\longrightarrow}\mathcal{L}_{1}^{-1}\left(\overline{{B}w_{1}{B}}\times_{{\rm Spec}\ k}\cdots\times_{{\rm Spec}\ k}\overline{{B}w_{r}{B}}\right)/({B})^{r}.$
-Furthermore, $G^{F}$ acts transitively on the connected components of $X(w_{1},\ldots,w_{r})$ (resp.
+Furthermore, $G^{F}$ acts transitively on the irreducible components of $X(w_{1},\ldots,w_{r})$ (resp.
 $X(\underline{w_{1}},\ldots,\underline{w_{r}})$).
-In particular, $X(w_{1},\ldots,w_{r})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ are equi-dimensional.
+In particular, $X(w_{1},\ldots,w_{r})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ are equidimensional.
 Proof.
@@
 In this proof, we denote $\mathcal{B}:=\overline{{B}w_{1}{B}}\times_{{\rm Spec}\ k}\cdots\times_{{\rm Spec}\ k}\overline{{B}w_{r}{B}}$.
-Observe that $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is isomorphic to a closed subscheme of $\left(G/B\right)^{r}$ via the projection ${\rm pr}_{0,\ldots,r-1}:\left(G/B\right)^{r+1}\to\left(G/B\right)^{r}$ to the first $r$-entries.
-The diagonal right $({B})^{r}$-action on $G^{r}$ is free.
-The subscheme $\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)$ of $G^{r}$ is stable under the right $({B})^{r}$-action, so we may take the quotient of $\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)$ with respect to the induced free right $({B})^{r}$-action.
+By Lang’s theorem as formulated in [CabanEngue, Theorem 7.1], we deduce that $\mathcal{L}_{1}:G^{r}\to G^{r}$ is surjective and is a Galois cover with Galois group $G^{F}$.
+The subscheme $\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)$ of $G^{r}$ is stable under the free $({B})^{r}$-action from the right, so we may take the quotient of $\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)$ with respect to the induced free right $({B})^{r}$-action.
 It follows that $\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$ is a closed subscheme of $\left(G/B\right)^{r}$.
-The isomorphism above then follows from verifying the definitions cf.
-[DL].
-With the diagonal $G^{F}$-action on $\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$, the isomorphism is $G^{F}$-equivariant.
-We show the second assertion for $\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$.
+Observe that $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is also isomorphic to a closed subscheme of $\left(G/B\right)^{r}$ via projection to the first $r$-entries.
+The isomorphisms in this lemma then follows from verifying the definitions, cf.
+[DL, §1.11].
+Moreover,
+$G^{F}\backslash(\mathcal{L}_{1}^{-1}(\mathcal{B}))\overset{\sim}{\longrightarrow}\mathcal{B}$
+is irreducible.
 Since the quotient map $\pi:\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)\to\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$ is equivariant under the left $G^{F}$-action and surjective, it induces a surjective morphism $G^{F}\backslash(\mathcal{L}_{1}^{-1}(\mathcal{B}))\to G^{F}\backslash\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$.
-Via Lemma 2.7 we know that
-$G^{F}\backslash(\mathcal{L}_{1}^{-1}(\mathcal{B}))\overset{\sim}{\longrightarrow}\mathcal{B}$
-is irreducible, so $G^{F}\backslash\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$ is irreducible and $G^{F}$ acts transitively on the set of connected components of $\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$.
-Therefore $\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$ is equidimensional.
-∎
-Remark 2.9.
+Thus $G^{F}\backslash\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$ is irreducible and $G^{F}$ acts transitively on the set of irreducible components of $\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$.
+Remark 2.2.2.
 The $G^{F}$-action on $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is not free.
@@
 2.3. Criterion for connectedness
-There is a simple criterion for classical Deligne–Lusztig varieties to be connected, proved by Lusztig, Bonnafé–Rouquier [BR], Görtz [Gör], and many others.
+There is a simple criterion for classical Deligne–Lusztig varieties to be connected, proved by Lusztig, Bonnafé–Rouquier [BonnafeRouquierOntheirred], Görtz [GoertzOntheconn], and many others.
 In this section we prove that $X(w_{1},\ldots,w_{r})$ (resp.
-$X(\underline{w_{1}},\ldots,\underline{w_{r}})$) is connected when the tuple $(w_{1},\ldots,w_{r})$ has full $F$-support by modifying the proof of [BR].
+$X(\underline{w_{1}},\ldots,\underline{w_{r}})$) is connected when the tuple $(w_{1},\ldots,w_{r})$ has full $F$-support by modifying the proof of [BonnafeRouquierOntheirred].
 To begin, we fix the notations for the set of $F$-orbits in $S$ and the $F$-support for elements in $W$ and $S^{*}$.
-Definition 2.10.
-The $F$-orbit of $w\in W$ is the set ${{\rm Orb}_{F}(w)}:=\{F^{k}(w)|k\in\mathbb{Z}_{\geq 0}\}$. Under the assumption that $F(S)=S$, we define the set
+Definition 2.3.1.
+The $F$-orbit of $w\in W$ is the set ${{\rm Orb}_{F}(w)}:=\{F^{n}(w)|n\in\mathbb{Z}_{\geq 0}\}$. Under the assumption that $F(S)=S$, we define the set
 $S_{F}:=\{{{\rm Orb}_{F}(s)}|s\in S\}$
@@
 ${{\text{supp}}}_{F}(w):=\{\tilde{s}\in S_{F}|\text{there exists }t\in\tilde{s}\text{ such that }t\leq w\}.$
-An element $w\in W$ is called of full $F$-support if ${{\text{supp}}}_{F}(w)=S_{F}$. For a tuple $(w_{1},\ldots w_{r})\in W^{r}$, we denote
+An element $w\in W$ is called of full $F$-support if ${{\text{supp}}}_{F}(w)=S_{F}$.
+For a tuple $(w_{1},\ldots w_{r})\in W^{r}$, we denote
 ${{\text{supp}}}_{F}(w_{1},\ldots,w_{r}):={{\text{supp}}}_{F}(w_{1})\cup\cdots\cup{{\text{supp}}}_{F}(w_{r}).$
@@
 ${{\text{supp}}}_{F}(v):={{\text{supp}}}_{F}(s_{1},\ldots,s_{m}).$
-Remark 2.11.
+Remark 2.3.2.
 We say that $w\in W$ is a Coxeter element when $w=s_{1}\cdots s_{r}$ is of full $F$-support and each $s_{i}\in S$ is from a distinct $F$-orbit.
-Proposition 2.12.
-[…]
+Proposition 2.3.3.
+Let $w_{1},\ldots,w_{r}\in W$ and denote $I={{\text{supp}}}_{F}(w_{1},\ldots,w_{r})$.
+The Deligne–Lusztig varieties $X(w_{1},\ldots,w_{r})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ are irreducible if and only if $I=S_{F}$.
+In general, they have $|G^{F}/P_{I}^{F}|$ connected components, where $W_{I}$ is the $F$-stable parabolic subgroup of $W$ generated by $\{s\in W|{{\rm Orb}_{F}(s)}\in I\}$ and $P_{I}={B}W_{I}{B}$.
 Proof.
-Take the product $w_{1}\cdots w_{r}$ in $S^{*}$ and denote it by $v$.
+For certain reduced expressions of $w_{1},\ldots,w_{r}$, take the product $w_{1}\cdots w_{r}$ in $S^{*}$ and denote it by $v$.
 This product depends on the choices of the reduced expressions of $w_{1},\ldots,w_{r}$, but for each such $v$, there is an isomorphism
 $X^{S^{*}}(v)\overset{\sim}{\longrightarrow}X(w_{1},\ldots,w_{r})$
-via projection.
+via projection, cf.
+[DMR, §2.3.1].
 It also follows that $X(w_{1},\ldots,w_{r})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ are dense open subschemes of ${\overline{X}^{S^{*}}(v)}$.
-Thus it suffices to prove connectedness for ${\overline{X}^{S^{*}}(v)}$.
-The proof showing that ${\overline{X}^{S^{*}}(v)}$ is connected is analogous to the proof of [BR] for classical Deligne–Lusztig varieties.
-Suppose ${{\text{supp}}}_{F}(w_{1},\ldots,w_{r})=S_{F}$ and thus ${{\text{supp}}}_{F}(v)=S_{F}$.
+Thus it suffices to prove that ${\overline{X}^{S^{*}}(v)}$ is connected if and only if $v$ has full $F$-support.
+The proof showing that ${\overline{X}^{S^{*}}(v)}$ is connected is analogous to the proof of [BonnafeRouquierOntheirred] for classical Deligne–Lusztig varieties.
+Suppose $I=S_{F}$ and thus ${{\text{supp}}}_{F}(v)=S_{F}$.
 Let $Z$ be the connected component of ${\overline{X}^{S^{*}}(v)}$ which contains the point $({B},\ldots,{B})$.
@@
 Fix $v^{\prime}$ to be of minimal length such that $v^{\prime}\preceq v$ and ${{\text{supp}}}_{F}(v^{\prime})=S_{F}$.
-In particular, $v^{\prime}$ has the same expression as $\alpha(v^{\prime})\in W$, which is a Coxeter element.
-This implies that ${\overline{X}^{S^{*}}(v^{\prime})}$ is irreducible [Lus], [BR].
+In particular, $v^{\prime}$ is a reduced expression of $\alpha(v^{\prime})\in W$, which is a Coxeter element.
+This implies that ${\overline{X}^{S^{*}}(v^{\prime})}$ is irreducible [LusztigCoxeterOrbits, Proposition 4.8].
 On the other hand, there is a $G^{F}$-equivariant embedding ${\overline{X}^{S^{*}}(v^{\prime})}\to{\overline{X}^{S^{*}}(v)}$.
@@
 Thus ${\overline{X}^{S^{*}}(v)}$ is irreducible.
+In this case, its dense open subschemes $X(w_{1},\ldots,w_{r})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ are also irreducible [GoerWedhAlgGeom1, Proposition 3.24].
 Suppose ${{\text{supp}}}_{F}(w_{1},\ldots,w_{r})=I$ and so ${{\text{supp}}}_{F}(v)=I$. By [DMR, Proposition 2.3.8], we know that there is a $G^{F}$-equivariant isomorphism
@@
 ∎
-2.4. Criterion for smoothness
-In [Wan, Proposition A.6], we showed that some compactified Deligne–Lusztig varieties are strongly $F$-regular. This property generalises to arbitrary compactified Deligne–Lusztig varieties.
-Proposition 2.13.
-For $w_{1},\ldots,w_{r}\in W$, the compactified Deligne–Lusztig variety $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is strongly $F$-regular.
-In particular, $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is $F$-rational and thus normal and Cohen-Macaulay.
+2.4. $F_{p}$-regularity
+We denote by $F_{p}$ the (absolute) $p$-Frobenius to distinguish our notations from the notations coming from the theory of $F_{p}$-singularities in commutative algebra and algebraic geometry.
+In [Wang22, Proposition A.6], we showed that some compactified Deligne–Lusztig varieties are strongly $F_{p}$-regular.
+This property generalises to arbitrary compactified Deligne–Lusztig varieties.
+As in loc.
+cit., we use a similar strategy as in the proof of [Lusztigunipotentch, Lemma 4.3], which uses faithfully flat descent of the smoothness property along the Lang map.
+Proposition 2.4.1.
+For $w_{1},\ldots,w_{r}\in W$, the compactified Deligne–Lusztig variety $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is strongly $F_{p}$-regular.
+In particular, $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is $F_{p}$-rational and thus normal and Cohen-Macaulay.
+If $\overline{Bw_{i}B}$ is smooth for all $i$, then $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is smooth.
 Proof.
-The proof is similar to the proof of [Wan, Proposition A.6]. For each $i=1,\ldots,r$, we know that $\overline{{B}w_{i}{B}}$ is strongly $F$-regular [BT, Cor. 4.1, p. 958]. Thus the product
+When $\overline{Bw_{i}B}$ is smooth for all $i$, this is proved by faithfully flat descent of smoothness in [Lusztigunipotentch, Lemma 4.3] and [DMR, Proposition 2.3.5], cf.
+[GoerWedhAlgGeom1, Proposition 14.59].
+Analogously, we use faithfully flat descent of strongly $F_{p}$-regularity to show that any compactified Deligne–Lusztig variety is strongly $F_{p}$-regular.
+For each $i=1,\ldots,r$, we know that $\overline{{B}w_{i}{B}}$ is strongly $F_{p}$-regular [BrTh, Cor. 4.1, p. 958] and thus Cohen-Macaulay [BrTh, Cor. 4.2, p. 959]. The product
 $\overline{{B}w_{1}{B}}\times_{{\rm Spec}\ k}\cdots\times_{{\rm Spec}\ k}\overline{{B}w_{r}{B}}$
-is strongly $F$-regular by faithfully flat descent [HH, Theorem 5.5].
-Recall that the Lang map $\mathcal{L}_{1}$ is étale, so its closed fibres are smooth and thus strongly $F$-regular cf.
-[Smi, Ch 3, Remark 1.6, Theorem 4.4(a)].
-We may now apply [Abe, Theorem 3.6] and conclude that
+is strongly $F_{p}$-regular by faithfully flat descent along the projection map ${\rm pr}_{i}:\overline{{B}w_{1}{B}}\times_{{\rm Spec}\ k}\cdots\times_{{\rm Spec}\ k}\overline{{B}w_{r}{B}}\to\overline{{B}w_{i}{B}}$ [HoHu, Theorem 5.5.b].
+Recall that the Lang map $\mathcal{L}_{1}$ is étale, so its closed fibres are smooth and thus strongly $F_{p}$-regular, cf.
+[KSmith, Ch 3, Remark 1.6, Theorem 4.4(a)].
+We may now apply [Aber, Theorem 3.6] and conclude that
 $\mathcal{L}_{1}^{-1}\left(\overline{{B}w_{1}{B}}\times_{{\rm Spec}\ k}\cdots\times_{{\rm Spec}\ k}\overline{{B}w_{r}{B}}\right)$
-is strongly $F$-regular.
-Finally, by Lemma 2.8 and the faithfully flat descent of strong $F$-regularity [HH, Theorem 5.5], $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is strongly $F$-regular.
-Strongly $F$-regular rings are $F$-rational, so $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is normal and Cohen-Macaulay by [HH, Theorem 4.2, 6.27].
-∎
-Remark 2.14.
+is strongly $F_{p}$-regular.
+Finally, by Lemma 2.2.1 and the faithfully flat descent of strong $F_{p}$-regularity [HoHu, Theorem 5.5], $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is strongly $F_{p}$-regular.
+Strongly $F_{p}$-regular rings are $F_{p}$-rational, so $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is normal and Cohen-Macaulay by [HoHu, Theorem 4.2, 6.27]. ∎
+Remark 2.4.2.
 For $w\in W$, the Zariski closure $\overline{X(w)}$ in $G/B$ is not always smooth, but for any $v\in S^{*}$ that gives a reduced expression of $w$, the projection map ${\overline{X}^{S^{*}}(v)}\to\overline{X(w)}$ is a proper birational morphism [DL, Lemma 9.11].
-Since $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is normal by 2.13, the connectedness criterion for $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ in 2.12 applies readily for irreducibility.
-Since smoothness can be verified étale locally and satisfies faithfully flat descent, the same proof as in 2.13 gives an alternative proof for the smoothness result in Lemma 2.3.
-Corollary 2.15.
-Let $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ be as in 2.13.
-If for all $i=1,\ldots,r$, the Schubert variety $\overline{{B}w_{i}{B}}/{B}$ is smooth, then $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is smooth.
-In particular, ${\overline{X}^{S^{*}}(w)}$ is smooth for all $w\in S^{*}$.
-Example 2.16.
-If for all $i=1,\ldots,r$, $w_{i}$ is the longest element in $W_{{\text{supp}}_{F}(w_{i})}$, then $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is smooth.
 3. Braid relations in the Weyl group and birational morphisms
@@
 We consider some specific elements in the Weyl group for which the compactified Deligne–Lusztig variety is smooth.
-To begin, we recall the relations in the Weyl group of finite reductive groups cf.
-[GP].
-All the relations in the group presentation of the Weyl group (as a finite Coxeter group $(W,S)$) can be written as
-$(st)^{m_{st}}=1,\ s,t\in S,$ (3.1)
+To begin, we recall the relations in the Weyl group of finite reductive groups, cf.
+[GeckPfeiffer].
+All the relations in the group presentation of the Weyl group as a finite Coxeter group $(W,S)$ can be written as
+$(st)^{m_{st}}=1,\ s,t\in S,$ (1)
 where $m_{st}\in\mathbb{Z}$ such that $m_{ss}=1$ and $m_{st}\geq 2$ if $s\neq t$.
-In the case of $s\neq t$, we call (3.1) a braid relation in the Weyl group $W$.
-The classification of finite groups of Lie type ([DM, Classification 4.3.6]) yields that for $s\neq t$, we have $m_{st}\in\{2,3,4,6\}$.
-Remark 3.1.
-Whenever $(st)^{m_{st}}=1,\ s,t\in S$, is a braid relation in $W$,
-$(ts)^{m_{st}}=1$ (3.2)
-also holds. This is a consequence of the relation $ss=1$ for all $s\in S$. We also have relations
-$(F(s)F(t))^{m_{st}}=1\hskip 14.22636pt\text{and}\hskip 14.22636pt(F(t)F(s))^{m_{st}}=1$ (3.3)
-corresponding to the braid relation (3.1). When there is no ambiguity, we refer to any equation among (3.1), (3.2) and (3.3) as a braid relation in $W$.
-Definition 3.2.
-Let $w\in S^{*}$. If $w=w_{1}w_{2}$, for some $w_{1},w_{2}\in S^{*}$, we say that $w_{i}$ divides $w$, $i=1,2$. Moreover, we call $w_{1}$ a prefix of $w$ and $w_{2}$ a suffix of $w$.
-Definition 3.3.
+In the case of $s\neq t$, we call (1) a braid relation in the Weyl group $W$.
+We refer to such an equation (1) as a braid relation in $W$.
+Remark 3.1.1.
+If $F(S)=S$, then whenever $(st)^{m_{st}}=1,\ s,t\in S$ is a braid relation in $W$, the equations
+$(ts)^{m_{st}}=1,\hskip 14.22636pt\ (F(s)F(t))^{m_{st}}=1,\hskip 14.22636pt\text{and}\hskip 14.22636pt(F(t)F(s))^{m_{st}}=1$
+also hold.
+For $w=w_{1}w_{2}\in S^{*}$, we call $w_{1}$ a prefix of $w$ and $w_{2}$ a suffix of $w$.
+Definition 3.1.2.
 Let $(st)^{m}=1$, with $\ s,t\in S$ and $s\neq t$, be a braid relation in $W$. We define
-$P_{m}(s,t):=st\cdots\in S^{*}$ (3.4)
-such that ${{\ell^{S^{*}}}(P_{m}(s,t))}=m$ and $P_{m}(s,t)$ is a prefix of $(st)^{m}$ in $S^{*}$.
-When there is no ambiguity, we also use $P_{m}(s,t)$ to denote its canonical projection $\alpha(P_{m}(s,t))\in W$.
-Lemma 3.4.
+$P(s,t):=st\cdots\in S^{*},$ (2)
+such that ${{\ell^{S^{*}}}(P(s,t))}=m$ and $P(s,t)$ is a prefix of $(st)^{m}$ in $S^{*}$. When there is no ambiguity, we also use $P(s,t)$ to denote its canonical projection $\alpha(P(s,t))\in W$.
+Lemma 3.1.3.
 Let $w_{1},\ldots,w_{r}\in W$.
-Suppose that for all $i\in\{1,\ldots,r\}$, we have $w_{i}=P_{m}(s,t)$ for some braid relation in $W$.
+Suppose that for all $j\in\{1,\ldots,r\}$, we have $w_{j}=P(s,t)$ for some braid relation in $W$.
 Then the compactified Deligne–Lusztig variety $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is smooth.
 Proof.
-By Lemma 2.3 (resp. Corollary 2.15), it suffices to show that for all $i=1,\ldots,r$, the associated Schubert variety $\overline{{B}w_{i}{B}}/{B}$ is smooth.
-We show this by using the classification of finite groups of Lie type and the classification of finite Coxeter groups.
-Denote $I=\{s,t\}\subseteq S$, $W_{I}\subseteq W$ the corresponding subgroup generated by $I$, and $P_{I}\subseteq G$, $L_{I}\subseteq P_{I}$ the parabolic and Levi subgroups respectively.
-Note that $L_{I}$ is reductive and $P_{I}/{B}$ is isomorphic to $L_{I}/({B}\cap L_{I})$, which is a flag variety and thus is smooth.
-We proceed by showing that $S(P_{m}(s,t))=P_{I}/{B}$.
-Since $W_{I}$ is generated by two elements, the longest element is determined by the braid relation.
-When $L_{I}$ is of type $A_{1}\times A_{1}$ (resp.
-${{\ell^{S^{*}}}(P_{2}(s,t))}=2$), $st\in W_{I}$ is the longest element.
-When $L_{I}$ is of type $A_{2}$ or ${}^{2}A_{2}$ (resp.
-${{\ell^{S^{*}}}(P_{3}(s,t))}=3$), $sts\in W_{I}$ is the longest element.
-When $L_{I}$ is of type $B_{2}$ or ${}^{2}B_{2}$ (resp.
-${{\ell^{S^{*}}}(P_{4}(s,t))}=4$), $stst\in W_{I}$ is the longest element.
-When $L_{I}$ is of type $G_{2}$ or ${}^{2}G_{2}$ (resp.
-${{\ell^{S^{*}}}(P_{6}(s,t))}=6$), $ststst\in W_{I}$ is the longest element.
-Therefore $S(P_{m}(s,t))=P_{I}/{B}$ for $m=2,3,4,6$.
+By Proposition 2.4.1, it suffices to show that for all $j=1,\ldots,r$, the associated shceme $\overline{{B}w_{j}{B}}$ is smooth.
+Consider $w_{j}=P(s,t)$ with $(st)^{m}=1$. Denote $J=\{s,t\}\subseteq S$. The the corresponding subgroup $W_{J}\subseteq W$ generated by $J$ has a group presentation
+$W_{J}=\langle s,t|(st)^{m}=1,s^{2}=1,t^{2}=1\rangle.$
+As a consequence, $P(s,t)$ is a reduced expression of the longest element of $W_{J}$.
+Let $P_{J}=BW_{J}B\subseteq G$ be the parabolic subgroup corresponding to $W_{J}$.
+Since $w_{j}$ is the longest element in $W_{J}$, we know that $\overline{{B}w_{j}{B}}=P_{J}$ and is smooth cf.
+[DigneMichelBook, Proposition 3.2.8].
 ∎
-Definition 3.5.
-Denote by $B^{+}$ the monoid generated by $S$ and all the relations of the form
-$P_{m}(s,t)=P_{m}(t,s)$
-with $s,t\in S$, $t\neq s$, such that $(st)^{m}=1$ is a braid relation in $W$.
-More details about this monoid can be found in [GP, §4].
 3.2. Higher direct images of the structure sheaf for birational morphisms
-Definition 3.6.
-For any $v\in S^{*}$ such that $v=P_{m}(s,t)$ for a braid relation (3.1, 3.2, 3.3) in $W$, we set $R_{st}(v):=P_{m}(t,s)$.
-Proposition 3.7.
-Let $w_{1},\ldots,w_{r}\in W$ such that their corresponding Schubert varieties $\overline{{B}w_{i}{B}}/{B}$ are smooth for all $i=1,\ldots,r$.
-Suppose for $j\in\{1,\ldots,r\}$, the reduced expression of $w_{j}$ is of the form $w_{j}=P_{m}(s,t)$ for a braid relation in $W$.
-Then for all $k\geq 0$, there are $G^{F}$-equivariant isomorphisms
-$H^{k}\left(X_{1},\mathcal{O}_{X_{1}}\right)\overset{\sim}{\longrightarrow}H^{k}\left(X_{2},\mathcal{O}_{X_{2}}\right)$
+Definition 3.2.1.
+For any $v\in S^{*}$ such that $v=P(s,t)$ for a braid relation (1) in $W$, we set $R(s,t):=P(t,s)$.
+Proposition 3.2.2.
+Let $w_{1},\ldots,w_{r}\in W$ such that $\overline{{B}w_{j}{B}}$ is smooth for all $j=1,\ldots,r$.
+Suppose for $j\in\{1,\ldots,r\}$, the reduced expression of $w_{j}$ is of the form $w_{j}=P(s,t)$ for a braid relation in $W$.
+Then for all $i\geq 0$, there are $G^{F}$-equivariant isomorphisms
+$H^{i}\left(X_{1},\mathcal{O}_{X_{1}}\right)\overset{\sim}{\longrightarrow}H^{i}\left(X_{2},\mathcal{O}_{X_{2}}\right)$
 and
-$H^{k}\left(X_{1},\omega_{X_{1}}\right)\overset{\sim}{\longrightarrow}H^{k}\left(X_{2},\omega_{X_{2}}\right),$
-for $X_{i}={\overline{X}^{S^{*}}(w_{1}\cdots w_{r})}$, ${\overline{X}^{S^{*}}(w_{1}\cdots R_{st}(w_{j})\cdots w_{r})}$, $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ and $i=1,2$.
+$H^{i}\left(X_{1},\omega_{X_{1}}\right)\overset{\sim}{\longrightarrow}H^{i}\left(X_{2},\omega_{X_{2}}\right),$
+for $X_{n}$ $(n\in\{1,2\})$ being one of the $k$-schemes ${\overline{X}^{S^{*}}(w_{1}\cdots w_{r})}$, ${\overline{X}^{S^{*}}(w_{1}\cdots R(s,t)\cdots w_{r})}$, and $X(\underline{w_{1}},\ldots,\underline{w_{r}})$, where $R(s,t)$ replaces $w_{j}$.
 Proof.
-By assumption, we have $w_{j}=R_{st}(w_{j})$ in $W$, so
-$X(\underline{w_{1}},\ldots,\underline{w_{r}})=X(\underline{w_{1}},\ldots,\underline{R_{st}(w_{j})},\ldots,\underline{w_{r}}).$
-For $X_{i}={\overline{X}^{S^{*}}(w_{1}\cdots w_{r})}$, ${\overline{X}^{S^{*}}(w_{1}\cdots R_{st}(w_{j})\cdots w_{r})}$, $X(\underline{w_{1}},\ldots,\underline{w_{r}})$, it follows from the assumptions and Lemma 2.3 that $X_{i}$ is smooth and projective.
-Note that we have $I:={\text{supp}}_{F}(w_{1}\cdots w_{r})={\text{supp}}_{F}(w_{1}\cdots R_{st}(w_{j})\cdots w_{r})={\text{supp}}_{F}(w_{1},\ldots,w_{r})$, and by 2.12 $X_{i}$ has $|G^{F}/P_{I}^{F}|$ irreducible components.
-There are natural projection maps ${\rm pr}_{0,\ell(w_{i})}:\overline{O}^{S^{*}}(w_{i})\to O(\underline{w_{i}})$ for all $i=1,\ldots r$ and ${\rm pr}_{0,\ell(w_{j})}:\overline{O}^{S^{*}}(R_{st}(w_{j}))\to O(\underline{w_{j}})$.
+By assumption, we have $w_{j}=P(s,t)=R(s,t)$ in $W$, so replacing $w_{j}$ by $R(s,t)$ does not affect the following compactified Deligne–Lusztig variety
+$X(\underline{w_{1}},\ldots,\underline{w_{r}})=X(\underline{w_{1}},\ldots,\underline{R(s,t)},\ldots,\underline{w_{r}}).$
+Replace $w_{j}$ by $R(s,t)$ in the expression $w_{1}\cdots w_{r}\in S^{*}$ and denote the resulting element by $w_{1}\cdots R(s,t)\cdots w_{r}$.
+For $X_{n}={\overline{X}^{S^{*}}(w_{1}\cdots w_{r})}$, ${\overline{X}^{S^{*}}(w_{1}\cdots R(s,t)\cdots w_{r})}$, $X(\underline{w_{1}},\ldots,\underline{w_{r}})$, it follows from the assumptions and Lemma 3.1.3 that $X_{n}$ is smooth and projective.
+Note that we have $I:={\text{supp}}_{F}(w_{1}\cdots w_{r})={\text{supp}}_{F}(w_{1}\cdots R(s,t)\cdots w_{r})={\text{supp}}_{F}(w_{1},\ldots,w_{r})$, and by Proposition 2.3.3 $X_{n}$ has $|G^{F}/P_{I}^{F}|$ irreducible components.
+There are natural projection maps ${\rm pr}_{0,\ell(w_{j})}:\overline{O}^{S^{*}}(w_{j})\to O(\underline{w_{j}})$ and ${\rm pr}_{0,\ell(w_{j})}:\overline{O}^{S^{*}}(R(s,t))\to O(\underline{w_{j}})$ for all $j=1,\ldots r$.
 Similarly, the projection to the 00-th, $\ell(w_{1})$-th, $\ell(w_{1})+\ell(w_{2})$-th, .
@@
 ., and $\ell(w_{1})+\cdots+\ell(w_{r})$-th coordinate yields the vertical maps in the cartesian diagrams
-[…]
+${{\overline{X}^{S^{*}}(w_{1}\cdots w_{r})}}$${\overline{O}^{S^{*}}(w_{1}\cdots w_{r})}$${X(\underline{w_{1}},\ldots,\underline{w_{r}})}$${O(\underline{w_{1}},\ldots,\underline{w_{r}}),}$$\scriptstyle{f}$
 and
-[…]
+${{\overline{X}^{S^{*}}(w_{1}\cdots R(s,t)\cdots w_{r})}}$${\overline{O}^{S^{*}}(w_{1}\cdots R(s,t)\cdots w_{r})}$${X(\underline{w_{1}},\ldots,\underline{w_{r}})}$${O(\underline{w_{1}},\ldots,\underline{w_{r}}).}$$\scriptstyle{f^{\prime}}$
 The maps $f$ and $f^{\prime}$ are $G^{F}$-equivariant proper birational morphisms.
-It follows from the proof of Zariski’s main theorem ([Har, Corollary 11.4]) that $\phi_{*}\mathcal{O}_{X_{1}}\cong\mathcal{O}_{X_{2}}$, for $\phi=f$ (resp.
+It follows from the proof of Zariski’s main theorem ([Hartshorne, Corollary 11.4]) that $\phi_{*}\mathcal{O}_{X_{1}}\cong\mathcal{O}_{X_{2}}$, for $\phi=f$ (resp.
 $\phi=f^{\prime}$), $X_{1}={\overline{X}^{S^{*}}(w_{1}\cdots w_{r})}$ (resp.
-$X_{1}={\overline{X}^{S^{*}}(w_{1}\cdots R_{st}(w_{j})\cdots w_{r})}$) and $X_{2}=X(\underline{w_{1}},\ldots,\underline{w_{r}})$.
-Moreover, applying [CR, Corollary 3.2.10] yields $R\phi_{*}\mathcal{O}_{X_{1}}\cong\mathcal{O}_{X_{2}}$.
+$X_{1}={\overline{X}^{S^{*}}(w_{1}\cdots R(s,t)\cdots w_{r})}$) and $X_{2}=X(\underline{w_{1}},\ldots,\underline{w_{r}})$.
+Moreover, applying [CR11, Corollary 3.2.10] yields $R\phi_{*}\mathcal{O}_{X_{1}}\cong\mathcal{O}_{X_{2}}$.
 Hence the Leray spectral sequence associated to $\phi$ degenerates.
 Since $\phi$ is $G^{F}$-equivariant, the isomorphism $\phi_{*}\mathcal{O}_{X_{1}}\cong\mathcal{O}_{X_{2}}$ is $G^{F}$-equivariant, thus functoriality of the associated Leray spectral sequence yields $G^{F}$-equivariant isomorphisms on cohomology groups
-$H^{k}\left(X_{1},\mathcal{O}_{X_{1}}\right)\overset{\sim}{\longrightarrow}H^{k}\left(X_{2},\mathcal{O}_{X_{2}}\right),$
-for all $k\geq 0$. The $G^{F}$-equivariant isomorphisms for the cohomology of the canonical sheaf follow from an equivariant version of Serre duality [Nir, Corollary 2.10]. ∎
-4. $\mathbb{P}^{1}$-fibrations and Cyclic shifting
-We construct $\mathbb{P}^{1}$-fibrations between compactified Deligne–Lusztig varieties and relate their cohomology for the structure sheaf and canonical sheaf.
-Moreover, we use these $\mathbb{P}^{1}$-fibrations to indirectly show that the partial Frobenius on compactified Deligne–Lusztig varieties does not affect the cohomology for the structure sheaf and canonical sheaf.
+$H^{i}\left(X_{1},\mathcal{O}_{X_{1}}\right)\overset{\sim}{\longrightarrow}H^{i}\left(X_{2},\mathcal{O}_{X_{2}}\right),$
+for all $i\geq 0$. The $G^{F}$-equivariant isomorphisms for the cohomology of the canonical sheaf follow from an equivariant version of Serre duality [Nironi, Corollary 2.10]. ∎
+4. $\mathbb{P}^{1}$-fibrations and $F$-conjugacy classes
+We construct $\mathbb{P}^{1}$-fibrations between compactified Deligne–Lusztig varieties and relate their cohomology for the structure sheaf and the canonical sheaf.
+Moreover, we use these $\mathbb{P}^{1}$-fibrations to show that pairs of compactified Deligne–Lusztig varieties of the form ${\overline{X}^{S^{*}}(sv)}$ and ${\overline{X}^{S^{*}}(vF(s))}$ have isomorphic cohomology groups for the the structure sheaf and the canonical sheaf.
 4.1. $\mathbb{P}^{1}$-fibrations over Deligne–Lusztig varieties
-Lemma 4.1.
-Let $w_{1},\ldots,w_{r}\in W$ and $s\in S$ such that $F(s)\in S$. Then there is a $G^{F}$-equivariant morphism of $k$-schemes
-$\pi_{1}:X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})\longrightarrow X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)}),$
-which is Zariski-locally a $\mathbb{P}^{1}$-bundle.
+Lemma 4.1.1.
+Let $w_{1},\ldots,w_{r},w_{1}^{\prime},\ldots,w_{m}^{\prime}\in W$ and $s\in S$. Then the projection map
+$\phi:X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{s},\underline{s},\underline{w_{1}^{\prime}},\ldots,\underline{w_{m}^{\prime}})\longrightarrow X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{s},\underline{w_{1}^{\prime}},\ldots,\underline{w_{m}^{\prime}})$
+given by
+$(B_{0},\ldots,B_{r},B_{r+1},B_{r+2},\ldots,B_{r+m+1},FB_{0})\mapsto(B_{0},\ldots,B_{r},B_{r+2},\ldots,B_{r+m+1},FB_{0})$ is a $G^{F}$-equivariant Zariski $\mathbb{P}^{1}$-bundle.
+If $F(s)\in S$, then the $G^{F}$-equivariant morphism of $k$-schemes
+$\pi:X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})\longrightarrow X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)}),$
+given by $(B_{0},\ldots,B_{r+1},FB_{0})\mapsto(B_{1},\ldots,B_{r+1},FB_{1})$, is a Zariski $\mathbb{P}^{1}$-bundle.
 Proof.
@@
 [DL, §1.2].
-By setting $t=F(s)$ and base change to $O(\underline{w_{1}},\ldots,\underline{w_{r}})$, we get again a $\mathbb{P}^{1}$-bundle
-${\rm pr}_{F(s)}:O(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)},\underline{F(s)})\longrightarrow O(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)}).$
-The partial Frobenius map induces a morphism of $k$-schemes
-$F_{1}:X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})\longrightarrow X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)},\underline{F(s)}).$
-As $X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)},\underline{F(s)})$ is a closed subscheme of $O(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)},\underline{F(s)})$, we may compose ${\rm pr}$ and $F_{1}$.
-It follows from the construction that
-${\rm pr}_{F(s)}(X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)},\underline{F(s)}))\subseteq{\rm pr}_{0,r+1}^{-1}(\Gamma_{F}),$
-where ${\rm pr}_{0,r+1}$ is the projection map
-${\rm pr}_{0,r+1}:O(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})\longrightarrow G/B\times G/B.$
-Thus the image of the composition ${\rm pr}_{F(s)}\circ F_{1}$ is contained in $X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$.
-We define
-$\pi_{1}:={\rm pr}_{F(s)}\circ F_{1}:X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})\longrightarrow X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)}),$
-which is $G^{F}$-equivariant by construction.
-For any $(B_{0},B_{1})\in O(\underline{s})$, we have $(FB_{0},FB_{1})\in O(\underline{F(s)})$.
-Recall that via the projection to the first $(r+2)$-entries $X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$ identifies with a closed subscheme of $(G/B)^{r+1}$.
+For $t=s$, we see that $\phi$ is the base change of $O(\underline{t},\underline{t})\to O(\underline{t})$ along the projection map $X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{t},\underline{w_{1}^{\prime}},\ldots,\underline{w_{m}^{\prime}})\to O(\underline{t})$ given by $(B_{0}^{\prime},\ldots,FB_{0}^{\prime})\mapsto(B_{r}^{\prime},B_{r+1}^{\prime})$.
+The first claim follows.
+Assume $F(s)\in S$ and consider $t=F(s)$.
+One can verify that $\pi$ is $G^{F}$-equivariant by its definition.
+Let $(B_{0},B_{1})\in O(\underline{s})$, so $(FB_{0},FB_{1})\in O(\underline{F(s)})$.
+Recall that via the projection to the first $(r+2)$-entries, $X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$ can be identified with a closed subscheme of $(G/B)^{r+2}$.
 Under this identification, we have an embedding
@@
 Thus $\iota$ is surjective with inverse given by $((B_{0}^{\prime},B_{1}^{\prime}),(B_{1}^{\prime},\ldots,B_{r+1}^{\prime}))\mapsto(B_{0}^{\prime},\ldots,B_{r+1}^{\prime})$.
-Note that here we have $(B_{r+1}^{\prime},FB_{0}^{\prime})\in O(\underline{F(s)})$.
-This yields that $X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$ is isomorphic to $O(\underline{s})\times_{G/B}X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$ and that $\pi_{1}$ is the composition of the isomorphism $\iota$ with the projection
+Therefore $X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$ is isomorphic to $O(\underline{s})\times_{G/B}X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$.
+It follows that $\pi$ is the composition of the isomorphism $\iota$ with the projection
 $O(\underline{s})\times_{G/B}X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})\longrightarrow X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)}),$
-which is a Zariski locally trivial $\mathbb{P}^{1}$-bundle with a trivialising section. ∎
-Remark 4.2.
-The reason for the fibre product of $X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$ with $O(\underline{s})$ yielding a Deligne–Lusztig variety is that the orbit of pairs $(B_{0},B_{1})$ determines the orbit of $(FB_{0},FB_{1})$.
-This allows us to verify that the tuples of the Borel subgroups live in the fibre product of the graph of Frobenius and a product of orbits.
+which is a Zariski $\mathbb{P}^{1}$-bundle. ∎
+Remark 4.1.2.
+The reason for the fibre product of $X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$ with $O(\underline{s})$ to yield a Deligne–Lusztig variety is that the orbit of pairs $(B_{0},B_{1})$ determines the orbit of $(FB_{0},FB_{1})$.
+This allows us to verify that the tuples $(B_{0},\ldots,B_{r+1},FB_{0})$ live in the fibre product of the graph of $F$ and a product of orbits.
 In general, taking a fibre product of a Deligne-Lusztig variety with $O(\underline{s})$ does not necessarily yield another Deligne–Lusztig variety.
-Lemma 4.3.
-Let $w_{1},\ldots,w_{r}\in W$ and $s\in S$ such that $F(s)\in S$. There are $G^{F}$-equivariant étale $\mathbb{P}^{1}$-bundles
-$\pi_{2}:X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})},\underline{s})\longrightarrow X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}}),$
-and
-$F\circ\pi_{1}:X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})},\underline{s})\longrightarrow X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)}).$
+Lemma 4.1.3.
+Let $w_{1},\ldots,w_{r}\in W$ and $s\in S$ such that $F(s)\in S$ and $\overline{Bw_{n}B}$ is smooth for $n=1,\ldots,r$. The morphisms
+$\tau_{2}:X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})},\underline{s})\longrightarrow X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}}),$
+given by $(B_{0},B_{1},\ldots,B_{r+1},FB_{0})\longmapsto(B_{r+1},FB_{1},\ldots,FB_{r+1})$, and
+$\tau_{1}:X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})},\underline{s})\longrightarrow X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)}),$
+given by $(B_{0},B_{1},\ldots,B_{r+1},FB_{0})\longmapsto(FB_{1},FB_{2}\ldots,F^{2}B_{1})$, are $G^{F}$-equivariant fppf $\mathbb{P}^{1}$-bundles.
 Proof.
-Define $\pi_{2}$ by
-$(B_{0},B_{1},\ldots,B_{r+1},FB_{0})\longmapsto(B_{r+1},FB_{1},\ldots,FB_{r+1}).$
-This is equivariant under the diagonal $G^{F}$-action.
-Since $(B_{0},B_{1})\in O(\underline{s})$, we know that $(FB_{0},FB_{1})\in O(F(\underline{s}))$.
-By construction we have $(B_{r+1},FB_{0})\in O(F(\underline{s}))$, so $(B_{r+1},FB_{1})\in O(F(\underline{s}))$ [DL, §1.2].
-Let $Z$ be given by the fibre product
+Given $(B_{0},B_{1})\in O(\underline{s})$, we know that $(FB_{0},FB_{1})\in O(F(\underline{s}))$.
+By construction we have $(B_{r+1},FB_{0})\in O(F(\underline{s}))$, so $(B_{r+1},FB_{1})\in O(F(\underline{s}))$ and $(FB_{r+1},F^{2}B_{1})\in O(F(\underline{s}))$ [DL, §1.2].
+Thus $\tau_{1}$ and $\tau_{2}$ are well-defined.
+They are $G^{F}$-equivariant by construction.
+Let $Z$ be the fibre product
 […]
-where $F=F_{1}^{r+1}$ is the Frobenius map. Points in $Z$ are denoted as tuples
+Geometric points in $Z$ are denoted as tuples
 $((B_{0},B_{1},\ldots,B_{r+1},FB_{0}),(B_{0}^{\prime},B_{1}^{\prime},\ldots,B_{r}^{\prime},FB_{0}^{\prime})),$
@@
 $(B_{0}^{\prime},B_{1}^{\prime},\ldots,B_{r}^{\prime},FB_{0}^{\prime})\in X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})})$ such that $FB_{0}^{\prime}=B_{r+1}$, $FB_{1}^{\prime}=FB_{1},\ldots,FB_{r}^{\prime}=FB_{r},F^{2}B_{0}^{\prime}=FB_{r+1}$.
-Here the only condition for $B_{0}$ is that $(B_{0},B_{0}^{\prime})\in O(\underline{F^{-1}(s)})$.
+Here the only condition on $B_{0}$ is that $(B_{0},B_{0}^{\prime})\in O(\underline{F^{-1}(s)})$.
 The morphism $d_{1}$ also fits into another commutative diagram
-[…]
+${Z}$${O(\underline{F^{-1}(s)})}$${X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})})}$${G/B,}$$\scriptstyle{r_{2}}$$\scriptstyle{d_{1}}$$\scriptstyle{{\rm pr}_{1}}$$\scriptstyle{{\rm pr}_{0}}$ (4)
 where ${\rm pr}_{1}:O(\underline{s})\to G/B$ is given by $(B_{0},B_{1})\mapsto B_{1}$, and $r_{2}$ is given by
 $((B_{0},B_{1},\ldots,B_{r+1},FB_{0}),(B_{0}^{\prime},B_{1}^{\prime},\ldots,B_{r}^{\prime},FB_{0}^{\prime}))\longmapsto(B_{0},B_{0}^{\prime}).$
-The diagram (4.2) yields an injective morphism
-$\iota:Z\longrightarrow O(\underline{s})\times_{G/B}X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})}).$
-On the other hand, given $(B_{0}^{\prime},B_{1}^{\prime},\ldots,B_{r}^{\prime},FB_{0}^{\prime})\in X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})})$ and arbitrary $B_{0}\in G/B$ with $(B_{0},B_{0}^{\prime})\in O(\underline{F^{-1}(s)})$, we have $(B_{0},B_{1}^{\prime})\in O(\underline{F^{-1}(s)})$ and $(FB_{0},FB_{0}^{\prime})\in O(\underline{s})$.
+The diagram (4) yields a morphism
+$\iota:Z\longrightarrow O(\underline{F^{-1}(s)})\times_{G/B}X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})}).$
+Given $(B_{0}^{\prime},B_{1}^{\prime},\ldots,B_{r}^{\prime},FB_{0}^{\prime})\in X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})})$ and arbitrary $B_{0}\in G/B$ with $(B_{0},B_{0}^{\prime})\in O(\underline{F^{-1}(s)})$, we have $(B_{0},B_{1}^{\prime})\in O(\underline{F^{-1}(s)})$ and $(FB_{0},FB_{0}^{\prime})\in O(\underline{s})$.
 This gives
@@
 and that $\iota$ is surjective.
-Hence $\iota$ is an isomorphism and $d_{1}$ is a Zariski $\mathbb{P}^{1}$-bundle.
-Let $\{U_{\alpha}\}_{\alpha}$ be a Zariksi covering of $X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})})$ over which $d_{1}$ trivialises.
-Since $F:X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})})\to X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}})$ is surjective and étale, $\{F(U_{\alpha})\}_{\alpha}$ is an étale covering of $X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}})$ over which $\pi_{2}$ trivialises as a $\mathbb{P}^{1}$-bundle.
-Let $\{V_{\alpha}\}_{\alpha}$ be a Zariski covering of $X(\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})},\underline{s})$ over which $\pi_{1}$ trivialises.
-Since $F:X(\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})},\underline{s})\to X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$ is surjective and étale, $\{F(V_{\alpha})\}_{\alpha}$ is an étale covering of $X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$ over which $F\circ\pi_{1}$ trivialises as a $\mathbb{P}^{1}$-bundle.
-Remark 4.4.
-As morphisms for $k$-schemes, the absolute Frobenius may not have a well-defined inverse, but the induced morphism $F:W\to W$ is an automorphism of the Weyl group.
-This is one of the key facts used in Lemma 4.3.
-Our constructions are analogous to the geometric constructions for classical Deligne–Lusztig varieties in the proof of [DL, Theorem 1.6].
-Remark 4.5.
-The maps constructed in Lemma 4.1 and Lemma 4.3 fit into a commutative diagram
-[…]
-Although there is a morphism $F_{1}:X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}})\to X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$, the morphism $F_{1}$ does not directly yield a clear relation between the cohomology of the structure sheaf for $X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$.
-The reason is that $F_{1}$ has degree divisible by $p$, which implies that $F_{1,*}\mathcal{O}_{X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}})}$ is not isomorphic to $\mathcal{O}_{X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})}$.
-However, the étale $\mathbb{P}^{1}$-bundles $\pi_{1}$ and $\pi_{2}$ are projective and flat with geometrically connected and geometrically reduced fibres, so we could establish a clear relation between the cohomology of the structure sheaves of the source and the target under pushforward.
-This leads to the result of the following proposition that relates the cohomology of the structure sheaf (resp.
-canonical sheaf) between $X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$.
-Proposition 4.6.
-Let $w_{1},\ldots,w_{r}\in W$ and $s\in S$ such that $F(s)\in S$. There are $G^{F}$-equivariant isomorphisms for all $k\geq 0$,
-$H^{k}\big{(}X_{1},\mathcal{O}_{X_{1}}\big{)}\overset{\sim}{\longrightarrow}H^{k}\big{(}X_{2},\mathcal{O}_{X_{2}}\big{)}$
+By the Bruhat decomposition, for every $i=1,\ldots,r$, there exists $v_{i}\in W$ such that $(B_{i},B_{i}^{\prime})\in O(v_{i})$.
+This implies $(FB_{i},FB_{i}^{\prime})\in O(F(v_{i}))$.
+We know $FB_{i}=FB_{i}^{\prime}$ by assumption, so $F(v_{i})=e$ is the neutral element in $W$ and $v_{i}=e$ because $F:W\to W$ is an automorphism, cf.
+[DL, §1.2].
+It follows that $\iota$ is also injective and an isomorphism.
+Hence $d_{1}$ is a Zariski $\mathbb{P}^{1}$-bundle.
+Since $F:X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})})\to X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}})$ is faithfully flat by our assumption on smoothness [GoerWedhAlgGeom1, Corollary 14.129], we deduce that $\tau_{2}$ is faithfully flat by faithfully flat descent [GoerWedhAlgGeom1, Corollary 14.12, Proposition 14.3(2)].
+For each affine open (Zariski) covering $\{U_{\alpha}\}_{\alpha}$ of $X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})})$ that trivialises $d_{1}$, we can construct an fppf covering $\{\left.F\right|_{U_{\alpha}}:U_{\alpha}\to X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}})\}_{\alpha}$ of $X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}})$ over which $\tau_{2}$ trivialises as a $\mathbb{P}^{1}$-bundle.
+After taking the fibre product of $\tau_{1}$ with $F:X(\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})},\underline{s})\to X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$, the claim for $\tau_{1}$ can be verified by the analogous argument we used for $\tau_{2}$ above.
+∎
+Proposition 4.1.4.
+Let $w_{1},\ldots,w_{r}\in W$ and $s\in S$ such that $F(s)\in S$ and $\overline{Bw_{n}B}$ is smooth for $n=1,\ldots,r$. There are $G^{F}$-equivariant isomorphisms for all $i\geq 0$,
+$H^{i}\big{(}X_{1},\mathcal{O}_{X_{1}}\big{)}\overset{\sim}{\longrightarrow}H^{i}\big{(}X_{2},\mathcal{O}_{X_{2}}\big{)}$
 and
-$H^{k}\big{(}X_{1},\omega_{X_{1}}\big{)}\overset{\sim}{\longrightarrow}H^{k}\big{(}X_{2},\omega_{X_{2}}\big{)},$
-where for $i=1,2$, $X_{i}=X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}})$, $X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$, $X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$, and $X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})},\underline{s})$.
+$H^{i}\big{(}X_{1},\omega_{X_{1}}\big{)}\overset{\sim}{\longrightarrow}H^{i}\big{(}X_{2},\omega_{X_{2}}\big{)}$
+induced by $f:X_{1}\to X_{2}$, where $f$ is one of the $\mathbb{P}^{1}$-bundles $\phi,\pi,\tau_{1}$ or $\tau_{2}$ from Lemma 4.1.1 and 4.1.3.
 Proof.
-We continue to use the notations of Lemma 4.3.
-By Lemma 2.8, 2.13, and 2.12, the schemes $X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$ are normal, reduced, and equidimensional with $|G^{F}/P_{I}^{F}|$ irreducible components, where $I={\text{supp}}_{F}(s,w_{1},\ldots,w_{r})$.
-From the proof of Lemma 4.3 we know that
-$d_{1}:Z\to X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})})$
-is a Zariski $\mathbb{P}^{1}$-bundle.
-It follows that all fibres of $d_{1}$ are isomorphic to $\mathbb{P}^{1}$ and that $d_{1}$ preserves connected components.
-Thus we may assume without loss of generality that $Z$ and $X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})})$ are irreducible.
-Since $X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})})$ is normal and all fibres of $d_{1}$ are geometrically reduced, a criterion for flatness [GW, Theorem 14.131] implies that $d_{1}$ is flat and thus $\pi_{2}$ is flat by base change.
-Similarly, we have a Zariski $\mathbb{P}^{1}$-bundle
-$\pi_{1}:X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})},\underline{s})\longrightarrow X(\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})},\underline{s}),$
-with $X(\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})},\underline{s})$ normal, so $\pi_{1}$ is flat by loc. cit. Hence the composition
-$F\circ\pi_{1}:X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})},\underline{s})\longrightarrow X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{s})$
-is flat.
-Denote $f:=\pi_{2}$ (resp.
-$F\circ\pi_{1}$), $X_{2}:=X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}})$ (resp.
-$X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{s})$) and $X_{1}:=X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})},\underline{s})$.
-The rest of the proof is similar to [Wan, Proposition 4.4].
-Consider the Leray spectral sequence for $\mathcal{O}_{X_{2}}$:
+The proof is simpler for the Zariski $\mathbb{P}^{1}$-bundles $\phi$ and $\pi$, so we only record the proof for the fppf $\mathbb{P}^{1}$-bundles $\tau_{1}$ and $\tau_{2}$.
+We use the notations of Lemma 4.1.3.
+Since $d_{1}$ is a Zariski $\mathbb{P}^{1}$-bundle, it is proper and faithfully flat.
+Hence $\tau_{1}$ and $\tau_{2}$ are proper and faithfully flat by faithfully flat descent [GoerWedhAlgGeom1, Proposition 14.53].
+Set $f:=\tau_{2}$ (resp.
+$\tau_{1}$), $X_{1}:=X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})},\underline{s})$, $X_{2}:=X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}})$ (resp.
+$X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$), and $X_{2}^{\prime}:=X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})})$ (resp.
+$X(\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})},\underline{s})$).
+Consider the Leray spectral sequence for $\mathcal{O}_{X_{1}}$:
 $E_{2}^{i,j}=H^{i}\left(X_{2},R^{j}f_{*}\mathcal{O}_{X_{1}}\right)\implies H^{i+j}\left(X_{1},\mathcal{O}_{X_{1}}\right).$
-Since $f$ is proper, and all fibres of $f$ are isomorphic to $\mathbb{P}^{1}_{k}$, we have $f_{*}\mathcal{O}_{X_{1}}\cong\mathcal{O}_{X_{2}}$ [SP, Lemma 0E0L].
-This isomorphism is $G^{F}$-equivariant since $f$ is $G^{F}$-equivariant.
-For all $x\in X_{2}$, we have $X_{1}\times_{X_{2}}x\cong\mathbb{P}^{1}_{k}$, so $H^{j}(X_{1}\times_{X_{2}}x,\mathcal{O}_{X_{1}\times_{X_{2}}x})=0$ for all $j>0$.
-Hence the stalk of $R^{j}f_{*}(\mathcal{O}_{X_{1}})$ at $x$ is 00 for all $x\in X_{2}$ and $j>0$ by Grauert’s theorem cf.
-[Vak, Theorem 25.1.6].
-We may conclude that $R^{j}f_{*}\mathcal{O}_{X_{1}}=0$ for all $j>0$.
-Therefore the Leray spectral sequence degenerates and we have isomorphisms
-$H^{i}\left(X_{2},\mathcal{O}_{X_{2}}\right)\overset{\sim}{\longrightarrow}H^{i}\left(X_{1},\mathcal{O}_{X_{1}}\right),$
-for all $i\geq 0$, which are $G^{F}$-equivariant by functoriality of the Leray spectral sequence.
-The same argument applies for the Zariski locally trivial $\mathbb{P}^{1}$-bundle
-$\pi_{1}^{\prime}:X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})\longrightarrow X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)}).$
-The $G^{F}$-equivariant isomorphisms for the cohomology of $\omega_{X_{1}}$ and $\omega_{X_{2}}$ follow from Serre duality for Deligne–Mumford stacks [Nir, Corollary 2.10].
-4.2. $F$-conjugacy classes and $F$-cyclic shifts
+Let $x\in X_{2}$ and $y\in X_{2}^{\prime}$ such that $F(y)=x$ as in diagram (3).
+By Lemma 4.1.3, the base change of the fibre $x\times_{X_{2}}X_{1}$ to ${\rm Spec}(\kappa(y))$ is isomorphic to $\mathbb{P}^{1}_{\kappa(y)}$.
+It follows from flat base change that $H^{j}(x\times_{X_{2}}X_{1},\mathcal{O}_{x\times_{X_{2}}X_{1}})=0$ for all $j\geq 0$ [GoerWedhAlgGeom2, Corollary 22.91].
+We conclude by Grauert’s theorem [GoerWedhAlgGeom2, Proposition 23.142] (cf.
+[GoerWedhAlgGeom1, Proposition 3.35]) that $R^{j}f_{*}\mathcal{O}_{X_{1}}=0$ for all $j>0$.
+Hence the Leray spectral sequence degenerates.
+We have an fppf covering $\mathfrak{U}=\{\gamma_{\alpha}:U_{\alpha}\to X_{2}\}_{\alpha}$ of $X_{2}$ in Lemma 4.1.3, such that the base change $f_{\alpha}:X_{1}\times_{X_{2}}U_{\alpha}\to U_{\alpha}$ of $f$ along $\gamma_{\alpha}$ is a trivial $\mathbb{P}^{1}$-bundle for all $\alpha$.
+Since $U_{\alpha}$ is affine and $\gamma_{\alpha}$ is an affine morphism for all $\alpha$ [GoerWedhAlgGeom1, Proposition 12.3(3)], any simplex $V$ in the Čech nerve of $\mathfrak{U}$ is affine and for any coherent sheaf $\mathcal{F}$ on $X_{2}$, we have $H^{i}(X_{2},\mathcal{F})\overset{\sim}{\longrightarrow}\check{H}^{i}(\mathfrak{U},\mathcal{F})$ for all $i\geq 0$ [Godement, Theorem II.5.4.1, Corollary].
+Note that $f_{*}\mathcal{O}_{X_{1}}$ is a coherent $\mathcal{O}_{X_{2}}$-module because $f$ is proper and cohomology of coherent sheaves is independent of topology [stacks-project, 03DW].
+For each simplex $V$ in the Čech nerve of $\mathfrak{U}$, we have $V\times_{X_{2}}X_{1}\overset{\sim}{\longrightarrow}V\times_{{\rm Spec}\ k}\mathbb{P}^{1}$, so $f_{*}\mathcal{O}_{X_{1}}(V)=\mathcal{O}_{V}(V)$ by flat base change [GoerWedhAlgGeom2, Proposition 22.90].
+This identification is compatible with restriction of sections, so the Čech complexes on $\mathfrak{U}$ for $f_{*}\mathcal{O}_{X_{1}}$ and $\mathcal{O}_{X_{2}}$ are identical and $H^{i}(X_{2},\mathcal{O}_{X_{2}})\overset{\sim}{\longrightarrow}H^{i}(X_{2},f_{*}\mathcal{O}_{X_{1}})$ for all $i\geq 0$.
+Finally, degeneration of the Leray spectral sequence yields $G^{F}$-equivariant isomorphisms
+$H^{i}\left(X_{2},\mathcal{O}_{X_{1}}\right)\overset{\sim}{\longrightarrow}H^{i}\left(X_{1},\mathcal{O}_{X_{1}}\right),$
+for all $i\geq 0$.
+The $G^{F}$-equivariant isomorphisms for the cohomology of $\omega_{X_{1}}$ and $\omega_{X_{2}}$ follow from Serre duality for Deligne–Mumford stacks [Nironi, Corollary 2.10].
+Corollary 4.1.5.
+Let $s\in S$ and assume $F(s)\in S$. Let $w=sv$ and $w^{\prime}=vF(s)$ be elements of $S^{*}$. There are $G^{F}$-equivariant isomorphisms for all $i\geq 0$,
+$H^{i}\Big{(}{\overline{X}^{S^{*}}(w)},\mathcal{O}_{{\overline{X}^{S^{*}}(w)}}\Big{)}\overset{\sim}{\longrightarrow}H^{i}\Big{(}{\overline{X}^{S^{*}}(w^{\prime})},\mathcal{O}_{{\overline{X}^{S^{*}}(w^{\prime})}}\Big{)}$
+and
+$H^{i}\Big{(}{\overline{X}^{S^{*}}(w)},\omega_{{\overline{X}^{S^{*}}(w)}}\Big{)}\overset{\sim}{\longrightarrow}H^{i}\Big{(}{\overline{X}^{S^{*}}(w^{\prime})},\omega_{{\overline{X}^{S^{*}}(w^{\prime})}}\Big{)}.$
+Proof.
+This follows from Proposition 4.1.4 by composing the corresponding isomorphisms of cohomology groups for $\tau_{1}$ and $\tau_{2}$. ∎
+4.2. $F$-conjugacy classes in $W$
 Two elements $w,w^{\prime}\in W$ are $F$-conjugate if $w^{\prime}=v^{-1}wF(v)$ for some $v\in W$.
@@
 An element $w\in W$ is of minimal length if it has the smallest Bruhat length in its $F$-conjugacy class.
-Minimal length elements were studied by Geck, Kim, and Pfeiffer [GKP], as well as He and Nie [HN].
-Under the assumption $F(S)=S$, these two papers showed that for any $w\in W$, there exists a minimal length element $w^{\prime}$ in the $F$-conjugacy class of $w$ such that $w\to w^{\prime}$ (cf.
-4.7).
-In combination with 4.6, this result gives an inductive approach for studying cohomology groups of Deligne–Lusztig varieties.
-Let us recall the definition of $w\to w^{\prime}$.
-Definition 4.7.
-We write $w\to w^{\prime}$ if there exist $w_{1},\ldots,w_{r}\in W$ and $s_{1},\ldots,s_{r-1}\in S$ such that $w_{1}=w,w_{r}=w^{\prime}$, $w_{i}=s_{i-1}w_{i-1}F(s_{i-1})$, and $\ell(w_{i})\leq\ell(w_{i-1})$ for all $i=2,\ldots,r$.
-We say that $w$ and $w^{\prime}$ are conjugate by $F$-cyclic shifts and write $w\approx w^{\prime}$ if there exist $w_{1},\ldots,w_{r}\in W$ and $s_{1},\ldots,s_{r-1}\in S$ such that $w_{1}=w,w_{r}=w^{\prime}$, $w_{i}=s_{i-1}w_{i-1}F(s_{i-1})$, and $\ell(w_{i})=\ell(w_{i-1})$ for all $i=2,\ldots,r$.
-Remark 4.8.
-Let $w,w^{\prime}\in W$ such that $w\approx w^{\prime}$.
-If $v\in S^{*}$ gives a fixed expression for $w$, then $w\approx w^{\prime}$ guarantees that there exists $w_{1},w_{2}\in S^{*}$ such that $v=w_{1}w_{2}$ and $w_{2}F(w_{1})\in S^{*}$ gives a reduced expression for $w^{\prime}$.
-The assumption $F(S)=S$ only excludes the cases when $G$ is of type ${}^{2}B_{2}$, ${}^{2}F_{4}$, or ${}^{2}G_{2}$, which will be treated separately in Section 5.2.
-Proposition 4.9.
-Assume $F(S)=S$. Let $w=w_{1}w_{2}$ and $w^{\prime}=w_{2}F(w_{1})$ for some $w_{1},w_{2}\in S^{*}$. For all $k\geq 0$, we have $G^{F}$-equivariant isomorphisms
-$H^{k}\Big{(}{\overline{X}^{S^{*}}(w)},\mathcal{O}_{{\overline{X}^{S^{*}}(w)}}\Big{)}\overset{\sim}{\longrightarrow}H^{k}\Big{(}{\overline{X}^{S^{*}}(w^{\prime})},\mathcal{O}_{{\overline{X}^{S^{*}}(w^{\prime})}}\Big{)}$
+Minimal length elements were studied by Geck, Kim, and Pfeiffer [GeckKimPfeifferMinlength], as well as He and Nie [HeNieminlength].
+Under the assumption $F(S)=S$, these two papers showed that for any $w\in W$, there exists a minimal length element $w^{\prime}$ in the $F$-conjugacy class of $w$ such that $w\to_{F}w^{\prime}$ (cf.
+Definition 4.2.1).
+In combination with Proposition 4.1.4, this result gives an inductive approach for studying cohomology groups of Deligne–Lusztig varieties.
+Definition 4.2.1.
+We write $w\to_{F}w^{\prime}$ if there exist $w_{1},\ldots,w_{r}\in W$ and $s_{1},\ldots,s_{r-1}\in S$ such that $w_{1}=w,w_{r}=w^{\prime}$, $w_{i}=s_{i-1}w_{i-1}F(s_{i-1})$, and $\ell(w_{i})\leq\ell(w_{i-1})$ for all $i=2,\ldots,r$.
+Furthermore, for any pairs of minimal length elements $w$ and $w^{\prime}$ in an elliptic $F$-conjugacy class $C$ in $W$, we have $w\to_{F}w^{\prime}$ and $w^{\prime}\to_{F}w$, cf.
+[HeNieminlength, Corollary 4.4].
+The definition of an elliptic $F$-conjugacy class, which we do not recall here, can be found in §4.1 of loc.
+cit.
+5. Cohomology of compactified Deligne–Lusztig varieties
+5.1. Main theorem and applications
+Recall that the braid monoid associated to $W$ is a monoid generated by $S$ and all the relations of the form $P(s,t)=P(t,s)$ with $s,t\in S$, $t\neq s$, such that $(st)^{m}=1$ is a braid relation in $W$, cf.
+[GeckPfeiffer, §4].
+We denote the braid monoid by $B^{+}$.
+There is a natural surjective morphism of monoids $\beta:S^{*}\twoheadrightarrow B^{+}$.
+Theorem 5.1.1.
+Assume $F(S)=S$.
+For each $w\in S^{*}$, there exists an $F$-conjugacy class $C$ in $W$ and a minimal length element $x_{0}\in C$ such that for any reduced expression $x\in S^{*}$ of $x_{0}$, there are $G^{F}$-equivariant isomorphisms
+$H^{i}\Big{(}{\overline{X}^{S^{*}}(w)},\mathcal{O}_{{\overline{X}^{S^{*}}(w)}}\Big{)}\overset{\sim}{\longrightarrow}H^{i}\Big{(}{\overline{X}^{S^{*}}(x)},\mathcal{O}_{{\overline{X}^{S^{*}}(x)}}\Big{)}$
 and
-$H^{k}\Big{(}{\overline{X}^{S^{*}}(w)},\omega_{{\overline{X}^{S^{*}}(w)}}\Big{)}\overset{\sim}{\longrightarrow}H^{k}\Big{(}{\overline{X}^{S^{*}}(w^{\prime})},\omega_{{\overline{X}^{S^{*}}(w^{\prime})}}\Big{)}.$
+$H^{i}\Big{(}{\overline{X}^{S^{*}}(w)},\omega_{{\overline{X}^{S^{*}}(w)}}\Big{)}\overset{\sim}{\longrightarrow}H^{i}\Big{(}{\overline{X}^{S^{*}}(x)},\omega_{{\overline{X}^{S^{*}}(x)}}\Big{)},$
+for all $i\geq 0$.
+If $C$ is moreover an elliptic conjugacy class in $W$, then the isomorphisms above hold if we replace $x_{0}$ by any other minimal length element in $C$.
 Proof.
-This follows from 4.6. We have $w^{\prime}=F_{1}^{{{\ell^{S^{*}}}(w_{1})}}(w)$ and the result follows from composing the corresponding isomorphisms on cohomology ${{\ell^{S^{*}}}(w_{1})}$-times. ∎
-Corollary 4.10.
-Assume $F(S)=S$. Let $w\in S^{*}$ such that $w=w_{1}ssw_{2}$ for some $s\in S,w_{1},w_{2}\in S^{*}$. There are $G^{F}$-equivariant isomorphisms for all $k\geq 0$,
-$H^{k}\Big{(}{\overline{X}^{S^{*}}(w)},\mathcal{O}_{{\overline{X}^{S^{*}}(w)}}\Big{)}\overset{\sim}{\longrightarrow}H^{k}\Big{(}{\overline{X}^{S^{*}}(sw_{2}F(w_{1})F(s))},\mathcal{O}_{{\overline{X}^{S^{*}}(sw_{2}F(w_{1})F(s))}}\Big{)},$ (4.3)
-and
-$H^{k}\Big{(}{\overline{X}^{S^{*}}(w)},\omega_{{\overline{X}^{S^{*}}(w)}}\Big{)}\overset{\sim}{\longrightarrow}H^{k}\Big{(}{\overline{X}^{S^{*}}(sw_{2}F(w_{1})F(s))},\omega_{{\overline{X}^{S^{*}}(sw_{2}F(w_{1})F(s))}}\Big{)}.$ (4.4)
-Moreover,
-$H^{k}\big{(}{\overline{X}^{S^{*}}(w)},\mathcal{F}\big{)}\overset{\sim}{\longrightarrow}H^{k}\big{(}{\overline{X}^{S^{*}}(w_{2}F(w_{1})F(s))},\mathcal{F}\big{)}\overset{\sim}{\longrightarrow}H^{k}\big{(}{\overline{X}^{S^{*}}(sw_{2}F(w_{1}))},\mathcal{F}\big{)},$ (4.5)
-for all $k\geq 0$ and $\mathcal{F}$ being the corresponding structure sheaf and canonical sheaf.
-Proof.
-As $sw_{2}F(w_{1})F(s)=F_{1}^{{{\ell^{S^{*}}}(w_{1}s)}}(w)$, the isomorphisms (4.3) and (4.4) follow from 4.9, and the isomorphisms (4.5) follow from composing (4.3) (resp.
-(4.4)) with the isomorphisms in 4.6.
+We prove the theorem for the structure sheaf. The results for the canonical sheaf follows from Serre duality for Deligne–Mumford stacks [Nironi, Corollary 2.10].
+Let $w\in S^{*}$. Consider the image $\beta(w)\in B^{+}$ of $w$ under the projection $\beta:S^{*}\twoheadrightarrow B^{+}$ and let $u\in S^{*}$ be another lift of $\beta(w)$, then we have
+$H^{i}\Big{(}{\overline{X}^{S^{*}}(u)},\mathcal{O}_{{\overline{X}^{S^{*}}(u)}}\Big{)}\overset{\sim}{\longrightarrow}H^{i}\Big{(}{\overline{X}^{S^{*}}(w)},\mathcal{O}_{{\overline{X}^{S^{*}}(w)}}\Big{)}$ (5)
+for all $i\geq 0$ by Proposition 3.2.2. Note that this proposition can be applied to any $t_{1}\ldots t_{r}\in S^{*}$ by setting $w_{j}=t_{j}$ for $j=1,\ldots r$.
+Using the description of the submonoid of reduced elements in $B^{+}$ given by Matsumoto’s theorem [GeckPfeiffer, Theorem 1.2.2, Exercise 4.1], we know that the criterion for $u$ to come from a reduced expression in the Weyl group is that $u$ has no subexpression of the form $ss$ for any $s\in S$.
+If $u=u_{1}ssu_{2}$, we apply Lemma 4.1.1 and Proposition 4.1.4 to obtain $u_{1}su_{2}$ such that ${\overline{X}^{S^{*}}(u)}$ and ${\overline{X}^{S^{*}}(u_{1}su_{2})}$ have the same cohomology for the structure sheaf.
+Inductively, this yields an element $v\in S^{*}$ with ${{\ell^{S^{*}}}(v)}\leq{{\ell^{S^{*}}}(u)}$ such that $v$ has no subexpression of the form $ss$ for any $s\in S$ and for all $i\geq 0$,
+$H^{i}\Big{(}{\overline{X}^{S^{*}}(v)},\mathcal{O}_{{\overline{X}^{S^{*}}(v)}}\Big{)}\overset{\sim}{\longrightarrow}H^{i}\Big{(}{\overline{X}^{S^{*}}(w)},\mathcal{O}_{{\overline{X}^{S^{*}}(w)}}\Big{)}.$
+In other words, we may assume $v\in S^{*}$ gives a reduced expression of $\alpha(v)$ in $W$.
+By [GeckKimPfeifferMinlength, Theorem 2.6], there exists $v^{\prime}\in W$ which is of minimal length in its $F$-conjugacy class such that $\alpha(v)\to_{F}v^{\prime}$ and $\ell(v^{\prime})\leq\ell(\alpha(v))$.
+Suppose $\alpha(v)$ is a minimal length element in its $F$-conjugacy class.
+By Matsumoto’s theorem [GeckPfeiffer, Theorem 1.2.2], there exists a lift $y\in B^{+}$ of $\alpha(v)$ such that any reduced expression $x\in S^{*}$ of $\alpha(v)$ satisfies $\beta(x)=y$.
+The statement of the theorem thus follows from Proposition 3.2.2.
+In general, $\alpha(v)\to_{F}v^{\prime}$ gives $v_{1}^{\prime},\ldots,v_{m}^{\prime}\in W,s_{1},\ldots s_{m-1}\in S$ such that $v_{1}^{\prime}=\alpha(v),v_{m}^{\prime}=v^{\prime}$ and $v_{j}^{\prime}=s_{j-1}v_{j-1}^{\prime}F(s_{j-1})$ for all $j=2,\ldots,m$.
+Let $a\in\{2,\ldots,m\}$ be the smallest integer such that $\ell(v_{a}^{\prime})<\ell(v_{a-1}^{\prime})$.
+For all $j=1,\ldots,a$, let $v_{j}\in S^{*}$ be the canonical lift of the corresponding reduced expression of $v_{j}^{\prime}$.
+Up to taking different lifts of braid relations, for each $j=2,\ldots,a-1$, we have $v_{j}=s_{j-1}z_{j-1}$ (resp.
+$v_{j}=z_{j-1}F(s_{j-1})$) and $v_{j-1}=z_{j-1}F(s_{j-1})$ (resp.
+$v_{j-1}=s_{j-1}z_{j-1}$) for some $z_{j-1}\in S^{*}$ (cf.
+[GeckKimPfeifferMinlength, Remark 2.3]).
+Thus it follows from Corollary 4.1.5 that ${\overline{X}^{S^{*}}(v_{j})}$ all have the same cohomology for the structure sheaf as ${\overline{X}^{S^{*}}(v)}$ for $j=2,\ldots,a-1$.
+On the other hand, we have $v_{a-1}=s_{a-1}v_{a}F(s_{a-1})$ in $S^{*}$.
+We know that ${\overline{X}^{S^{*}}(v)}$ has the same cohomology as ${\overline{X}^{S^{*}}(s_{a-1}v_{a})}$ and ${\overline{X}^{S^{*}}(v_{a}F(s_{a-1}))}$ by Proposition 4.1.4.
+In this case, we reset $v$ to be $s_{a-1}v_{a}$.
+This process will terminate and yields an element $x_{0}\in W$ which is of minimal length in its $F$-conjugacy class, cf.
+[GeckKimPfeifferMinlength, Theorem 2.6].
+It remains to consider the case when $C$ is an elliptic conjugacy class in $W$.
+By [HeNieminlength, Corollary 4.4], we have $x\to_{F}y$ for any two minimal length elements $x,y\in C$.
+The corresponding cohomology of ${\overline{X}^{S^{*}}(x)}$ and ${\overline{X}^{S^{*}}(y)}$ for the structure sheaf are isomorphic by Corollary 4.1.5.
 ∎
-Example 4.11.
-In the case when $w=ss$, we have the following isomorphisms on cohomology for for all $k\geq 0$ and for $\mathcal{F}$ being the structure sheaf (resp. canonical sheaf)
-$H^{k}\big{(}{\overline{X}^{S^{*}}(ss)},\mathcal{F}\big{)}\overset{\sim}{\longrightarrow}H^{k}\big{(}{\overline{X}^{S^{*}}(sF(s))},\mathcal{F}\big{)}\overset{\sim}{\longrightarrow}H^{k}\big{(}{\overline{X}^{S^{*}}(F(s))},\mathcal{F}\big{)}\overset{\sim}{\longrightarrow}H^{k}\big{(}{\overline{X}^{S^{*}}(s)},\mathcal{F}\big{)}.$
-5. Cohomology of compactified Deligne–Lustig varieties
-5.1. Main theorem and applications
-Theorem 5.1.
-Assume $F(S)=S$.
-For each $w\in S^{*}$, there exists an $F$-conjugacy class $C$ in $W$ such that for any $x\in C_{\rm min}$ and any reduced expression of $x$, there are $G^{F}$-equivariant isomorphisms
-$H^{k}\Big{(}{\overline{X}^{S^{*}}(w)},\mathcal{O}_{{\overline{X}^{S^{*}}(w)}}\Big{)}\overset{\sim}{\longrightarrow}H^{k}\Big{(}\overline{X}(x),\mathcal{O}_{\overline{X}(x)}\Big{)}$
-and
-$H^{k}\Big{(}{\overline{X}^{S^{*}}(w)},\omega_{{\overline{X}^{S^{*}}(w)}}\Big{)}\overset{\sim}{\longrightarrow}H^{k}\Big{(}\overline{X}(x),\omega_{\overline{X}(x)}\Big{)},$
-[…]
-$H^{k}\Big{(}\overline{X}(x),\mathcal{F}\Big{)}\overset{\sim}{\longrightarrow}{\rm Ind}_{P_{I}(\mathbb{F}_{q})}^{G(\mathbb{F}_{q})}H^{k}\Big{(}\overline{X}_{L_{I}}(x),\left.\mathcal{F}\right|_{\overline{X}_{L_{I}}(x)}\Big{)},$
-where $\mathcal{F}$ is the structure sheaf (resp. canonical sheaf) of ${\overline{X}(x)}$.
-Proof.
-We give the proof for the structure sheaf. The case for the canonical sheaf follows from Serre duality for Deligne–Mumford stacks [Nir, Corollary 2.10].
-By 3.7, we may assume ${\overline{X}^{S^{*}}(w)}$ has the same cohomology as an element in $B^{+}$ (3.5).
-In other words, let $u^{\prime}\in B^{+}$ be the image of $w$ under $S^{*}\twoheadrightarrow B^{+}$ and let $u\in S^{*}$ be a lift of $u^{\prime}$ with the same expression as $u^{\prime}$, then we have
-$H^{k}\Big{(}{\overline{X}^{S^{*}}(u)},\mathcal{O}_{{\overline{X}^{S^{*}}(u)}}\Big{)}\overset{\sim}{\longrightarrow}H^{k}\Big{(}{\overline{X}^{S^{*}}(w)},\mathcal{O}_{{\overline{X}^{S^{*}}(w)}}\Big{)}$
-for all $k\geq 0$.
-Using the description of the submonoid of reduced elements in $B^{+}$ [GP, Remark 4.1.1, Exercise 4.1], we know that the criterion for $u$ to come from a reduced expression in the Weyl group is that $u$ has no subexpression of the form $ss$ for all $s\in S$.
-It follows from Corollary 4.10 that there exists $v\in S^{*}$ such that $v$ has no subexpression of the form $ss$ for all $s\in S$ and for all $k\geq 0$,
-$H^{k}\Big{(}{\overline{X}^{S^{*}}(v)},\mathcal{O}_{{\overline{X}^{S^{*}}(v)}}\Big{)}\overset{\sim}{\longrightarrow}H^{k}\Big{(}{\overline{X}^{S^{*}}(w)},\mathcal{O}_{{\overline{X}^{S^{*}}(w)}}\Big{)}.$
-In other words, we may assume $v\in S^{*}$ gives a reduced expression of $\alpha(v)$ in $W$.
-By [GKP, Theorem 2.6], there exists $v^{\prime}\in W$ with $\alpha(v)\to v^{\prime}$ and $\ell(v^{\prime})\leq\ell(\alpha(v))$.
-If $\alpha(v)$ is already the $F$-minimal length element in its conjugacy class, then there is nothing to prove.
-In general, $\alpha(v)\to v^{\prime}$ gives $v_{1}^{\prime},\ldots,v_{m}^{\prime}\in W,s_{1},\ldots s_{m-1}\in S$ such that $v_{1}^{\prime}=\alpha(v),v_{m}^{\prime}=v^{\prime}$ and $v_{i}^{\prime}=s_{i-1}v_{i-1}^{\prime}F(s_{i-1})$ for all $i=2,\ldots,m$.
-Let $j\in\{1,\ldots,m\}$ be the smallest integer such that $\ell(v_{j})<\ell(v_{j-1})$.
-For all $i=1,\ldots j$, let $v_{i}\in S^{*}$ be the canonical lift of the corresponding reduced expression of $v_{i}^{\prime}$.
-It follows from 4.9 that ${\overline{X}^{S^{*}}(v_{i})}$ all have the same cohomology for the structure sheaf as ${\overline{X}^{S^{*}}(v)}$ for $i=2,\ldots,j-1$.
-On the other hand, we have $v_{j-1}=s_{j-1}v_{j}F(s_{j-1})$ in $S^{*}$.
-Via 4.6 we know that ${\overline{X}^{S^{*}}(v)}$ has the same cohomology as ${\overline{X}^{S^{*}}(s_{j-1}v_{j})}$ and ${\overline{X}^{S^{*}}(v_{j}F(s_{j-1}))}$.
-In this case, we reset $v$ to be $s_{j-1}v_{j}$.
-This process will only terminate and yields a minimal length element $x$ cf.
-loc.
-cit..
-Any two minimal length elements $x$, $x^{\prime}$ in an $F$-conjugacy class $C$ are conjugate by $F$-cyclic shifts by [GKP, Theorem 2.6] and the corresponding cohomology of ${\overline{X}^{S^{*}}(x)}$ and ${\overline{X}^{S^{*}}(x^{\prime})}$ for the structure sheaf (resp.
-canonical sheaf) are isomorphic by 4.9.
-For the last assertion, there are $G^{F}$-equivariant morphisms (cf. [DMR, Proposition 2.3.8])
-${\overline{X}^{S^{*}}(x)}\overset{\sim}{\longrightarrow}G^{F}/U_{I}^{F}\times^{L_{I}^{F}}{\overline{X}^{S^{*}}_{L_{I}}(x)}\overset{\pi_{I}}{\longrightarrow}G^{F}/P_{I}^{F},$
-where $P_{I}=\bigcup_{v\in W_{I}}{B}v{B}$, $U_{I}$ is the unipotent radical of $P_{I}$, and $L_{I}$ is the Levi subgroup of $G$ satisfying $P_{I}=L_{I}\ltimes U_{I}$.
-Since $G^{F}$ acts on the fibres of $\pi_{I}$ transitively, we have an $G^{F}$-equivariant isomorphisms
-$H^{k}\Big{(}{\overline{X}^{S^{*}}(x)},\mathcal{O}_{{\overline{X}^{S^{*}}(x)}}\Big{)}\overset{\sim}{\longrightarrow}{\rm Ind}_{P_{I}(\mathbb{F}_{q})}^{G(\mathbb{F}_{q})}H^{k}\Big{(}\overline{X}^{S^{*}}_{L_{I}}(x),\mathcal{O}_{\overline{X}^{S^{*}}_{L_{I}}(x)}\Big{)}.$
-for all $k\geq 0$. ∎
-Remark 5.2.
-With the same notations as above, it follows from the proof of the theorem that ${\rm supp}_{F}(x)={\rm supp}_{F}(w)$ for all $x\in C_{\rm min}$.
-Note that $x$ is of full $F$-support in $L_{I}$, so ${\overline{X}(x)}$ is irreducible by 2.12.
-Corollary 5.3.
+Corollary 5.1.2.
 Assume $F(S)=S$.
 Let $v_{1},\ldots,v_{r}\in W$.
-There exists an $F$-conjugacy class $C$ in $W$ such that for any $x\in C_{\rm min}$ and any reduced expression of $x$, there are $G^{F}$-equivariant isomorphisms
-$H^{k}\Big{(}X(\underline{v_{1}},\ldots,\underline{v_{r}}),\mathcal{O}_{X(\underline{v_{1}},\ldots,\underline{v_{r}})}\Big{)}\overset{\sim}{\longrightarrow}H^{k}\Big{(}\overline{X}(x),\mathcal{O}_{\overline{X}(x)}\Big{)}$
+There exists an $F$-conjugacy class $C$ in $W$ and a minimal length element $x_{0}\in C$ such that for any reduced expression $x\in S^{*}$ of $x_{0}$, there are $G^{F}$-equivariant isomorphisms
+$H^{i}\Big{(}X(\underline{v_{1}},\ldots,\underline{v_{r}}),\mathcal{O}_{X(\underline{v_{1}},\ldots,\underline{v_{r}})}\Big{)}\overset{\sim}{\longrightarrow}H^{i}\Big{(}{\overline{X}^{S^{*}}(x)},\mathcal{O}_{{\overline{X}^{S^{*}}(x)}}\Big{)}$
 and
-$H^{k}\Big{(}X(\underline{v_{1}},\ldots,\underline{v_{r}}),\omega_{X(\underline{v_{1}},\ldots,\underline{v_{r}})}\Big{)}\overset{\sim}{\longrightarrow}H^{k}\Big{(}\overline{X}(x),\omega_{\overline{X}(x)}\Big{)},$
-[…]
-$H^{k}\Big{(}\overline{X}(x),\mathcal{F}\Big{)}\overset{\sim}{\longrightarrow}{\rm Ind}_{P_{I}(\mathbb{F}_{q})}^{G(\mathbb{F}_{q})}H^{k}\Big{(}\overline{X}_{L_{I}}(x),\left.\mathcal{F}\right|_{\overline{X}_{L_{I}}(x)}\Big{)},$
-where $\mathcal{F}$ is the structure sheaf (resp. canonical sheaf) of $\overline{X}(x)$.
+$H^{i}\Big{(}X(\underline{v_{1}},\ldots,\underline{v_{r}}),\omega_{X(\underline{v_{1}},\ldots,\underline{v_{r}})}\Big{)}\overset{\sim}{\longrightarrow}H^{i}\Big{(}{\overline{X}^{S^{*}}(x)},\omega_{{\overline{X}^{S^{*}}(x)}}\Big{)}$
+for all $i\geq 0$.
+If $C$ is moreover an elliptic conjugacy class in $W$, then the isomorphisms above hold if we replace $x_{0}$ by any other minimal length element in $C$.
 Proof.
 Fix reduced expressions for $v_{1},\ldots,v_{r}\in W$, and take lifts $v_{1}^{\prime},\ldots,v_{r}^{\prime}\in S^{*}$ corresponding to these reduced expressions.
-By 5.1, it suffices to show that the cohomology of $X(\underline{v_{1}},\ldots,\underline{v_{r}})$ for the structure sheaf is given by the cohomology of ${\overline{X}^{S^{*}}(v_{1}^{\prime}\cdots v_{r}^{\prime})}$.
-Via projection, there is a proper birational morphism $f:{\overline{X}^{S^{*}}(v_{1}^{\prime}\cdots v_{r}^{\prime})}\to X(\underline{v_{1}},\ldots,\underline{v_{r}})$ of projective and strongly $F$-regular schemes (2.13).
+By Theorem 5.1.1, it suffices to show that the cohomology of $X(\underline{v_{1}},\ldots,\underline{v_{r}})$ for the structure sheaf is given by the cohomology of ${\overline{X}^{S^{*}}(v_{1}^{\prime}\cdots v_{r}^{\prime})}$.
+Via projection, there is a proper birational morphism $f:{\overline{X}^{S^{*}}(v_{1}^{\prime}\cdots v_{r}^{\prime})}\to X(\underline{v_{1}},\ldots,\underline{v_{r}})$ of projective and strongly $F_{p}$-regular schemes (Proposition 2.4.1).
 Thus there is a $G^{F}$-equivariant isomorphism
-$\phi_{*}\mathcal{O}_{{\overline{X}^{S^{*}}(v_{1}^{\prime}\cdots v_{r}^{\prime})}}\overset{\sim}{\longrightarrow}\mathcal{O}_{X(\underline{v_{1}},\ldots,\underline{v_{r}})},$
-by the proof of Zariski’s main theorem in [Har, Corollary 11.4].
-The definition of local rings being strongly $F$-regular implies that they are $F$-rational.
-Smith has shown that $F$-rational implies pseudo-rational [Smi, Theorem 3.1].
-Moreover, recall that compactified Deligne–Lusztig varieties are normal and Cohen–Macaulay (2.13).
-By [Kov, Theorem 8.13], the higher direct image $R^{i}\phi_{*}\mathcal{O}_{{\overline{X}^{S^{*}}(v_{1}^{\prime}\cdots v_{r}^{\prime})}}$ for $i>0$ vanish.
+$f_{*}\mathcal{O}_{{\overline{X}^{S^{*}}(v_{1}^{\prime}\cdots v_{r}^{\prime})}}\overset{\sim}{\longrightarrow}\mathcal{O}_{X(\underline{v_{1}},\ldots,\underline{v_{r}})},$
+by the proof of Zariski’s main theorem in [Hartshorne, Corollary III.11.4].
+The definition of local rings being strongly $F_{p}$-regular implies that they are $F_{p}$-rational.
+Smith has shown that $F_{p}$-rational implies pseudo-rational [Ksmith97, Theorem 3.1].
+Moreover, recall that compactified Deligne–Lusztig varieties are normal and Cohen–Macaulay (Proposition 2.4.1).
+By [Kovacs, Theorem 8.13], the higher direct image $R^{j}f_{*}\mathcal{O}_{{\overline{X}^{S^{*}}(v_{1}^{\prime}\cdots v_{r}^{\prime})}}$ for $j>0$ vanish.
 Functoriality of the associated Leray spectral sequence yields $G^{F}$-equivariant isomorphisms on cohomology groups
-$H^{k}\Big{(}X(\underline{v_{1}},\ldots,\underline{v_{r}}),\mathcal{O}_{X(\underline{v_{1}},\ldots,\underline{v_{r}})}\Big{)}\overset{\sim}{\longrightarrow}H^{k}\Big{(}{\overline{X}^{S^{*}}(v_{1}^{\prime}\cdots v_{r}^{\prime})},\mathcal{O}_{{\overline{X}^{S^{*}}(v_{1}^{\prime}\cdots v_{r}^{\prime})}}\Big{)},$
-for all $k\geq 0$. The $G^{F}$-equivariant isomorphisms for the cohomology of the canonical sheaf follow from Serre duality for Deligne–Mumford stacks [Nir, Corollary 2.10].
-The following proposition shows that taking the image of a compactified Deligne–Lusztig variety in a partial flag variety does not affect its cohomology for the structure sheaf and canonical sheaf.
-In particular, Corollary 5.3 applies.
-Proposition 5.4.
-Assume $F(S)=S$. Let $P$ be a parabolic subgroup of $G$ containing ${B}$ and denote $\pi_{P}:G/B\to G/P$ the projection map. Then there are $G^{F}$-equivariant isomorphisms
-$H^{k}\Big{(}X(\underline{v}),\mathcal{F}_{X(\underline{v})}\Big{)}\overset{\sim}{\longrightarrow}H^{k}\Big{(}\pi_{P}(X(\underline{v})),\mathcal{F}_{\pi_{P}(X(\underline{v}))}\Big{)},$
-for all $k\geq 0$, where $\mathcal{F}$ is the structure sheaf (resp. canonical sheaf).
-Proof.
-The map $\pi_{P}:G/B\to G/P$ is Zariski locally trivial with fibres isomorphic to $P/{B}\cong L/(L\cap{B})$, where $L$ is the Levi subgroup of $P$.
-The induced map $\left.\pi_{P}\right|_{X(\underline{v})}:X(\underline{v})\to\pi_{P}(X(\underline{v}))$ is proper and Zariski locally trivial with fibres isomorphic to $L/(L\cap{B})$.
-The Levi subgroup $L$ is a reductive group scheme over $k$ with finitely many connected components.
-Without loss of generality, we assume that $L$ is connected.
-Moreover, $L/(L\cap{B})$ is a flag variety over $k$, so its cohomology for the structure sheaf is given by Kempf’s vanishing theorem: $H^{j}(L/(L\cap{B}),\mathcal{O}_{L/(L\cap{B})})=0$ for $j>0$ and $H^{0}(L/(L\cap{B}),\mathcal{O}_{L/(L\cap{B})})=k$.
-Hence the proof of 4.6 applies.
-∎
+$H^{i}\Big{(}X(\underline{v_{1}},\ldots,\underline{v_{r}}),\mathcal{O}_{X(\underline{v_{1}},\ldots,\underline{v_{r}})}\Big{)}\overset{\sim}{\longrightarrow}H^{i}\Big{(}{\overline{X}^{S^{*}}(v_{1}^{\prime}\cdots v_{r}^{\prime})},\mathcal{O}_{{\overline{X}^{S^{*}}(v_{1}^{\prime}\cdots v_{r}^{\prime})}}\Big{)},$
+for all $i\geq 0$. The $G^{F}$-equivariant isomorphisms for the cohomology of the canonical sheaf follow from Serre duality for Deligne–Mumford stacks [Nironi, Corollary 2.10].
+Remark 5.1.3.
+We use the same notations as in Theorem 5.1.1 and set $I={{\text{supp}}}_{F}(x)$. There are $G^{F}$-equivariant morphisms given by [DMR, Proposition 2.3.8]:
+${\overline{X}^{S^{*}}(x)}\overset{\sim}{\longrightarrow}G^{F}/U_{I}^{F}\times^{L_{I}^{F}}{\overline{X}^{S^{*}}_{L_{I}}(x)}\overset{\pi_{I}}{\longrightarrow}G^{F}/P_{I}^{F},$
+where $P_{I}=BW_{I}B$ is the parabolic subgroup with unipotent radical $U_{I}$ satisfying $P_{I}=L_{I}\ltimes U_{I}$ for a Levi subgroup $L_{I}$ of $G$.
+Since $G^{F}$ acts on the fibres of $\pi_{I}$ transitively, we have a $G^{F}$-equivariant isomorphism
+$H^{i}\Big{(}{\overline{X}^{S^{*}}(x)},\mathcal{O}_{{\overline{X}^{S^{*}}(x)}}\Big{)}\overset{\sim}{\longrightarrow}{\rm Ind}_{P_{I}^{F}}^{G^{F}}H^{i}\Big{(}\overline{X}^{S^{*}}_{L_{I}}(x),\mathcal{O}_{\overline{X}^{S^{*}}_{L_{I}}(x)}\Big{)}.$
+for all $i\geq 0$.
 5.2. Suzuki and Ree groups
-In the classification of finite reductive groups, the Frobenius endomorphism does not necessarily fix the length of elements of $S$ for groups of type ${}^{2}B_{2}$, ${}^{2}F_{4}$, and ${}^{2}G_{2}$.
-Analogous to [DL, §11], we work with groups of type ${}^{2}B_{2}$, ${}^{2}F_{4}$ (resp.
-${}^{2}G_{2}$) assuming $k$ is of characteristic $2$ (resp.
-characteristic $3$).
-In this situation, let $q$ be an odd power of $\sqrt{2}$ (resp.
-$\sqrt{3}$) and let $G$ be defined over $\mathbb{F}_{q^{2}}$.
-Let $F^{\prime}:G\to G$ be a Frobenius root (cf.
-[DM, §4.2]) such that $F^{\prime 2}$ is the the $q^{2}$-Frobenius on $G$.
-In order to make $F^{\prime}$ fix the length of the simple roots, fix a $F^{\prime}$-stable maximal torus $T^{*}\subseteq G$ and a Borel ${B}$ containing $T^{*}$.
-A Frobenius root shares all the properties of the Frobenius morphism we used in our proofs.
-In particular, it is étale and its corresponding Lang map is surjective.
-A systematic referece for Frobenius roots can be found in loc.
-cit.
-Therefore, once we fix the datum above, 5.1 and Corollary 5.3 readily yield $G^{F^{\prime}}$-equivariant isomorphism of cohomology groups of compactified Deligne–Lustig varieties.
-Note that the group $G^{F^{\prime}}$ for type ${}^{2}B_{2}$ is known as the Suzuki group and $G^{F^{\prime}}$ is known as the Ree groups for type ${}^{2}F_{4}$ and ${}^{2}G_{2}$.
+We recall here the assumptions needed on $d$ and $q$ for which $(G,F)$ of type ${}^{2}B_{2}$, ${}^{2}F_{4}$, and ${}^{2}G_{2}$ can arise with $F(S)=S$.
+In addition to the assumptions Section 1.1, let $d=2$ and let $q$ be an odd power of $\sqrt{2}$ (resp.
+$\sqrt{2}$, $\sqrt{3}$).
+Let $G$ be of type $B_{2}$ (resp.
+$F_{4}$, $G_{2}$).
+Then $F:G\to G$ as constructed in Section 1.1 satisfies $F(S)=S$ and $F^{2}$ is the $q^{2}$-Frobenius [DigneMichelBook, §4.3.7].
+The group of fixed points $G^{F}$ is the Suzuki group ${}^{2}B_{2}(q)$ (resp.
+Ree groups ${}^{2}F_{4}(q)$, and ${}^{2}G_{2}(q)$).
+Acknowledgements
+The author thanks Jochen Heinloth and Georg Linden for their comments on an earlier draft of this paper. Thanks to Ulrich Görtz for helpful conversations.
+Parts of the work are carried out when the author was a member of the trimester program "The Arithmetic of the Langlands Program" at the Hausdorff Research Institute for Mathematics in Bonn and when the author was a member of the Research Training Group 2553 in Essen, both funded by the DFG (Deutsche Forschungsgemeinschaft).
+[…]
+Fakultät für Mathematik, Universität Duisburg-Essen, 45127 Essen, Germany
+E-mail address:

### v3

@@
 Abstract.
-For connected reductive groups together with a Frobenius root $F$ that preserves the set of simple reflections, we show that for every tuple $(w_{1},\ldots,w_{r})$ of Weyl group elements there exists a minimal length element in an $F$-conjugacy class in $W$ such that the cohomology of the structure sheaf and the canonical sheaf for the associated compactified Deligne–Lusztig varieties are isomorphic respectively.
+For connected reductive groups together with a Frobenius root $F$, we show that the cohomology of the structure sheaf and respectively the canonical sheaf for compactified Deligne–Lusztig varieties associated to an element in the free monoid generated by the simple reflections is isomorphic to that of a minimal length element in an $F$-conjugacy class in the Weyl group.
 Key words and phrases:
@@
 1.1. Background
-Let $q$ be a power of a prime $p$ and let $k_{0}=\mathbb{F}_{q^{d}}$ for a positive integer $d$.
-Let $k$ be an algebraic closure of $\mathbb{F}_{p}$ containing $k_{0}$.
-Let $G_{0}$ be a connected reductive $k_{0}$-group with $q^{d}$-Frobenius $F_{0}:G_{0}\to G_{0}$.
-Let $G$ be the base change of $G_{0}$ to $k$ and let $F:G\to G$ be an endomorphism such that $F^{d}=F_{0}\otimes_{k_{0}}{\rm Id}$.
+Let $q$ be a power of a prime $p$, $k_{0}=\mathbb{F}_{q^{d}}$ for a positive integer $d$, and $k$ an algebraic closure of $\mathbb{F}_{p}$ containing $k_{0}$.
+Let $G_{0}$ be a connected reductive $k_{0}$-group with $q^{d}$-Frobenius $F_{0}:G_{0}\to G_{0}$, and let $G$ be the base change of $G_{0}$ to $k$ together with an endomorphism $F:G\to G$ such that $F^{d}=F_{0}\otimes_{k_{0}}{\rm Id}$.
 The group of fixed points $G^{F}$ is known as a finite group of Lie type.
+Examples of such groups include ${\rm GL}_{n}(\mathbb{F}_{q})$ and the subgroup of unitary matrices in ${\rm GL}_{n}(\mathbb{F}_{q^{2}})$.
 Let $T\subset B$ be an $F$-stable maximal torus and an $F$-stable Borel subgroup of $G$.
@@
 For any $w\in W$, the associated Deligne–Lusztig variety is denoted by $X(w)$.
-Throughout the introduction, we fix a reduced expression of $w$ and denote by $\overline{X}(w)$ the corresponding smooth compactification of $X(w)$.
-In [DL], Deligne and Lusztig considered the $\ell$-adic cohomology of $X(w)$ and certain Galois coverings of $X(w)$ to study the irreducible $G^{F}$-representations over fields of characteristic zero.
+Throughout the introduction, we fix a reduced expression of $w$ and denote by $\overline{X}(w)$ the corresponding smooth compactification of $X(w)$ (Section˜2.1).
+In [7], Deligne and Lusztig considered the $\ell$-adic cohomology of $X(w)$ and certain Galois coverings of $X(w)$ to study the irreducible $G^{F}$-representations over fields of characteristic zero.
 The de Rham and mod $p$ cohomology of $X(w)$ and $\overline{X}(w)$ are less understood than their $\ell$-adic counterpart.
-However, for $G^{F}={\rm GL}_{2}(\mathbb{F}_{q})$ and $w$ the Coxeter element, the de Rham cohomology of $\overline{X}(w)$ is not mysterious because $\overline{X}(w)$ is isomorphic to the projective line.
+However, for ${\rm GL}_{2}(\mathbb{F}_{q})$ and $w$ the Coxeter element, the de Rham cohomology of $\overline{X}(w)$ is not mysterious because $\overline{X}(w)$ is isomorphic to the projective line.
 The simplicity in the situation of $\overline{X}(w)$ for ${\rm GL}_{2}(\mathbb{F}_{q})$ does not carry over to higher dimensions.
@@
 Thus the cohomology groups of $\overline{X}(w)$ cannot be computed in the same way.
-In [Wang22], we described the cohomology of the structure sheaf of $\overline{X}(w)$ for any $w\in W$ in the case of ${\rm GL}_{n}(\mathbb{F}_{q})$.
+In [26], we described the cohomology of the structure sheaf of $\overline{X}(w)$ for any $w\in W$ in the case of ${\rm GL}_{n}(\mathbb{F}_{q})$.
 Our main theorem showed that the cohomology of the structure sheaf of $\overline{X}(w)$ is always isomorphic to that of $\overline{X}(c)$, where $c$ is a Coxeter element in a parabolic subgroup of $W$ with ${\text{supp}}(c)={\text{supp}}(w):=\{s\in S\ |\ s\leq w\}$.
@@
 1.2. Main results
-The aim of this paper is to extend the main theorem of [Wang22] to all pairs $(G,F)$ for which $F(S)=S$.
-The assumption on $F$, for example, is satisfied when $G$ is split over $k_{0}$, or when $G^{F}$ is the group of unitary matrices in ${\rm GL}_{n}(\mathbb{F}_{q^{2}})$.
-If $G^{F}$ is a finite simple group of Lie type arising from a pair $(G,F)$ as in in Section 1.1, then the classification of such groups implies that $F(S)=S$ is satisfied.
-In fact, $F$ would be the Frobenius endomorphism (i.e.
-$d=1$) except for the cases for Suzuki and Ree groups, for which $d=2$ and additional assumptions on $q$ are required (Section 5.2).
-In our setting, we encounter the minimal length elements in the $F$-conjugacy classes of $W$ (cf.
-Section 4.2), which is a special class of elements in $W$ with nice properties [GeckKimPfeifferMinlength, Theorem 2.6], [HeNieminlength, Theorem 3.1].
-For example, if $G^{F}$ is ${\rm GL}_{n}(\mathbb{F}_{q})$ or the group of unitary matrices in ${\rm GL}_{n}(\mathbb{F}_{q^{2}})$, then such elements are precisely the Coxeter elements in parabolic subgroups of $W$.
-We show that the study of the cohomology of the structure sheaf and the canonical sheaf of $\overline{X}(w)$ for any $w\in W$ can always be reduced to that of a minimal length element in an $F$-conjugacy class in $W$.
-Theorem 5.1.1.
-Assume $F(S)=S$ and let $w\in W$. There exists an $F$-conjugacy class $C$ in $W$ and a minimal length element $x_{0}\in C$ such that for all $i\geq 0$, there are $G^{F}$-equivariant isomorphisms
+The aim of this paper is to extend the main theorem of [26] to all pairs $(G,F)$ as in §1.1.
+After fixing a suitable pair $T\subset B$ of $F$-stable maximal torus and Borel subgroup, we may assume that $F(S)=S$ without loss of generality [9, §4.2].
+In fact, when $G$ is semisimple, $F$ would be the Frobenius endomorphism (i.e.
+$d=1$) except for the cases for Suzuki and Ree groups, for which $d=2$ and additional assumptions on $q$ are required (Section˜5.2).
+In our setting, we encounter the minimal length elements in the $F$-conjugacy classes of $W$ (Section˜4.2), which are special elements in $W$ with nice properties [10, Theorem 2.6], [19, Theorem 3.1].
+When $G^{F}$ is ${\rm GL}_{n}(\mathbb{F}_{q})$, they are precisely the Coxeter elements in parabolic subgroups of $W$.
+In general, Coxeter elements of $W$ (in the sense of Remark˜2.3.2) are examples of minimal length elements in $F$-conjugacy classes.
+Theorem˜5.1.1.
+For every $w\in W$ there exists an $F$-conjugacy class $C$ in $W$ and a minimal length element $x_{0}\in C$ such that for all $i\geq 0$, there are $G^{F}$-equivariant isomorphisms
 $H^{i}\left({\overline{X}(w)},\mathcal{F}_{{\overline{X}(w)}}\right)\overset{\sim}{\longrightarrow}H^{i}\left({\overline{X}(x_{0})},\mathcal{F}_{{\overline{X}(x_{0})}}\right),$
@@
 Note that the $F$-conjugacy class $C$ in the theorem may not contain $w$.
-However, for any given $w$, one can determine $x_{0}$ and $C$ explicitly by following our recipe given in the proof of Theorem 5.1.1.
+However, for any given $w$, one can determine $x_{0}$ and $C$ explicitly by following our recipe given in the proof of Theorem˜5.1.1.
 We will also show that the cohomology groups in the theorem do not depend on the choice of reduced expressions of $w$ and $x_{0}$ and their corresponding smooth compactifications.
@@
 To avoid this ambiguity, we work with compactified Deligne–Lusztig varieties $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ associated to a tuple of Weyl group elements, cf.
-Definition 2.1.1, [DMR, Definition 2.3.2].
+Definition 2.1.1, [8, Definition 2.3.2].
 This notion makes the construction of compactifications of Deligne–Lusztig varieties more straightforward and uniform.
@@
 Setting $w_{1},\ldots,w_{r}\in S$ such that $w_{1}\cdots w_{r}\in W$ is a reduced expression yields the smooth compactification $X(\underline{w_{1}},\ldots,\underline{w_{r}})=\overline{X}(w_{1}\cdots w_{r})$ defined by Deligne and Lusztig.
-In this generality, we give a connectedness criterion for compactified Deligne–Lusztig varieties $X(\underline{w_{1}},\ldots,\underline{w_{r}})$, generalising a result of Bonnafé and Rouquier [BonnafeRouquierOntheirred].
-We also show that they have $F_{p}$-rational singularities (Section 2.4) and thus are normal, Cohen–Macaulay, and pseudo-rational.
-Furthermore, we provide a list of compactifications $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ of Deligne–Lusztig varieties which are smooth.
+In this generality, we give an irreducibility criterion for compactified Deligne–Lusztig varieties $X(\underline{w_{1}},\ldots,\underline{w_{r}})$, generalising a result of Bonnafé and Rouquier [3].
+We also show that they have $F_{p}$-rational singularities (Section˜2.4).
 The structure of the paper is as follows.
-We fix our notations for Deligne–Lusztig varieties and compactified Deligne–Lusztig varieties and describe their geometric properties in Section 2.
-In Section 3, we show that $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is smooth when each $w_{i},i=1,\ldots,r,$ comes from a braid relation in the Weyl group $W$, and we establish relations between cohomology groups for the structure sheaf and the canonical sheaf on the associated compactified Deligne–Lusztig varieties.
-Section 4 contains the constructions of morphisms which are $\mathbb{P}^{1}$-fibrations between certain compactified Deligne–Lusztig varieties.
+We fix our notations for Deligne–Lusztig varieties and compactified Deligne–Lusztig varieties and describe their geometric properties in Section˜2.
+In Section˜3, we show that $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is smooth when each $w_{i},i=1,\ldots,r,$ comes from a braid relation in the Weyl group $W$, and we establish relations between cohomology groups for the structure sheaf and the canonical sheaf on the associated compactified Deligne–Lusztig varieties.
+Section˜4 contains the constructions of morphisms which are $\mathbb{P}^{1}$-fibrations between certain compactified Deligne–Lusztig varieties.
 They are associated with $F$-cyclic conjugations in the Weyl group.
@@
 canonical sheaf) between these compactified Deligne–Lusztig varieties.
-Finally, we prove Theorem 5.1.1 and extend the result to arbitrary compactified Deligne–Lusztig varieties of the form $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ in Section 5.
+Finally, we prove Theorem˜5.1.1 in Section˜5.
 2. Compactified Deligne–Lusztig varieties
 We begin with fixing notations related to Deligne–Lusztig varieties associated to a tuple of Weyl group elements.
-This construction has been treated in [DL, §9] and [DMR, §2].
-We review here the definitions and properties that are relevant for our context.
-Moreover, by adapting classical proofs, we give criterions for compactified Deligne–Lusztig varieties to be connected and show that they are $F_{p}$-rational.
+This construction has been treated in [7, §9] and [8, §2].
+We review the definitions and properties that are relevant for our context.
+Moreover, by adapting classical proofs, we give criterions for compactified Deligne–Lusztig varieties to be irreducible and show that they are $F_{p}$-rational.
 2.1. Definitions and basic properties
-Let $G$, $F:G\to G$, $T\subset B$, $S\subset W$ be as in Section 1.1.
+Let $G$, $F:G\to G$, $T\subset B$, $S\subset W$ be as in Section˜1.1.
+As above, we take a suitable pair of $F$-stable $T\subset B$ such that $F(S)=S$ [9, §4.2].
 In the literature, the endomorphism $F$ is referred to as a Frobenius root or a Steinberg endomorphism.
@@
 Recall that by the Bruhat decomposition, the diagonal $G$-action on the product $G/B\times G/B$ yields a bijection between the Weyl group $W$ and $G\backslash\left(G/B\times G/B\right)$.
-The orbit corresponding to $w\in W$ is denoted by ${O(w)}$.
+The orbit corresponding to $w\in W$ is denoted by O(w){O(w)}.
 It can be regarded as a fibre bundle over the flag variety $G/B$ with fibres isomorphic to the Schubert cell $BwB/B$.
@@
 Since $O(\underline{w_{1}},\ldots,\underline{w_{r}})$ is an iterated fibre bundle over $G/B$ with fibres isomorphic to Schubert varieties $\overline{Bw_{i}B}/B$, cf.
-[Jantzen, F.23], it is an integral, separated, normal, Cohen–Macaulay $k$-scheme locally of finite type.
-A proof of normality can be found in [DMR, Lemma 2.2.13].
-The Cohen–Macaulay property behaves well under base change [stacks-project, Lemma 045T], so $O(\underline{w_{1}},\ldots,\underline{w_{r}})$ is Cohen–Macaulay follows from the classical result that Schubert varieties are Cohen–Macaulay (see the introduction of [Hashimoto]).
+[21, F.23], it is an integral, separated, normal, Cohen–Macaulay $k$-scheme locally of finite type.
+A proof of normality can be found in [8, Lemma 2.2.13].
+The Cohen–Macaulay property behaves well under base change [2, Lemma 045T], so $O(\underline{w_{1}},\ldots,\underline{w_{r}})$ is Cohen–Macaulay follows from the classical result that Schubert varieties are Cohen–Macaulay (see the introduction of [18]).
 Denote by $F:G/B\to G/B$ the map induced by the endomorphism $F:G\to G$ and $\Gamma_{F}\subseteq G/B\times G/B$ its graph.
-Since $F^{d}$ is a Frobenius endomorphism, we know that $F$ is surjective, finite ([DigneMichelBook, Proposition 8.1.2]), separated ([GoerWedhAlgGeom1, Proposition 9.13]), and purely inseparable ([EGA1, (3.7.6)]).
-Moreover, the induced map $F:W\to W$ is a group automorphism [DigneMichelBook, Lemma 4.2.21].
+Since $F^{d}$ is a Frobenius endomorphism, we know that $F$ is surjective, finite ([9, Proposition 8.1.2]), separated ([13, Proposition 9.13]), and purely inseparable ([16, (3.7.6)]).
+Moreover, the induced map $F:W\to W$ is a group automorphism [9, Lemma 4.2.21].
 Definition 2.1.1.
@@
 If for all $i=1,\ldots,r$, the associated Schubert variety $\overline{{B}w_{i}{B}}/{B}$ is smooth, then $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is smooth of dimension $\ell(w_{1})+\cdots+\ell(w_{r})$, cf.
-[DMR, Proposition 2.3.5, 2.3.6 (i)].
+[8, Proposition 2.3.5, 2.3.6 (i)].
 In general, compactified Deligne–Lusztig varieties have $F_{p}$-rational singularities, as we will see in Proposition 2.4.1.
@@
 Moreover, we have a stratification
-$X(\underline{w_{1}},\ldots,\underline{w_{r}})=\bigcup_{w_{i}^{\prime}\leq w_{i},\text{ for all }i}X(w_{1}^{\prime},\ldots w_{r}^{\prime}).$
+[…]
 When $s_{i}\in S$ for all $i=1,\ldots,r$, the tuple $(s_{1},\ldots,s_{r})$ can be regarded as an element in the free monoid generated by $S$, which we denote by $S^{*}$.
 For any $s\in S$, the Zariski closure $\overline{BsB}=B\langle s\rangle B$ is the parabolic subgroup corresponding to the subgroup $\langle s\rangle\subset W$, cf.
-[DMR, Corollary 2.2.10], so the compactified Deligne–Lusztig variety $X(\underline{s_{1}},\ldots,\underline{s_{r}})$ is always smooth.
+[8, Corollary 2.2.10], so the compactified Deligne–Lusztig variety $X(\underline{s_{1}},\ldots,\underline{s_{r}})$ is always smooth.
 To simplify the notations, given any $w=s_{1}\cdots s_{m}\in S^{*}$ with $s_{i}\in S,i=1,\ldots m$, we denote $X(s_{1},\ldots,s_{m})$ by $X^{S^{*}}(w)$ and $X(\underline{s_{1}},\ldots,\underline{s_{m}})$ by $\overline{X}^{S^{*}}(w)$.
 The notations $O^{S^{*}}(w)$ and $\overline{O}^{S^{*}}(w)$ for the orbits are defined analogously.
+If $s_{1}\cdots s_{r}\in S^{*}$ is a reduced expression of an element in $w\in W$, then $\overline{X}^{S^{*}}(s_{1}\cdots s_{r})$ recovers the smooth compactification $\overline{X}(w)$ as defined in [7, §9.10].
 Furthermore, denote by $\alpha:S^{*}\to W$ be the natural projection morphism of monoids.
@@
 We consider an alternative description of Deligne–Lusztig varieties via the Lang map.
-Using this description, we show that $G^{F}$ acts transitively on the set of connected components of $X(w_{1},\ldots,w_{r})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}})$.
+Using this description, we show that $G^{F}$ acts transitively on the set of irreducible components of $X(w_{1},\ldots,w_{r})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}})$.
 Denote by $G^{r}$ the $r$-fold product $G\times\cdots\times G$ over ${\rm Spec}\ k$.
@@
 Note that $F_{1}$ can be viewed as a Frobenius root for the algebraic group $G^{r}$.
-Let $\mathcal{L}_{1}:G^{r}\to G^{r}$ be the Lang map associated to $F_{1}$ given by $x\mapsto x^{-1}F_{1}(x)$ for $x\in G^{r}$. In other words, we have
+The Lang map $\mathcal{L}_{1}$ associated to $F_{1}$ is the endomorphism $x\mapsto x^{-1}F_{1}(x)$ of $G^{r}$. In other words, we have
 $\mathcal{L}_{1}:(g_{0},\ldots,g_{r-1})\longmapsto(g_{0}^{-1}g_{1},\ldots,g_{r-2}^{-1}g_{r-1},g_{r-1}^{-1}F(g_{0})).$
@@
 $X(\underline{w_{1}},\ldots,\underline{w_{r}})\overset{\sim}{\longrightarrow}\mathcal{L}_{1}^{-1}\left(\overline{{B}w_{1}{B}}\times_{{\rm Spec}\ k}\cdots\times_{{\rm Spec}\ k}\overline{{B}w_{r}{B}}\right)/({B})^{r}.$
-Furthermore, $G^{F}$ acts transitively on the irreducible components of $X(w_{1},\ldots,w_{r})$ (resp.
+Moreover, $G^{F}$ acts transitively on the irreducible components of $X(w_{1},\ldots,w_{r})$ (resp.
 $X(\underline{w_{1}},\ldots,\underline{w_{r}})$).
@@
 In this proof, we denote $\mathcal{B}:=\overline{{B}w_{1}{B}}\times_{{\rm Spec}\ k}\cdots\times_{{\rm Spec}\ k}\overline{{B}w_{r}{B}}$.
-By Lang’s theorem as formulated in [CabanEngue, Theorem 7.1], we deduce that $\mathcal{L}_{1}:G^{r}\to G^{r}$ is surjective and is a Galois cover with Galois group $G^{F}$.
+By Lang’s theorem as formulated in [5, Theorem 7.1], we deduce that $\mathcal{L}_{1}:G^{r}\to G^{r}$ is surjective and is a Galois cover with Galois group $G^{F}$.
 The subscheme $\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)$ of $G^{r}$ is stable under the free $({B})^{r}$-action from the right, so we may take the quotient of $\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)$ with respect to the induced free right $({B})^{r}$-action.
@@
 The isomorphisms in this lemma then follows from verifying the definitions, cf.
-[DL, §1.11].
+[7, §1.11].
 Moreover,
-$G^{F}\backslash(\mathcal{L}_{1}^{-1}(\mathcal{B}))\overset{\sim}{\longrightarrow}\mathcal{B}$
+$G^{F}\backslash\mathcal{L}_{1}^{-1}(\mathcal{B})\overset{\sim}{\longrightarrow}\mathcal{B}$
 is irreducible.
-Since the quotient map $\pi:\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)\to\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$ is equivariant under the left $G^{F}$-action and surjective, it induces a surjective morphism $G^{F}\backslash(\mathcal{L}_{1}^{-1}(\mathcal{B}))\to G^{F}\backslash\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$.
+Since the quotient map $\pi:\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)\to\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$ is equivariant under the left $G^{F}$-action and surjective, it induces a surjective morphism $G^{F}\backslash\mathcal{L}_{1}^{-1}(\mathcal{B})\to G^{F}\backslash\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$.
 Thus $G^{F}\backslash\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$ is irreducible and $G^{F}$ acts transitively on the set of irreducible components of $\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$.
@@
 The $G^{F}$-action on $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is not free.
-In $G/B$, the point ${B}$ already has nontrivial stabiliser $B^{F}=B(\mathbb{F}_{q})$.
+In $G/B$, the point B{B} already has nontrivial stabiliser $B^{F}=B(\mathbb{F}_{q})$.
 The point $(B,\ldots,B)$ in $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is also stabilised by $B^{F}=B(\mathbb{F}_{q})$.
-2.3. Criterion for connectedness
-There is a simple criterion for classical Deligne–Lusztig varieties to be connected, proved by Lusztig, Bonnafé–Rouquier [BonnafeRouquierOntheirred], Görtz [GoertzOntheconn], and many others.
-In this section we prove that $X(w_{1},\ldots,w_{r})$ (resp.
-$X(\underline{w_{1}},\ldots,\underline{w_{r}})$) is connected when the tuple $(w_{1},\ldots,w_{r})$ has full $F$-support by modifying the proof of [BonnafeRouquierOntheirred].
+2.3. Criterion for irreducibility
+There is a simple criterion for the irreducibility of classical Deligne–Lusztig varieties, proved by Lusztig, Bonnafé–Rouquier [3], Görtz [15], and many others.
+In this section we show that $X(w_{1},\ldots,w_{r})$ (resp.
+$X(\underline{w_{1}},\ldots,\underline{w_{r}})$) is irreducible when the tuple $(w_{1},\ldots,w_{r})$ has full $F$-support by modifying the proof of [3].
 To begin, we fix the notations for the set of $F$-orbits in $S$ and the $F$-support for elements in $W$ and $S^{*}$.
 Definition 2.3.1.
-The $F$-orbit of $w\in W$ is the set ${{\rm Orb}_{F}(w)}:=\{F^{n}(w)|n\in\mathbb{Z}_{\geq 0}\}$. Under the assumption that $F(S)=S$, we define the set
+The $F$-orbit of $w\in W$ is the set ${{\rm Orb}_{F}(w)}:=\{F^{n}(w)|n\in\mathbb{Z}_{\geq 0}\}$. We define the set
 $S_{F}:=\{{{\rm Orb}_{F}(s)}|s\in S\}$
@@
 The Deligne–Lusztig varieties $X(w_{1},\ldots,w_{r})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ are irreducible if and only if $I=S_{F}$.
-In general, they have $|G^{F}/P_{I}^{F}|$ connected components, where $W_{I}$ is the $F$-stable parabolic subgroup of $W$ generated by $\{s\in W|{{\rm Orb}_{F}(s)}\in I\}$ and $P_{I}={B}W_{I}{B}$.
+In general, they have $|G^{F}/P_{I}^{F}|$ irreducible components, where $W_{I}$ is the $F$-stable parabolic subgroup of $W$ generated by $\{s\in W|{{\rm Orb}_{F}(s)}\in I\}$ and $P_{I}={B}W_{I}{B}$.
 Proof.
@@
 via projection, cf.
-[DMR, §2.3.1].
+[8, §2.3.1].
 It also follows that $X(w_{1},\ldots,w_{r})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ are dense open subschemes of $\overline{X}^{S^{*}}(v)$}.
-Thus it suffices to prove that ${\overline{X}^{S^{*}}(v)}$ is connected if and only if $v$ has full $F$-support.
-The proof showing that ${\overline{X}^{S^{*}}(v)}$ is connected is analogous to the proof of [BonnafeRouquierOntheirred] for classical Deligne–Lusztig varieties.
+Thus it suffices to prove that $\overline{X}^{S^{*}}(v)$} is irreducible if and only if $v$ has full $F$-support.
+The proof showing that $\overline{X}^{S^{*}}(v)$} is irreducible is analogous to the proof of [3] for classical Deligne–Lusztig varieties.
 Suppose $I=S_{F}$ and thus ${{\text{supp}}}_{F}(v)=S_{F}$.
@@
 In particular, $v^{\prime}$ is a reduced expression of $\alpha(v^{\prime})\in W$, which is a Coxeter element.
-This implies that ${\overline{X}^{S^{*}}(v^{\prime})}$ is irreducible [LusztigCoxeterOrbits, Proposition 4.8].
+This implies that $\overline{X}^{S^{*}}(v^{\prime})$} is irreducible [22, Proposition 4.8].
 On the other hand, there is a $G^{F}$-equivariant embedding ${\overline{X}^{S^{*}}(v^{\prime})}\to{\overline{X}^{S^{*}}(v)}$.
@@
 Thus $\overline{X}^{S^{*}}(v)$} is irreducible.
-In this case, its dense open subschemes $X(w_{1},\ldots,w_{r})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ are also irreducible [GoerWedhAlgGeom1, Proposition 3.24].
-Suppose ${{\text{supp}}}_{F}(w_{1},\ldots,w_{r})=I$ and so ${{\text{supp}}}_{F}(v)=I$. By [DMR, Proposition 2.3.8], we know that there is a $G^{F}$-equivariant isomorphism
+In this case, its dense open subschemes $X(w_{1},\ldots,w_{r})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ are also irreducible [13, Proposition 3.24].
+Suppose ${{\text{supp}}}_{F}(w_{1},\ldots,w_{r})=I$ and so ${{\text{supp}}}_{F}(v)=I$. By [8, Proposition 2.3.8], we know that there is a $G^{F}$-equivariant isomorphism
 ${\overline{X}^{S^{*}}(v)}\overset{\sim}{\longrightarrow}G^{F}/U_{I}^{F}\times^{L_{I}^{F}}\overline{X}^{S^{*}}_{L_{I}}(v),$
@@
 whose fibres are isomorphic to $\overline{X}^{S^{*}}_{L_{I}}(v)$.
-In this case, we know that $v$ has full $F$-support with respect to the Weyl group of $L_{I}$, so ${\overline{X}^{S^{*}}(v)}$ has $|G^{F}/P_{I}^{F}|$ connected components.
+In this case, we know that $v$ has full $F$-support with respect to the Weyl group of $L_{I}$, so $\overline{X}^{S^{*}}(v)$} has $|G^{F}/P_{I}^{F}|$ irreducible components.
 ∎
@@
 We denote by $F_{p}$ the (absolute) $p$-Frobenius to distinguish our notations from the notations coming from the theory of $F_{p}$-singularities in commutative algebra and algebraic geometry.
-In [Wang22, Proposition A.6], we showed that some compactified Deligne–Lusztig varieties are strongly $F_{p}$-regular.
+In [26, Proposition A.6], we showed that some compactified Deligne–Lusztig varieties are strongly $F_{p}$-regular.
 This property generalises to arbitrary compactified Deligne–Lusztig varieties.
 As in loc.
-cit., we use a similar strategy as in the proof of [Lusztigunipotentch, Lemma 4.3], which uses faithfully flat descent of the smoothness property along the Lang map.
+cit., we use a similar strategy as in the proof of [23, Lemma 4.3], which uses faithfully flat descent of the smoothness property along the Lang map.
 Proposition 2.4.1.
@@
 Proof.
-When $\overline{Bw_{i}B}$ is smooth for all $i$, this is proved by faithfully flat descent of smoothness in [Lusztigunipotentch, Lemma 4.3] and [DMR, Proposition 2.3.5], cf.
-[GoerWedhAlgGeom1, Proposition 14.59].
+When $\overline{Bw_{i}B}$ is smooth for all $i$, this is proved by faithfully flat descent of smoothness in [23, Lemma 4.3] and [8, Proposition 2.3.5], cf.
+[13, Proposition 14.59].
 Analogously, we use faithfully flat descent of strongly $F_{p}$-regularity to show that any compactified Deligne–Lusztig variety is strongly $F_{p}$-regular.
-For each $i=1,\ldots,r$, we know that $\overline{{B}w_{i}{B}}$ is strongly $F_{p}$-regular [BrTh, Cor. 4.1, p. 958] and thus Cohen-Macaulay [BrTh, Cor. 4.2, p. 959]. The product
+For each $i=1,\ldots,r$, we know that $\overline{{B}w_{i}{B}}$ is strongly $F_{p}$-regular [4, Cor. 4.1, p. 958] and thus Cohen-Macaulay [4, Cor. 4.2, p. 959]. The product
 $\overline{{B}w_{1}{B}}\times_{{\rm Spec}\ k}\cdots\times_{{\rm Spec}\ k}\overline{{B}w_{r}{B}}$
-is strongly $F_{p}$-regular by faithfully flat descent along the projection map ${\rm pr}_{i}:\overline{{B}w_{1}{B}}\times_{{\rm Spec}\ k}\cdots\times_{{\rm Spec}\ k}\overline{{B}w_{r}{B}}\to\overline{{B}w_{i}{B}}$ [HoHu, Theorem 5.5.b].
+is strongly $F_{p}$-regular by faithfully flat descent along the projection map ${\rm pr}_{i}:\overline{{B}w_{1}{B}}\times_{{\rm Spec}\ k}\cdots\times_{{\rm Spec}\ k}\overline{{B}w_{r}{B}}\to\overline{{B}w_{i}{B}}$ [20, Theorem 5.5.b].
 Recall that the Lang map $\mathcal{L}_{1}$ is étale, so its closed fibres are smooth and thus strongly $F_{p}$-regular, cf.
-[KSmith, Ch 3, Remark 1.6, Theorem 4.4(a)].
-We may now apply [Aber, Theorem 3.6] and conclude that
+[25, Ch 3, Remark 1.6, Theorem 4.4(a)].
+We may now apply [1, Theorem 3.6] and conclude that
 $\mathcal{L}_{1}^{-1}\left(\overline{{B}w_{1}{B}}\times_{{\rm Spec}\ k}\cdots\times_{{\rm Spec}\ k}\overline{{B}w_{r}{B}}\right)$
 is strongly $F_{p}$-regular.
-Finally, by Lemma 2.2.1 and the faithfully flat descent of strong $F_{p}$-regularity [HoHu, Theorem 5.5], $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is strongly $F_{p}$-regular.
-Strongly $F_{p}$-regular rings are $F_{p}$-rational, so $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is normal and Cohen-Macaulay by [HoHu, Theorem 4.2, 6.27]. ∎
+Finally, by Lemma 2.2.1 and the faithfully flat descent of strong $F_{p}$-regularity [20, Theorem 5.5], $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is strongly $F_{p}$-regular.
+Strongly $F_{p}$-regular rings are $F_{p}$-rational, so $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is normal and Cohen-Macaulay by [20, Theorem 4.2, 6.27]. ∎
 Remark 2.4.2.
-For $w\in W$, the Zariski closure $\overline{X(w)}$ in $G/B$ is not always smooth, but for any $v\in S^{*}$ that gives a reduced expression of $w$, the projection map ${\overline{X}^{S^{*}}(v)}\to\overline{X(w)}$ is a proper birational morphism [DL, Lemma 9.11].
+For $w\in W$, the Zariski closure $\overline{X(w)}$ in $G/B$ is not always smooth, but for any $v\in S^{*}$ that gives a reduced expression of $w$, the projection map ${\overline{X}^{S^{*}}(v)}\to\overline{X(w)}$ is a proper birational morphism [7, Lemma 9.11].
 3. Braid relations in the Weyl group and birational morphisms
@@
 To begin, we recall the relations in the Weyl group of finite reductive groups, cf.
-[GeckPfeiffer].
-All the relations in the group presentation of the Weyl group as a finite Coxeter group $(W,S)$ can be written as
-$(st)^{m_{st}}=1,\ s,t\in S,$ (1)
-where $m_{st}\in\mathbb{Z}$ such that $m_{ss}=1$ and $m_{st}\geq 2$ if $s\neq t$.
-In the case of $s\neq t$, we call (1) a braid relation in the Weyl group $W$.
-We refer to such an equation (1) as a braid relation in $W$.
-Remark 3.1.1.
-If $F(S)=S$, then whenever $(st)^{m_{st}}=1,\ s,t\in S$ is a braid relation in $W$, the equations
-$(ts)^{m_{st}}=1,\hskip 14.22636pt\ (F(s)F(t))^{m_{st}}=1,\hskip 14.22636pt\text{and}\hskip 14.22636pt(F(t)F(s))^{m_{st}}=1$
-also hold.
-For $w=w_{1}w_{2}\in S^{*}$, we call $w_{1}$ a prefix of $w$ and $w_{2}$ a suffix of $w$.
-Definition 3.1.2.
-Let $(st)^{m}=1$, with $\ s,t\in S$ and $s\neq t$, be a braid relation in $W$. We define
-$P(s,t):=st\cdots\in S^{*},$ (2)
-such that ${{\ell^{S^{*}}}(P(s,t))}=m$ and $P(s,t)$ is a prefix of $(st)^{m}$ in $S^{*}$. When there is no ambiguity, we also use $P(s,t)$ to denote its canonical projection $\alpha(P(s,t))\in W$.
+[11].
+All the relations in the group presentation of the Weyl group as a finite Coxeter group $(W,S)$ can be written as $(st)^{m_{st}}=1,s,t\in S,$ where $m_{st}$ is a positive integer such that $m_{ss}=1$ and $m_{st}\geq 2$ if $s\neq t$.
+When $(st)^{m}=1$, with $s,t\in S$ and $s\neq t$ is a relation in $W$, we define
+$P(s,t):=st\cdots\in S^{*},$
+such that ${{\ell^{S^{*}}}(P(s,t))}=m$.
+When there is no ambiguity, we also use $P(s,t)$ to denote its canonical projection $\alpha(P(s,t))\in W$.
+Moreover, the equation $(st)^{m}=1$ can be rewritten as
+$P(s,t)=P(t,s)$ (1)
+with ${{\ell^{S^{*}}}(P(s,t))}={{\ell^{S^{*}}}(P(t,s))}=m$. We call (1) a braid relation for $W$ following the convention of [11, §4].
 Lemma 3.1.1.
 Let $w_{1},\ldots,w_{r}\in W$.
-Suppose that for all $j\in\{1,\ldots,r\}$, we have $w_{j}=P(s,t)$ for some braid relation in $W$.
+Suppose for each $j\in\{1,\ldots,r\}$, there is a braid relation (1) of $W$ such that $w_{j}=P(s,t)=P(t,s)$.
 Then the compactified Deligne–Lusztig variety $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is smooth.
 Proof.
-By Proposition 2.4.1, it suffices to show that for all $j=1,\ldots,r$, the associated shceme $\overline{{B}w_{j}{B}}$ is smooth.
-Consider $w_{j}=P(s,t)$ with $(st)^{m}=1$. Denote $J=\{s,t\}\subseteq S$. The the corresponding subgroup $W_{J}\subseteq W$ generated by $J$ has a group presentation
+By Proposition 2.4.1, it suffices to show that for all $j=1,\ldots,r$, the associated scheme $\overline{{B}w_{j}{B}}$ is smooth.
+Consider $w_{j}=P(s,t)=P(t,s)$ given by $(st)^{m}=1$. Denote $J=\{s,t\}\subseteq S$. The the corresponding subgroup $W_{J}\subseteq W$ generated by $J$ has a group presentation
 $W_{J}=\langle s,t|(st)^{m}=1,s^{2}=1,t^{2}=1\rangle.$
@@
 Since $w_{j}$ is the longest element in $W_{J}$, we know that $\overline{{B}w_{j}{B}}=P_{J}$ and is smooth cf.
-[DigneMichelBook, Proposition 3.2.8].
+[9, Proposition 3.2.8].
 ∎
 3.2. Higher direct images of the structure sheaf for birational morphisms
-Definition 3.2.1.
-For any $v\in S^{*}$ such that $v=P(s,t)$ for a braid relation (1) in $W$, we set $R(s,t):=P(t,s)$.
+Recall that the braid monoid $B^{+}=B^{+}(W,S)$ associated to $W$ is a monoid generated by $S$ with respect to the braid relations (1) for $W$, cf.
+[11, §4].
+There is a natural surjective morphism of monoids $\beta:S^{*}\twoheadrightarrow B^{+}$.
+Note that the projection map $\alpha:S^{*}\to W$ from §2 factors through $\beta$.
 Proposition 3.2.1.
-Let $w_{1},\ldots,w_{r}\in W$ such that $\overline{{B}w_{j}{B}}$ is smooth for all $j=1,\ldots,r$.
-Suppose for $j\in\{1,\ldots,r\}$, the reduced expression of $w_{j}$ is of the form $w_{j}=P(s,t)$ for a braid relation in $W$.
-Then for all $i\geq 0$, there are $G^{F}$-equivariant isomorphisms
-$H^{i}\left(X_{1},\mathcal{O}_{X_{1}}\right)\overset{\sim}{\longrightarrow}H^{i}\left(X_{2},\mathcal{O}_{X_{2}}\right)$
+Let $w,v\in S^{*}$ such that $\beta(w)=\beta(v)$. Then for all $i\geq 0$, there are $G^{F}$-equivariant isomorphisms
+$H^{i}\left({\overline{X}^{S^{*}}(w)},\mathcal{O}_{{\overline{X}^{S^{*}}(w)}}\right)\overset{\sim}{\longrightarrow}H^{i}\left({\overline{X}^{S^{*}}(v)},\mathcal{O}_{{\overline{X}^{S^{*}}(v)}}\right)$
 and
-$H^{i}\left(X_{1},\omega_{X_{1}}\right)\overset{\sim}{\longrightarrow}H^{i}\left(X_{2},\omega_{X_{2}}\right),$
-for $X_{n}$ $(n\in\{1,2\})$ being one of the $k$-schemes ${\overline{X}^{S^{*}}(w_{1}\cdots w_{r})}$, ${\overline{X}^{S^{*}}(w_{1}\cdots R(s,t)\cdots w_{r})}$, and $X(\underline{w_{1}},\ldots,\underline{w_{r}})$, where $R(s,t)$ replaces $w_{j}$.
+$H^{i}\left({\overline{X}^{S^{*}}(w)},\omega_{{\overline{X}^{S^{*}}(w)}}\right)\overset{\sim}{\longrightarrow}H^{i}\left({\overline{X}^{S^{*}}(v)},\omega_{{\overline{X}^{S^{*}}(v)}}\right).$
 Proof.
-By assumption, we have $w_{j}=P(s,t)=R(s,t)$ in $W$, so replacing $w_{j}$ by $R(s,t)$ does not affect the following compactified Deligne–Lusztig variety
-$X(\underline{w_{1}},\ldots,\underline{w_{r}})=X(\underline{w_{1}},\ldots,\underline{R(s,t)},\ldots,\underline{w_{r}}).$
-Replace $w_{j}$ by $R(s,t)$ in the expression $w_{1}\cdots w_{r}\in S^{*}$ and denote the resulting element by $w_{1}\cdots R(s,t)\cdots w_{r}$.
-For $X_{n}={\overline{X}^{S^{*}}(w_{1}\cdots w_{r})}$, ${\overline{X}^{S^{*}}(w_{1}\cdots R(s,t)\cdots w_{r})}$, $X(\underline{w_{1}},\ldots,\underline{w_{r}})$, it follows from the assumptions and Lemma 3.1.3 that $X_{n}$ is smooth and projective.
-Note that we have $I:={\text{supp}}_{F}(w_{1}\cdots w_{r})={\text{supp}}_{F}(w_{1}\cdots R(s,t)\cdots w_{r})={\text{supp}}_{F}(w_{1},\ldots,w_{r})$, and by Proposition 2.3.3 $X_{n}$ has $|G^{F}/P_{I}^{F}|$ irreducible components.
-There are natural projection maps ${\rm pr}_{0,\ell(w_{j})}:\overline{O}^{S^{*}}(w_{j})\to O(\underline{w_{j}})$ and ${\rm pr}_{0,\ell(w_{j})}:\overline{O}^{S^{*}}(R(s,t))\to O(\underline{w_{j}})$ for all $j=1,\ldots r$.
-Similarly, the projection to the 00-th, $\ell(w_{1})$-th, $\ell(w_{1})+\ell(w_{2})$-th, .
+We can rewrite $w=w_{1}\cdots w_{r}$ in $S^{*}$ such that ${{\ell^{S^{*}}}(w_{j})}=\ell(\alpha(w_{j}))$ and $\overline{{B}\alpha(w_{j}){B}}$ is smooth for all $j=1,\ldots,r$.
+Suppose for an $l\in\{1,\ldots,r\}$, we have $w_{l}=P(s,t)$ given by a braid relation $P(s,t)=P(t,s)$ of $W$.
+Define the element $w^{\prime}:=w_{1}\cdots P(t,s)\cdots w_{r}$ in $S^{*}$ by replacing $w_{l}$ by $P(t,s)$.
+We show below that $\overline{X}^{S^{*}}(w)$}, $\overline{X}^{S^{*}}(w^{\prime})$}, and $X(\underline{\alpha(w_{1})},\ldots,\underline{\alpha(w_{r}}))$ have the same cohomology for the structure sheaf and canonical sheaf.
+Since $\beta(w)=\beta(v)$, the statement of the proposition follows from applying this procedure finitiely many times.
+By assumption, we have $\alpha(w_{l})=P(s,t)=P(t,s)$ in $W$, so replacing $w_{l}$ by $P(t,s)$ does not affect the construction of the following compactified Deligne–Lusztig variety
+$X(\underline{\alpha(w_{1})},\ldots,\underline{\alpha(w_{r}}))=X(\underline{\alpha(w_{1})},\ldots,\underline{P(t,s)},\ldots,\underline{\alpha(w_{r}})).$
+For $X_{n}={\overline{X}^{S^{*}}(w)}$, $\overline{X}^{S^{*}}(w^{\prime})$}, $X(\underline{\alpha(w_{1})},\ldots,\underline{\alpha(w_{r}}))$, it follows from the assumptions and Lemma 3.1.1 that $X_{n}$ is smooth and projective.
+Note that we have $I:={\text{supp}}_{F}(w)={\text{supp}}_{F}(w^{\prime})={\text{supp}}_{F}(\alpha(w_{1}),\ldots,\alpha(w_{r}))$, so $X_{n}$ has $|G^{F}/P_{I}^{F}|$ irreducible components by Proposition 2.3.3,.
+There are natural projection maps ${\rm pr}_{0,\ell(w_{j})}:\overline{O}^{S^{*}}(w_{j})\to O(\underline{w_{j}})$ for all $j=1,\ldots r$ and ${\rm pr}_{0,\ell(w_{l})}:\overline{O}^{S^{*}}(P(t,s))\to O(\underline{\alpha(w_{l})})$.
+Similarly, the projection to the 0-th, $\ell(w_{1})$-th, $\ell(w_{1})+\ell(w_{2})$-th, .
 .
 ., and $\ell(w_{1})+\cdots+\ell(w_{r})$-th coordinate yields the vertical maps in the cartesian diagrams
-${{\overline{X}^{S^{*}}(w_{1}\cdots w_{r})}}$${\overline{O}^{S^{*}}(w_{1}\cdots w_{r})}$${X(\underline{w_{1}},\ldots,\underline{w_{r}})}$${O(\underline{w_{1}},\ldots,\underline{w_{r}}),}$$\scriptstyle{f}$
+$\overline{X}^{S^{*}}(w)$$\overline{O}^{S^{*}}(w)$$X(\underline{\alpha(w_{1})},\ldots,\underline{\alpha(w_{r}}))$$O(\underline{\alpha(w_{1})},\ldots,\underline{\alpha(w_{r}})),$$\scriptstyle{f}$
 and
-${{\overline{X}^{S^{*}}(w_{1}\cdots R(s,t)\cdots w_{r})}}$${\overline{O}^{S^{*}}(w_{1}\cdots R(s,t)\cdots w_{r})}$${X(\underline{w_{1}},\ldots,\underline{w_{r}})}$${O(\underline{w_{1}},\ldots,\underline{w_{r}}).}$$\scriptstyle{f^{\prime}}$
+$\overline{X}^{S^{*}}(w^{\prime})$$\overline{O}^{S^{*}}(w^{\prime})$$X(\underline{\alpha(w_{1})},\ldots,\underline{\alpha(w_{r}}))$$O(\underline{\alpha(w_{1})},\ldots,\underline{\alpha(w_{r}})).$$\scriptstyle{f^{\prime}}$
 The maps $f$ and $f^{\prime}$ are $G^{F}$-equivariant proper birational morphisms.
-It follows from the proof of Zariski’s main theorem ([Hartshorne, Corollary 11.4]) that $\phi_{*}\mathcal{O}_{X_{1}}\cong\mathcal{O}_{X_{2}}$, for $\phi=f$ (resp.
-$\phi=f^{\prime}$), $X_{1}={\overline{X}^{S^{*}}(w_{1}\cdots w_{r})}$ (resp.
-$X_{1}={\overline{X}^{S^{*}}(w_{1}\cdots R(s,t)\cdots w_{r})}$) and $X_{2}=X(\underline{w_{1}},\ldots,\underline{w_{r}})$.
-Moreover, applying [CR11, Corollary 3.2.10] yields $R\phi_{*}\mathcal{O}_{X_{1}}\cong\mathcal{O}_{X_{2}}$.
+It follows from the proof of Zariski’s main theorem ([17, Corollary 11.4]) that $\phi_{*}\mathcal{O}_{X_{1}}\cong\mathcal{O}_{X_{2}}$, for $\phi=f$ (resp.
+$\phi=f^{\prime}$), $X_{1}={\overline{X}^{S^{*}}(w)}$ (resp.
+$X_{1}={\overline{X}^{S^{*}}(w^{\prime})}$) and $X_{2}=X(\underline{\alpha(w_{1})},\ldots,\underline{\alpha(w_{r}}))$.
+Moreover, applying [6, Corollary 3.2.10] yields $R\phi_{*}\mathcal{O}_{X_{1}}\cong\mathcal{O}_{X_{2}}$.
 Hence the Leray spectral sequence associated to $\phi$ degenerates.
@@
 $H^{i}\left(X_{1},\mathcal{O}_{X_{1}}\right)\overset{\sim}{\longrightarrow}H^{i}\left(X_{2},\mathcal{O}_{X_{2}}\right),$
-for all $i\geq 0$. The $G^{F}$-equivariant isomorphisms for the cohomology of the canonical sheaf follow from an equivariant version of Serre duality [Nironi, Corollary 2.10]. ∎
+for all $i\geq 0$. The $G^{F}$-equivariant isomorphisms for the cohomology of the canonical sheaf follow from an equivariant version of Serre duality [24, Corollary 2.10]. ∎
 4. $\mathbb{P}^{1}$-fibrations and $F$-conjugacy classes
@@
 Lemma 4.1.1.
-Let $w_{1},\ldots,w_{r},w_{1}^{\prime},\ldots,w_{m}^{\prime}\in W$ and $s\in S$. Then the projection map
+Let $w_{1},\ldots,w_{r},w_{1}^{\prime},\ldots,w_{m}^{\prime}\in W$ and $s\in S$. The projection map
 $\phi:X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{s},\underline{s},\underline{w_{1}^{\prime}},\ldots,\underline{w_{m}^{\prime}})\longrightarrow X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{s},\underline{w_{1}^{\prime}},\ldots,\underline{w_{m}^{\prime}})$
-given by
-$(B_{0},\ldots,B_{r},B_{r+1},B_{r+2},\ldots,B_{r+m+1},FB_{0})\mapsto(B_{0},\ldots,B_{r},B_{r+2},\ldots,B_{r+m+1},FB_{0})$ is a $G^{F}$-equivariant Zariski $\mathbb{P}^{1}$-bundle.
-If $F(s)\in S$, then the $G^{F}$-equivariant morphism of $k$-schemes
+given by $(B_{0},\ldots,B_{r},B_{r+1},B_{r+2},\ldots,B_{r+m+1},FB_{0})\mapsto(B_{0},\ldots,B_{r},B_{r+2},\ldots,B_{r+m+1},FB_{0})$ is a $G^{F}$-equivariant Zariski $\mathbb{P}^{1}$-bundle.
+The $G^{F}$-equivariant morphism of $k$-schemes
 $\pi:X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})\longrightarrow X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)}),$
@@
 Recall that it is a property of the Bruhat decomposition that if $(B_{0},B_{1})\in O(t)$ and $(B_{1},B_{2})\in O(t)$ then $(B_{0},B_{2})\in O(\underline{t})$, cf.
-[DL, §1.2].
+[7, §1.2].
 For $t=s$, we see that $\phi$ is the base change of $O(\underline{t},\underline{t})\to O(\underline{t})$ along the projection map $X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{t},\underline{w_{1}^{\prime}},\ldots,\underline{w_{m}^{\prime}})\to O(\underline{t})$ given by $(B_{0}^{\prime},\ldots,FB_{0}^{\prime})\mapsto(B_{r}^{\prime},B_{r+1}^{\prime})$.
 The first claim follows.
-Assume $F(s)\in S$ and consider $t=F(s)$.
+Consider $t=F(s)\in S$.
 One can verify that $\pi$ is $G^{F}$-equivariant by its definition.
@@
 Lemma 4.1.3.
-Let $w_{1},\ldots,w_{r}\in W$ and $s\in S$ such that $F(s)\in S$ and $\overline{Bw_{n}B}$ is smooth for $n=1,\ldots,r$. The morphisms
+Let $w_{1},\ldots,w_{r}\in W$ and $s\in S$ such that $\overline{Bw_{n}B}$ is smooth for $n=1,\ldots,r$. The morphisms
 $\tau_{2}:X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})},\underline{s})\longrightarrow X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}}),$
@@
 Given $(B_{0},B_{1})\in O(\underline{s})$, we know that $(FB_{0},FB_{1})\in O(F(\underline{s}))$.
-By construction we have $(B_{r+1},FB_{0})\in O(F(\underline{s}))$, so $(B_{r+1},FB_{1})\in O(F(\underline{s}))$ and $(FB_{r+1},F^{2}B_{1})\in O(F(\underline{s}))$ [DL, §1.2].
+By construction we have $(B_{r+1},FB_{0})\in O(F(\underline{s}))$, so $(B_{r+1},FB_{1})\in O(F(\underline{s}))$ and $(FB_{r+1},F^{2}B_{1})\in O(F(\underline{s}))$ [7, §1.2].
 Thus $\tau_{1}$ and $\tau_{2}$ are well-defined.
@@
 The morphism $d_{1}$ also fits into another commutative diagram
-${Z}$${O(\underline{F^{-1}(s)})}$${X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})})}$${G/B,}$$\scriptstyle{r_{2}}$$\scriptstyle{d_{1}}$$\scriptstyle{{\rm pr}_{1}}$$\scriptstyle{{\rm pr}_{0}}$ (4)
+[…]
 where ${\rm pr}_{1}:O(\underline{s})\to G/B$ is given by $(B_{0},B_{1})\mapsto B_{1}$, and $r_{2}$ is given by
@@
 We know $FB_{i}=FB_{i}^{\prime}$ by assumption, so $F(v_{i})=e$ is the neutral element in $W$ and $v_{i}=e$ because $F:W\to W$ is an automorphism, cf.
-[DL, §1.2].
+[7, §1.2].
 It follows that $\iota$ is also injective and an isomorphism.
 Hence $d_{1}$ is a Zariski $\mathbb{P}^{1}$-bundle.
-Since $F:X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})})\to X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}})$ is faithfully flat by our assumption on smoothness [GoerWedhAlgGeom1, Corollary 14.129], we deduce that $\tau_{2}$ is faithfully flat by faithfully flat descent [GoerWedhAlgGeom1, Corollary 14.12, Proposition 14.3(2)].
+Since $F:X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})})\to X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}})$ is faithfully flat by our assumption on smoothness [13, Corollary 14.129], we deduce that $\tau_{2}$ is faithfully flat by faithfully flat descent [13, Corollary 14.12, Proposition 14.3(2)].
 For each affine open (Zariski) covering $\{U_{\alpha}\}_{\alpha}$ of $X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})})$ that trivialises $d_{1}$, we can construct an fppf covering $\{\left.F\right|_{U_{\alpha}}:U_{\alpha}\to X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}})\}_{\alpha}$ of $X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}})$ over which $\tau_{2}$ trivialises as a $\mathbb{P}^{1}$-bundle.
@@
 Proposition 4.1.4.
-Let $w_{1},\ldots,w_{r}\in W$ and $s\in S$ such that $F(s)\in S$ and $\overline{Bw_{n}B}$ is smooth for $n=1,\ldots,r$. There are $G^{F}$-equivariant isomorphisms for all $i\geq 0$,
+Let $w_{1},\ldots,w_{r}\in W$ and $s\in S$ such that $\overline{Bw_{n}B}$ is smooth for $n=1,\ldots,r$. There are $G^{F}$-equivariant isomorphisms for all $i\geq 0$,
 $H^{i}\big(X_{1},\mathcal{O}_{X_{1}}\big)\overset{\sim}{\longrightarrow}H^{i}\big(X_{2},\mathcal{O}_{X_{2}}\big)$
@@
 Since $d_{1}$ is a Zariski $\mathbb{P}^{1}$-bundle, it is proper and faithfully flat.
-Hence $\tau_{1}$ and $\tau_{2}$ are proper and faithfully flat by faithfully flat descent [GoerWedhAlgGeom1, Proposition 14.53].
+Hence $\tau_{1}$ and $\tau_{2}$ are proper and faithfully flat by faithfully flat descent [13, Proposition 14.53].
 Set $f:=\tau_{2}$ (resp.
@@
 Let $x\in X_{2}$ and $y\in X_{2}^{\prime}$ such that $F(y)=x$ as in diagram (2).
-By Lemma 4.1.3, the base change of the fibre $x\times_{X_{2}}X_{1}$ to ${\rm Spec}(\kappa(y))$ is isomorphic to $\mathbb{P}^{1}_{\kappa(y)}$.
-It follows from flat base change that $H^{j}(x\times_{X_{2}}X_{1},\mathcal{O}_{x\times_{X_{2}}X_{1}})=0$ for all $j\geq 0$ [GoerWedhAlgGeom2, Corollary 22.91].
-We conclude by Grauert’s theorem [GoerWedhAlgGeom2, Proposition 23.142] (cf.
-[GoerWedhAlgGeom1, Proposition 3.35]) that $R^{j}f_{*}\mathcal{O}_{X_{1}}=0$ for all $j>0$.
+By Lemma 4.1.3, the base change of the fibre $x\times_{X_{2}}X_{1}$ to $y$ is isomorphic to $\mathbb{P}^{1}_{y}$.
+It follows from flat base change that $H^{j}(x\times_{X_{2}}X_{1},\mathcal{O}_{x\times_{X_{2}}X_{1}})=0$ for all $j\geq 0$ [14, Corollary 22.91].
+We conclude by Grauert’s theorem [14, Proposition 23.142] (cf.
+[13, Proposition 3.35]) that $R^{j}f_{*}\mathcal{O}_{X_{1}}=0$ for all $j>0$.
 Hence the Leray spectral sequence degenerates.
 We have an fppf covering $\mathfrak{U}=\{\gamma_{\alpha}:U_{\alpha}\to X_{2}\}_{\alpha}$ of $X_{2}$ in Lemma 4.1.3, such that the base change $f_{\alpha}:X_{1}\times_{X_{2}}U_{\alpha}\to U_{\alpha}$ of $f$ along $\gamma_{\alpha}$ is a trivial $\mathbb{P}^{1}$-bundle for all $\alpha$.
-Since $U_{\alpha}$ is affine and $\gamma_{\alpha}$ is an affine morphism for all $\alpha$ [GoerWedhAlgGeom1, Proposition 12.3(3)], any simplex $V$ in the Čech nerve of $\mathfrak{U}$ is affine and for any coherent sheaf $\mathcal{F}$ on $X_{2}$, we have $H^{i}(X_{2},\mathcal{F})\overset{\sim}{\longrightarrow}\check{H}^{i}(\mathfrak{U},\mathcal{F})$ for all $i\geq 0$ [Godement, Theorem II.5.4.1, Corollary].
-Note that $f_{*}\mathcal{O}_{X_{1}}$ is a coherent $\mathcal{O}_{X_{2}}$-module because $f$ is proper and cohomology of coherent sheaves is independent of topology [stacks-project, 03DW].
-For each simplex $V$ in the Čech nerve of $\mathfrak{U}$, we have $V\times_{X_{2}}X_{1}\overset{\sim}{\longrightarrow}V\times_{{\rm Spec}\ k}\mathbb{P}^{1}$, so $f_{*}\mathcal{O}_{X_{1}}(V)=\mathcal{O}_{V}(V)$ by flat base change [GoerWedhAlgGeom2, Proposition 22.90].
+Since $U_{\alpha}$ is affine and $\gamma_{\alpha}$ is an affine morphism for all $\alpha$ [13, Proposition 12.3(3)], any simplex $V$ in the Čech nerve of $\mathfrak{U}$ is affine and for any coherent sheaf $\mathcal{F}$ on $X_{2}$, we have $H^{i}(X_{2},\mathcal{F})\overset{\sim}{\longrightarrow}\check{H}^{i}(\mathfrak{U},\mathcal{F})$ for all $i\geq 0$ [12, Theorem II.5.4.1, Corollary].
+Note that $f_{*}\mathcal{O}_{X_{1}}$ is a coherent $\mathcal{O}_{X_{2}}$-module because $f$ is proper and cohomology of coherent sheaves is independent of topology [2, 03DW].
+For each simplex $V$ in the Čech nerve of $\mathfrak{U}$, we have $V\times_{X_{2}}X_{1}\overset{\sim}{\longrightarrow}V\times_{{\rm Spec}\ k}\mathbb{P}^{1}$, so $f_{*}\mathcal{O}_{X_{1}}(V)=\mathcal{O}_{V}(V)$ by flat base change [14, Proposition 22.90].
 This identification is compatible with restriction of sections, so the Čech complexes on $\mathfrak{U}$ for $f_{*}\mathcal{O}_{X_{1}}$ and $\mathcal{O}_{X_{2}}$ are identical and $H^{i}(X_{2},\mathcal{O}_{X_{2}})\overset{\sim}{\longrightarrow}H^{i}(X_{2},f_{*}\mathcal{O}_{X_{1}})$ for all $i\geq 0$.
 Finally, degeneration of the Leray spectral sequence yields $G^{F}$-equivariant isomorphisms
-$H^{i}\left(X_{2},\mathcal{O}_{X_{1}}\right)\overset{\sim}{\longrightarrow}H^{i}\left(X_{1},\mathcal{O}_{X_{1}}\right),$
+$H^{i}\left(X_{2},\mathcal{O}_{X_{2}}\right)\overset{\sim}{\longrightarrow}H^{i}\left(X_{1},\mathcal{O}_{X_{1}}\right),$
 for all $i\geq 0$.
-The $G^{F}$-equivariant isomorphisms for the cohomology of $\omega_{X_{1}}$ and $\omega_{X_{2}}$ follow from Serre duality for Deligne–Mumford stacks [Nironi, Corollary 2.10].
+The $G^{F}$-equivariant isomorphisms for the cohomology of $\omega_{X_{1}}$ and $\omega_{X_{2}}$ follow from Serre duality for Deligne–Mumford stacks [24, Corollary 2.10].
 Corollary 4.1.5.
-Let $s\in S$ and assume $F(s)\in S$. Let $w=sv$ and $w^{\prime}=vF(s)$ be elements of $S^{*}$. There are $G^{F}$-equivariant isomorphisms for all $i\geq 0$,
+Let $s\in S$. Let $w=sv$ and $w^{\prime}=vF(s)$ be elements of $S^{*}$. There are $G^{F}$-equivariant isomorphisms for all $i\geq 0$,
 $H^{i}\Big({\overline{X}^{S^{*}}(w)},\mathcal{O}_{{\overline{X}^{S^{*}}(w)}}\Big)\overset{\sim}{\longrightarrow}H^{i}\Big({\overline{X}^{S^{*}}(w^{\prime})},\mathcal{O}_{{\overline{X}^{S^{*}}(w^{\prime})}}\Big)$
@@
 An element $w\in W$ is of minimal length if it has the smallest Bruhat length in its $F$-conjugacy class.
-Minimal length elements were studied by Geck, Kim, and Pfeiffer [GeckKimPfeifferMinlength], as well as He and Nie [HeNieminlength].
-Under the assumption $F(S)=S$, these two papers showed that for any $w\in W$, there exists a minimal length element $w^{\prime}$ in the $F$-conjugacy class of $w$ such that $w\to_{F}w^{\prime}$ (cf.
+Minimal length elements were studied by Geck, Kim, and Pfeiffer [10], as well as He and Nie [19].
+In our context, the results of these two papers imply that that for any $w\in W$, there exists a minimal length element $w^{\prime}$ in the $F$-conjugacy class of $w$ such that $w\to_{F}w^{\prime}$ (cf.
 Definition 4.2.1).
-In combination with Proposition 4.1.4, this result gives an inductive approach for studying cohomology groups of Deligne–Lusztig varieties.
+In combination with Proposition 4.1.4, we obtain a framework for an inductive approach for studying cohomology groups of Deligne–Lusztig varieties.
 Definition 4.2.1.
@@
 Furthermore, for any pairs of minimal length elements $w$ and $w^{\prime}$ in an elliptic $F$-conjugacy class $C$ in $W$, we have $w\to_{F}w^{\prime}$ and $w^{\prime}\to_{F}w$, cf.
-[HeNieminlength, Corollary 4.4].
+[19, Corollary 4.4].
 The definition of an elliptic $F$-conjugacy class, which we do not recall here, can be found in §4.1 of loc.
@@
 5.1. Main theorem and applications
-Recall that the braid monoid associated to $W$ is a monoid generated by $S$ and all the relations of the form $P(s,t)=P(t,s)$ with $s,t\in S$, $t\neq s$, such that $(st)^{m}=1$ is a braid relation in $W$, cf.
-[GeckPfeiffer, §4].
-We denote the braid monoid by $B^{+}$.
-There is a natural surjective morphism of monoids $\beta:S^{*}\twoheadrightarrow B^{+}$.
 Theorem 5.1.1.
-Assume $F(S)=S$.
 For each $w\in S^{*}$, there exists an $F$-conjugacy class $C$ in $W$ and a minimal length element $x_{0}\in C$ such that for any reduced expression $x\in S^{*}$ of $x_{0}$, there are $G^{F}$-equivariant isomorphisms
@@
 Proof.
-We prove the theorem for the structure sheaf. The results for the canonical sheaf follows from Serre duality for Deligne–Mumford stacks [Nironi, Corollary 2.10].
-Let $w\in S^{*}$. Consider the image $\beta(w)\in B^{+}$ of $w$ under the projection $\beta:S^{*}\twoheadrightarrow B^{+}$ and let $u\in S^{*}$ be another lift of $\beta(w)$, then we have
-$H^{i}\Big{(}{\overline{X}^{S^{*}}(u)},\mathcal{O}_{{\overline{X}^{S^{*}}(u)}}\Big{)}\overset{\sim}{\longrightarrow}H^{i}\Big{(}{\overline{X}^{S^{*}}(w)},\mathcal{O}_{{\overline{X}^{S^{*}}(w)}}\Big{)}$ (5)
-for all $i\geq 0$ by Proposition 3.2.2. Note that this proposition can be applied to any $t_{1}\ldots t_{r}\in S^{*}$ by setting $w_{j}=t_{j}$ for $j=1,\ldots r$.
-Using the description of the submonoid of reduced elements in $B^{+}$ given by Matsumoto’s theorem [GeckPfeiffer, Theorem 1.2.2, Exercise 4.1], we know that the criterion for $u$ to come from a reduced expression in the Weyl group is that $u$ has no subexpression of the form $ss$ for any $s\in S$.
-If $u=u_{1}ssu_{2}$, we apply Lemma 4.1.1 and Proposition 4.1.4 to obtain $u_{1}su_{2}$ such that ${\overline{X}^{S^{*}}(u)}$ and ${\overline{X}^{S^{*}}(u_{1}su_{2})}$ have the same cohomology for the structure sheaf.
-Inductively, this yields an element $v\in S^{*}$ with ${{\ell^{S^{*}}}(v)}\leq{{\ell^{S^{*}}}(u)}$ such that $v$ has no subexpression of the form $ss$ for any $s\in S$ and for all $i\geq 0$,
+We prove the theorem for the structure sheaf. The results for the canonical sheaf follows from Serre duality for Deligne–Mumford stacks [24, Corollary 2.10].
+Let $w\in S^{*}$.
+We showed in Proposition 3.2.1 that the cohomology for the structure sheaf of $\overline{X}^{S^{*}}(w)$} and $\overline{X}^{S^{*}}(w^{\prime})$} are isomorphic for any other lift $w^{\prime}$ of $\beta(w)$ under the projection $\beta:S^{*}\twoheadrightarrow B^{+}$.
+Using the description of the submonoid of reduced elements in $B^{+}$ given by Matsumoto’s theorem [11, Theorem 1.2.2, Exercise 4.1], we know that the criterion for $u\in B^{+}$ to come from a reduced expression in the Weyl group is that $u$ has no subexpression of the form $ss$ for any $s\in S$.
+If $w=w_{1}ssw_{2}$, we apply Lemma 4.1.1 and Proposition 4.1.4 to obtain $w_{1}sw_{2}$ such that $\overline{X}^{S^{*}}(w)$} and $\overline{X}^{S^{*}}(w_{1}sw_{2})$} have the same cohomology for the structure sheaf.
+Inductively, this yields an element $v\in S^{*}$ with ${{\ell^{S^{*}}}(v)}\leq{{\ell^{S^{*}}}(w)}$ such that $v$ has no subexpression of the form $ss$ for any $s\in S$ and for all $i\geq 0$,
 $H^{i}\Big({\overline{X}^{S^{*}}(v)},\mathcal{O}_{{\overline{X}^{S^{*}}(v)}}\Big)\overset{\sim}{\longrightarrow}H^{i}\Big({\overline{X}^{S^{*}}(w)},\mathcal{O}_{{\overline{X}^{S^{*}}(w)}}\Big).$
 In other words, we may assume $v\in S^{*}$ gives a reduced expression of $\alpha(v)$ in $W$.
-By [GeckKimPfeifferMinlength, Theorem 2.6], there exists $v^{\prime}\in W$ which is of minimal length in its $F$-conjugacy class such that $\alpha(v)\to_{F}v^{\prime}$ and $\ell(v^{\prime})\leq\ell(\alpha(v))$.
+By [10, Theorem 2.6], there exists $v^{\prime}\in W$ which is of minimal length in its $F$-conjugacy class such that $\alpha(v)\to_{F}v^{\prime}$ and $\ell(v^{\prime})\leq\ell(\alpha(v))$.
 Suppose $\alpha(v)$ is a minimal length element in its $F$-conjugacy class.
-By Matsumoto’s theorem [GeckPfeiffer, Theorem 1.2.2], there exists a lift $y\in B^{+}$ of $\alpha(v)$ such that any reduced expression $x\in S^{*}$ of $\alpha(v)$ satisfies $\beta(x)=y$.
+By Matsumoto’s theorem [11, Theorem 1.2.2], there exists a lift $y\in B^{+}$ of $\alpha(v)$ such that any reduced expression $x\in S^{*}$ of $\alpha(v)$ satisfies $\beta(x)=y$.
 The statement of the theorem thus follows from Proposition 3.2.1.
@@
 $v_{j-1}=s_{j-1}z_{j-1}$) for some $z_{j-1}\in S^{*}$ (cf.
-[GeckKimPfeifferMinlength, Remark 2.3]).
+[10, Remark 2.3]).
 Thus it follows from Corollary 4.1.5 that $\overline{X}^{S^{*}}(v_{j})$} all have the same cohomology for the structure sheaf as $\overline{X}^{S^{*}}(v)$} for $j=2,\ldots,a-1$.
@@
 This process will terminate and yields an element $x_{0}\in W$ which is of minimal length in its $F$-conjugacy class, cf.
-[GeckKimPfeifferMinlength, Theorem 2.6].
+[10, Theorem 2.6].
 It remains to consider the case when $C$ is an elliptic conjugacy class in $W$.
-By [HeNieminlength, Corollary 4.4], we have $x\to_{F}y$ for any two minimal length elements $x,y\in C$.
+By [19, Corollary 4.4], we have $x\to_{F}y$ for any two minimal length elements $x,y\in C$.
 The corresponding cohomology of $\overline{X}^{S^{*}}(x)$} and $\overline{X}^{S^{*}}(y)$} for the structure sheaf are isomorphic by Corollary 4.1.5.
 ∎
-Corollary 5.1.2.
-Assume $F(S)=S$.
-Let $v_{1},\ldots,v_{r}\in W$.
-There exists an $F$-conjugacy class $C$ in $W$ and a minimal length element $x_{0}\in C$ such that for any reduced expression $x\in S^{*}$ of $x_{0}$, there are $G^{F}$-equivariant isomorphisms
-$H^{i}\Big{(}X(\underline{v_{1}},\ldots,\underline{v_{r}}),\mathcal{O}_{X(\underline{v_{1}},\ldots,\underline{v_{r}})}\Big{)}\overset{\sim}{\longrightarrow}H^{i}\Big{(}{\overline{X}^{S^{*}}(x)},\mathcal{O}_{{\overline{X}^{S^{*}}(x)}}\Big{)}$
-and
-$H^{i}\Big{(}X(\underline{v_{1}},\ldots,\underline{v_{r}}),\omega_{X(\underline{v_{1}},\ldots,\underline{v_{r}})}\Big{)}\overset{\sim}{\longrightarrow}H^{i}\Big{(}{\overline{X}^{S^{*}}(x)},\omega_{{\overline{X}^{S^{*}}(x)}}\Big{)}$
-for all $i\geq 0$.
-If $C$ is moreover an elliptic conjugacy class in $W$, then the isomorphisms above hold if we replace $x_{0}$ by any other minimal length element in $C$.
-Proof.
-Fix reduced expressions for $v_{1},\ldots,v_{r}\in W$, and take lifts $v_{1}^{\prime},\ldots,v_{r}^{\prime}\in S^{*}$ corresponding to these reduced expressions.
-By Theorem 5.1.1, it suffices to show that the cohomology of $X(\underline{v_{1}},\ldots,\underline{v_{r}})$ for the structure sheaf is given by the cohomology of ${\overline{X}^{S^{*}}(v_{1}^{\prime}\cdots v_{r}^{\prime})}$.
-Via projection, there is a proper birational morphism $f:{\overline{X}^{S^{*}}(v_{1}^{\prime}\cdots v_{r}^{\prime})}\to X(\underline{v_{1}},\ldots,\underline{v_{r}})$ of projective and strongly $F_{p}$-regular schemes (Proposition 2.4.1).
-Thus there is a $G^{F}$-equivariant isomorphism
-$f_{*}\mathcal{O}_{{\overline{X}^{S^{*}}(v_{1}^{\prime}\cdots v_{r}^{\prime})}}\overset{\sim}{\longrightarrow}\mathcal{O}_{X(\underline{v_{1}},\ldots,\underline{v_{r}})},$
-by the proof of Zariski’s main theorem in [Hartshorne, Corollary III.11.4].
-The definition of local rings being strongly $F_{p}$-regular implies that they are $F_{p}$-rational.
-Smith has shown that $F_{p}$-rational implies pseudo-rational [Ksmith97, Theorem 3.1].
-Moreover, recall that compactified Deligne–Lusztig varieties are normal and Cohen–Macaulay (Proposition 2.4.1).
-By [Kovacs, Theorem 8.13], the higher direct image $R^{j}f_{*}\mathcal{O}_{{\overline{X}^{S^{*}}(v_{1}^{\prime}\cdots v_{r}^{\prime})}}$ for $j>0$ vanish.
-Functoriality of the associated Leray spectral sequence yields $G^{F}$-equivariant isomorphisms on cohomology groups
-$H^{i}\Big{(}X(\underline{v_{1}},\ldots,\underline{v_{r}}),\mathcal{O}_{X(\underline{v_{1}},\ldots,\underline{v_{r}})}\Big{)}\overset{\sim}{\longrightarrow}H^{i}\Big{(}{\overline{X}^{S^{*}}(v_{1}^{\prime}\cdots v_{r}^{\prime})},\mathcal{O}_{{\overline{X}^{S^{*}}(v_{1}^{\prime}\cdots v_{r}^{\prime})}}\Big{)},$
-for all $i\geq 0$. The $G^{F}$-equivariant isomorphisms for the cohomology of the canonical sheaf follow from Serre duality for Deligne–Mumford stacks [Nironi, Corollary 2.10].
 Remark 5.1.2.
-We use the same notations as in Theorem 5.1.1 and set $I={{\text{supp}}}_{F}(x)$. There are $G^{F}$-equivariant morphisms given by [DMR, Proposition 2.3.8]:
+We use the same notations as in Theorem˜5.1.1 and set $I={{\text{supp}}}_{F}(x)$. There are $G^{F}$-equivariant morphisms given by [8, Proposition 2.3.8]:
 ${\overline{X}^{S^{*}}(x)}\overset{\sim}{\longrightarrow}G^{F}/U_{I}^{F}\times^{L_{I}^{F}}{\overline{X}^{S^{*}}_{L_{I}}(x)}\overset{\pi_{I}}{\longrightarrow}G^{F}/P_{I}^{F},$
@@
 5.2. Suzuki and Ree groups
-We recall here the assumptions needed on $d$ and $q$ for which $(G,F)$ of type ${}^{2}B_{2}$, ${}^{2}F_{4}$, and ${}^{2}G_{2}$ can arise with $F(S)=S$.
-In addition to the assumptions Section 1.1, let $d=2$ and let $q$ be an odd power of $\sqrt{2}$ (resp.
+Here we recall the assumptions needed on $d$ and $q$ for $(G,F)$ of type $^{2}B_{2}$, $^{2}F_{4}$, and $^{2}G_{2}$.
+In addition to the assumptions Section˜1.1, let $d=2$ and let $q$ be an odd power of $\sqrt{2}$ (resp.
 $\sqrt{2}$, $\sqrt{3}$).
@@
 $F_{4}$, $G_{2}$).
-Then $F:G\to G$ as constructed in Section 1.1 satisfies $F(S)=S$ and $F^{2}$ is the $q^{2}$-Frobenius [DigneMichelBook, §4.3.7].
+Then the endomorphism $F:G\to G$ induces an automorphism of $W$ which preserves $S$ and $F^{2}$ is the $q^{2}$-Frobenius [9, §4.3.7].
 The group of fixed points $G^{F}$ is the Suzuki group $^{2}B_{2}(q)$ (resp.
@@
 Parts of the work are carried out when the author was a member of the trimester program "The Arithmetic of the Langlands Program" at the Hausdorff Research Institute for Mathematics in Bonn and when the author was a member of the Research Training Group 2553 in Essen, both funded by the DFG (Deutsche Forschungsgemeinschaft).
-[…]
-Fakultät für Mathematik, Universität Duisburg-Essen, 45127 Essen, Germany
-E-mail address: